### v1

# Chen, Yan, and Jiang

[…]
Chen, Yan, and Jiang
[…]
On the LP for Dynamic Stochastic Matching and Its Application on Pricing
[…]
On the Linear Programming Model for Dynamic Stochastic Matching and Its Application on Pricing
[…]
Important pricing problems in centralized matching markets—such as carpooling and food delivery platforms—often exhibit a bi-level structure. At the upper level, the platform sets prices for different types of agents (e.g., riders with various origins and destinations, or delivery orders from diverse restaurants to customer locations). The lower level then matches converted agents to minimize operational costs, such as pooling riders with similar itineraries into the same vehicle or consolidating multiple delivery orders into a single trip. Motivated by these applications, we study the optimal value (cost) function of a linear programming model with respect to agent arrival rates, originally proposed by Aouad and Sarıtaç (2022) for cost-minimizing dynamic stochastic matching under limited time. In particular, we investigate the concavity of this cost function. We show that weak concavity holds when there are fewer than or equal to three agent types with identical patience level. For an arbitrary number of agent types, (weak) concavity is more likely to hold when agents have high or similar patience, or when agents are more heterogeneous in their matching efficiencies. Building on these theoretical insights, we develop a Minorize-Maximization (MM) algorithm for the pricing problem that requires little stepsize tuning, and demonstrate substantial performance improvements over projected gradient-based methods on a large-scale, real-world ridesharing dataset. This makes it a compelling algorithmic choice for solving such pricing problems in practice.
[…]
dynamic stochastic matching, pricing, fluid relaxation, weak concavity, minorize-maximization
1 Introduction
Centralized platforms that perform dynamic, non-bipartite matching are increasingly prevalent in modern marketplaces, where different types of agents arrive stochastically and can be matched with other agents within a limited time. A representative example is shared ride services offered by ridesharing platforms such as Uber, Lyft and Didi, which match passengers with similar origin-destination (OD) pairs into shared trips. Due to limited patience, unmatched passengers may eventually be served individually (Yan, Yan, and Shen 2025). Another example is the online food delivery platform with batching decisions, where two or more orders with similar customer locations can be matched together and delivered by the same courier (Ma, Ma, and Romero 2025).
This work is motivated by the pricing challenges faced by such platforms—for example, in setting upfront fares for shared rides, where passengers have different ODs. Pricing and matching are inherently interdependent, naturally forming a bi-level optimization problem: at the upper level, the platform determines optimal prices to maximize profit, often expressed as a concave revenue component minus a total cost component; whereas at the lower level, the platform matches agents to minimize total costs. For a detailed discussion of this application, see Yan, Yan, and Shen (2025).
Problem Formulation.
Consider a system with $N$ possible agent types, where $[N]$ denotes the set ${1,2,\ldots,N}$. Potential agents of type $i\in[N]$ observe a quoted price $p_{i}$ and decide whether to enter the platform. Each type-$i$ agent arrives according to an independent Poisson process with arrival rate $\lambda_{i}>0$, which depends on the quoted price. Once admitted, a type-$i$ agent’s sojourn time in the system is exponentially distributed with rate $\theta_{i}$. Here, $\theta_{i}$ characterizes the patience of type-$i$ agents—a smaller $\theta_{i}$ indicates a greater willingness to wait for a match.
The platform’s objective is to maximize total profit by optimally setting prices and conducting matching for all agents. Following common assumptions in the literature, we assume that for each agent type $i\in[N]$, there is a one-to-one correspondence between the arrival rate $\lambda_{i}$ and the quoted price $p_{i}$, allowing us to express the price as a function of the arrival rate: $p_{i}(\lambda_{i})$. We further assume that the revenue function, $\lambda_{i}p_{i}(\lambda_{i})$, is concave in $\lambda_{i}$. Thus, the pricing problem reduces to finding the optimal vector $\boldsymbol{\lambda}=(\lambda_{i})_{i\in[N]}$ that maximizes total profit, formulated as
$\max_{\boldsymbol{\lambda}\in\Lambda}~{}g(\boldsymbol{\lambda})=\max_{\boldsymbol{\lambda}\in\Lambda}{\sum_{i\in[N]}~{}\lambda_{i}p_{i}(\lambda_{i})-c(\boldsymbol{\lambda})},$ (1)
where $\Lambda$ denotes the feasible domain of $\boldsymbol{\lambda}$, and $c(\boldsymbol{\lambda})$ is the optimal value (cost) function of the lower-level matching problem.
The lower-level matching problem has received a lot of attention in the literature recently. Existing studies often formulate it as a Markov Decision Process (MDP), under either a continuous-time setting (Aouad and Sarıtaç 2022) or a discrete-time setting (Ashlagi et al. 2023, Gupta 2024, Eom and Toriello 2024, Kerimov, Ashlagi, and Gurvich 2024, 2025). In the MDP framework, the state of the system is defined as the number of agents of each type waiting in the system to be matched. When an agent arrives, the platform can (i) immediately match her with another existing agent, or (ii) keep her waiting in the system. However, the optimal policies are in general intractable to obtain due to the curse of dimensionality. Therefore, many studies use fluid relaxations based on linear programs (LPs) to give bounds for the optimal value function. For example, for the dynamic non-bipartite matching problem, Eom and Toriello (2024), Gupta (2024) and Kerimov, Ashlagi, and Gurvich (2024, 2025) give reward-maximizing LP formulations under infinite agent patience. Collina et al. (2020) and Aouad and Sarıtaç (2022) formulate LPs incorporating limited patience. In this work, we adopt the cost-minimizing LP by Aouad and Sarıtaç (2022) to model $c(\boldsymbol{\lambda})$.
While existing studies focus on developing near-optimal matching policies using these LPs, our contribution lies in understanding how the optimal cost function $c(\boldsymbol{\lambda})$ varies as a function of the arrival rate, which is a crucial aspect of solving the pricing problem (1). This, to our best knowledge, has not been covered in existing studies where arrival rates are typically treated as fixed inputs.
With infinite patience (as in the LP formulation of Eom and Toriello 2024), it can be shown that $c(\boldsymbol{\lambda})$ is convex, making the pricing problem (1) a well-behaved concave maximization. However, when limited patience is taken into account, characterizing $c(\boldsymbol{\lambda})$ becomes highly nontrivial. As we will show in Section 2, $\boldsymbol{\lambda}$ appears in both the coefficients and the right-hand sides of the LP constraints, departing from known results in classical parametric LP literature (Gal and Nedoma 1972, Adler and Monteiro 1992). This makes the pricing problem (1) non-convex in general under limited agent patience.
Interestingly, in sharp contrast to the convexity of the optimal cost function under infinite agent patience, we find that $c(\boldsymbol{\lambda})$ often exhibits weak concavity when agent patience is limited. In particular, we show that weak concavity holds when there are fewer than or equal to three agent types with identical patience level. For an arbitrary number of agent types, (weak) concavity is more likely to hold when agents have high or similar patience, or when agents are more heterogeneous in their matching efficiencies.
Motivated by this observation, we develop a Minorize-Maximization (MM) algorithm (Hunter and Lange 2004), which provides convergence guarantees for concave-minus-concave maximization problems. Specifically, in program (1), the revenue term $\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})$ is concave, and when $c(\boldsymbol{\lambda})$ also exhibits concave properties, the MM algorithm can be effectively applied. Our proposed MM algorithm requires little stepsize tuning, making it easy to implement. A case study using the Chicago ridesharing dataset demonstrates that the MM algorithm significantly outperforms the projected gradient-based method, achieving both shorter running times and better optimality. Moreover, while the performance of the projected gradient method is highly sensitive to step-size tuning, the MM algorithm delivers consistently robust results, making it a compelling choice for solving such pricing problems in practice.
Organization. The remainder of the paper is organized as follows. We introduce the model and preliminaries in Section 2. We perform concave analysis on the value function in 3 and discuss its application on price optimization in 4. We conclude in Section 5. All proofs, auxiliary results, and additional examples are provided in the appendix.
2 Preliminaries
First, we introduce some additional terminology and notation. Following Aouad and Sarıtaç (2022), when two agents are matched, we refer to the agent who arrives earlier as the active agent and the one who arrives later as the passive agent. When an active agent of type $i\in[N]$ is matched with a passive agent of type $j\in[N]$, a cost $c_{(i,j)}>0$ is incurred. If an agent of type $i\in[N]$ remains unmatched when her patience runs out, a cost $c_{(i)}>0$ is incurred. Consistent with Yan, Yan, and Shen (2025), we make the following assumptions about these costs. {assumption} $c_{(i)}=c_{(i,i)}\leq c_{(i,j)}=c_{(j,i)},\ \forall i,j\in[N]$.
Assumption 2 states that when two agents of the same type $i\in[N]$ are matched, the resulting cost equals that of serving a single unmatched agent of type $i$. In the context of ridesharing, this means that matching two identical trips incurs the same cost as serving either trip individually. The condition $c_{(i,j)}=c_{(j,i)}$ imposes symmetry and requires that the order of arrival—that is, which agent is active or passive—does not affect the matched cost. Additionally, $c_{(i,j)}\geq c_{(i)}$ requires that the cost of serving two matched agents is at least as great as serving a single agent alone. This is a natural assumption in ridesharing, where a pooled trip must cover at least the distance required by one of the solo trips.
We also define the matching efficiency between two types of agents as follows.
Definition 2.1
The matching efficiency between agent types $i\in[N]$ and $j\in[N]$ is defined as $e_{i,j}=1-c_{(i,j)}/(c_{(i)}+c_{(j)})$.
Matching efficiency quantifies the proportion of cost saved when two types of agents are matched. A low matching efficiency indicates greater heterogeneity between the agents—for example, when their OD pairs differ significantly in ridesharing. By Assumption 2, $c_{(i,j)}\geq c_{(i)}$ and $c_{(i,j)}\geq c_{(j)}$, so $e_{i,j}$ is bounded above by $0.5$. Additionally, since $c_{(i,i)}=c_{(i)}$, we have $e_{i,i}=0.5$ for all $i\in[N]$, implying that self-matching always yields the highest efficiency. The lower bound of $e_{i,j}$ can be negative when $c_{(i,j)}>c_{(i)}+c_{(j)}$, meaning that matching these two types of agents would result in higher costs than serving them individually—rendering such matches impractical in any scenario.
The dynamic stochastic matching problem can be formulated as an MDP. In our study, instead of directly focusing on solving the MDP, we analyze the following LP:
$c(\boldsymbol{\lambda})=\min_{\boldsymbol{x},\boldsymbol{y}}\quad$ $\sum_{i\in[N]}\sum_{j\in[N]}c_{(i,j)}x_{i,j}+\sum_{i\in[N]}c_{(i)}y_{i}$ (2a)
s.t. $\sum_{j\in[N]}x_{j,i}+\sum_{j\in[N]}x_{i,j}+y_{i}=\lambda_{i},$ $\forall i\in[N],$ (2b)
$\theta_{i}x_{i,j}\leq\lambda_{j}y_{i},$ $\forall i,j\in[N],$ (2c)
$x_{i,j}\geq 0,$ $\forall i,j\in[N],$ (2d)
$y_{i}\geq 0,$ $\forall i\in[N],$ (2e)
where $\boldsymbol{\lambda}:=(\lambda_{i})_{i\in[N]}$, $\boldsymbol{x}:=(x_{i,j})_{i,j\in[N]}$, and $\boldsymbol{y}:=(y_{i})_{i\in[N]}$.
This LP model is often called the fluid relaxation of the MDP model. Variable $x_{i,j}$ represents the average match rate between type $i\in[N]$ active agent and type $j\in[N]$ passive agent, and $y_{i}$ represents the rate of type $i\in[N]$ agent unmatched. The objective function (2a) is the average total cost per unit of time. Constraints (2b) can be viewed as the flow balance equation, that is, in the steady state, the arrival rate of type-$i$ agent should be the summation of its active match rate, passive match rate, and unmatched rate. Constraints (2c) give the lower bound of the unmatched rate $y_{i}$ in terms of $\lambda_{j}$, $x_{i,j}$, and $\theta_{i}$. Intuitively, this reflects that a type $i$-th agent may leave the system before the next type $j$-th agent arrives, with a probability of $\theta_{i}/(\theta_{i}+\lambda_{j})$, thus $y_{i}$ should be a factor of at least $\theta_{i}/\lambda_{j}$ of the match rate $x_{i,j}$. Lastly, (2d) and (2e) require that the variables are nonnegative.
This LP was originally introduced in Aouad and Sarıtaç (2022), where it was shown that its optimal value provides a lower bound on the expected cost under the optimal matching policy. While Aouad and Sarıtaç (2022) treat $\boldsymbol{\lambda}$ as exogenously given, we instead view this LP as a parametric programming model in $\boldsymbol{\lambda}$. Formally, we consider $c(\boldsymbol{\lambda})$ as a function of $\boldsymbol{\lambda}$ defined over a compact interval $\Lambda=[\underline{\boldsymbol{\lambda}},\overline{\boldsymbol{\lambda}}]$. These lower and upper bounds can be interpreted as minimum and maximum prices imposed on different agent types. For convenience, we assume $\underline{\boldsymbol{\lambda}}>\boldsymbol{0}$, where all vector inequalities in this paper are interpreted component-wise.
{assumption}
$\boldsymbol{\lambda}\in\Lambda=[\underline{\boldsymbol{\lambda}},\overline{\boldsymbol{\lambda}}]$, where $\underline{\boldsymbol{\lambda}}>\boldsymbol{0}$.
We first observe that, as an extreme case, when all agents have infinite patience (i.e., $\theta_{i}=0$ for all $i\in[N]$), constraints (2c) become redundant, and the resulting LP closely resembles the formulation in Eom and Toriello (2024). In this scenario, $\boldsymbol{\lambda}$ appears only on the right-hand side of constraint (2b), and it is well known that $c(\boldsymbol{\lambda})$ becomes a convex, piecewise linear function (Adler and Monteiro 1992). Moreover, under Assumption 2, we can show that $c(\boldsymbol{\lambda})$ is, in fact, linear.
Proposition 2.2
When $\theta_{i}=0$, $\forall i\in[N]$, $c(\boldsymbol{\lambda})$ is linear in $\boldsymbol{\lambda}\in\Lambda$.
However, when constraint (2c) is imposed, analyzing $c(\boldsymbol{\lambda})$ becomes significantly more challenging. In this case, $\boldsymbol{\lambda}$ also appears in the coefficients of constraint (2c), making the LP fundamentally different from the classical parametric LP literature, where the parameter appears only on the right-hand side of the constraints (Gal and Nedoma 1972, Adler and Monteiro 1992). In fact, as shown in the following proposition, the cost function exhibits economies of scale, implying that $c(\boldsymbol{\lambda})$ tends to be concave rather than convex.
Proposition 2.3
[…]
Proposition 2.3 states that the average cost per agent decreases when the arrival rates are scaled proportionally. This implies that $c(\boldsymbol{\lambda})$ tends to bend downward (or remain linear) as the arrival rate increases—a behavior typically associated with concavity rather than convexity. However, since $c(\boldsymbol{\lambda})$ is a multivariate function, establishing concavity requires a more rigorous analysis.
3 Concavity of Value Function
In this section, we examine the concavity of $c(\boldsymbol{\lambda})$. We start with the simplest case when there is only one type of agent (i.e., $N=1$). It is straightforward to derive the explicit solution and show that the value function is concave under this case. Note that the optimal solutions to (2) depend on $\boldsymbol{\lambda}$ and can be written as $\boldsymbol{x}^{*}(\boldsymbol{\lambda})=(x_{i,j}^{*}(\boldsymbol{\lambda}))_{i,j\in[N]}$ and $\boldsymbol{y}^{*}(\boldsymbol{\lambda})=(y_{i}^{*}(\boldsymbol{\lambda}))_{i\in[N]}$, but for notational simplicity, we just use the abbreviations $\boldsymbol{x}^{*}=(x_{i,j}^{*})_{i,j\in[N]}$ and $\boldsymbol{y}^{*}=(y_{i}^{*})_{i\in[N]}$ throughout this section.
Lemma 3.1
When $N=1$, the optimal solution to (2) is given by $y_{1}^{*}=\lambda_{1}\theta_{1}/(\theta_{1}+2\lambda_{1})$ and $x_{1,1}^{*}=\lambda_{1}^{2}/(\theta_{1}+2\lambda_{1})$. The value function $c(\lambda_{1})$ is concave.
To prove Lemma 3.1, it suffices to show that the inequality in (2c) is always binding at optimality, implying that the platform should match as much as possible. While Lemma 3.1 establishes concavity in the simple case of $N=1$, this result does not directly extend to larger $N$. When $N\geq 2$, different agent types interact, and the total cost involves a sum of heterogeneous pairwise terms $(c_{(i,j)})_{i,j\in[N]}$, making the structure of $c(\boldsymbol{\lambda})$ much more complex—and potentially non-smooth. This complexity invalidates the straightforward concavity observed for $N=1$. For explicit counterexamples of non-concavity, see Example 7.1 (for $N=2$ with $\theta_{1}\neq\theta_{2}$) and Example 7.2 (for $N>2$ with $\theta_{1}=\theta_{2}=\cdots=\theta_{N}$), both in Appendix 7.
Given these counterexamples for $N\geq 2$, it may be unrealistic to rely solely on the standard definition of concavity. Therefore, we also consider the notion of weak concavity, a relaxed alternative defined below. See also Proposition 4.3 in Vial (1983) for an equivalent definition.
Definition 3.2
(Weak concavity) $c(\boldsymbol{\lambda})$ is said to be weakly concave in $\boldsymbol{\lambda}\in\Lambda$ if there exists a finite value $\rho\geq 0$, such that $c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2$ is concave in $\boldsymbol{\lambda}\in\Lambda$.
If $c(\boldsymbol{\lambda})$ is weakly concave, it can be written as the sum of a concave function, $c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2$, and a convex function, $\rho\|\boldsymbol{\lambda}\|^{2}/2$. This decomposition allows us to express the objective as $g(\boldsymbol{\lambda})=(\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})-\rho\|\boldsymbol{\lambda}\|^{2}/2)-(c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2)$, thereby restoring a concave-minus-concave structure that is useful for algorithm design in Section 4. We also emphasize that the weakly concave property is nontrivial, even for bounded functions; for example, one can verify that $|x|$ (for $-1\leq x\leq 1$) and $-\sqrt{x}$ (for $0\leq x\leq 1$) are not weakly concave.
We now state a fundamental lemma, which provides a sufficient condition under which the value function $c(\boldsymbol{\lambda})$ is weakly concave.
Lemma 3.3
If $\forall\boldsymbol{\lambda}\in\Lambda$, all optimal solutions satisfy $\boldsymbol{y}^{*}>\boldsymbol{0}$, then $c(\boldsymbol{\lambda})$ is weakly concave in $\boldsymbol{\lambda}\in\Lambda$.
When proving Lemma 3.3, we first establish the connection between global weak concavity and local weak concavity (see Lemma 6.1 in Appendix 6.4), that is, it is sufficient for us to prove that $c(\boldsymbol{\lambda})$ is weakly concave in a small neighborhood around any point $\hat{\boldsymbol{\lambda}}\in\Lambda$. To prove the local weak concavity, we examine the optimal basic feasible solution (BFS) of program (2) at $\hat{\boldsymbol{\lambda}}$ (Bertsimas and Tsitsiklis 1997). Using the assumption that $\boldsymbol{y}^{*}>\boldsymbol{0}$, we can prove that any optimal BFS still produces a basic feasible solution within a sufficiently small neighborhood around $\hat{\boldsymbol{\lambda}}$, and the corresponding value function remains both bounded and weakly concave. Using the fact that $c(\boldsymbol{\lambda})$ can be seen as a minimization of all candidate BFS, this local concavity property extends to $c(\boldsymbol{\lambda})$ itself.
Given Lemma 3.3, to establish the weak concavity of $c(\boldsymbol{\lambda})$, it suffices to identify conditions under which $\boldsymbol{y}^{*}>\boldsymbol{0}$. This allows us to bypass the complexity and potential non-smoothness of $c(\boldsymbol{\lambda})$ when $N\geq 2$. Moreover, restricting our attention to the case where $\boldsymbol{y}^{*}>\boldsymbol{0}$ is also practically justifiable—the case $y_{i}^{*}=0$ is an artifact of the LP relaxation and cannot be realized in the original MDP framework, as the unmatched rate should never be exactly zero when agents have limited patience (i.e., $\theta_{i}>0$ for all $i\in[N]$).
Leveraging Lemma 3.3, we present results for the case $N=2$ in Section 3.1, and extend to the general case for arbitrary $N$ in Section 3.2.
3.1 Analysis of Two Types of Agents
In this subsection, we examine the case where $N=2$. We show that the optimal solution to the LP can be fully characterized, and that weak concavity can be established by identifying the region where $\boldsymbol{y}^{*}>\boldsymbol{0}$. Without loss of generality, we assume $\theta_{1}\leq\theta_{2}$ throughout this subsection.
It is well known that whenever an LP admits an optimal solution, there always exists an optimal basic feasible solution (BFS) (Bertsimas and Tsitsiklis 1997). In any such optimal BFS, a subset of the inequality constraints—specifically, those in (2c), (2d), and (2e)—will be binding. We leverage this fact to characterize the structure of optimal solutions, that is, to determine which constraints are binding for different values of $\lambda_{1}$ and $\lambda_{2}$. For ease of exposition, we define:
$\Delta_{1}(\lambda_{1},\lambda_{2})$ $:=c_{(1)}\frac{\theta_{1}+\lambda_{1}}{\theta_{1}+2\lambda_{1}}+c_{(2)}\frac{\theta_{2}+\lambda_{2}}{\theta_{2}+2\lambda_{2}}-c_{(1,2)},$
$\Delta_{2}(\lambda_{1},\lambda_{2})$ $:=\frac{c_{(1)}}{2}\left[1-\frac{\theta_{1}(\theta_{2}+\lambda_{1}+2\lambda_{2})}{\lambda_{2}(\theta_{2}+2\lambda_{2})}\right]+c_{(2)}\frac{\theta_{2}+\lambda_{2}}{\theta_{2}+2\lambda_{2}}-c_{(1,2)},$
$\Delta_{3}(\lambda_{1},\lambda_{2})$ $:=\lambda_{1}-\lambda_{2}-\theta_{1}.$
Then the following proposition describes the optimal solution to (2).
Lemma 3.4
When $N=2$, an optimal solution to (2) is characterized by Table 1. In addition,
(i) all optimal solutions to (2) satisfy $y_{1}^{*}>0$;
(ii) all optimal solutions to (2) satisfy $y_{2}^{*}>0$ if and only if $\Delta_{2}(\lambda_{1},\lambda_{2})<0$ or $\Delta_{3}(\lambda_{1},\lambda_{2})<0$.
Table 1: Characterization of optimal solutions to program (2) when $N=2$
Case Conditions Optimal Solutions
$\Delta_{1}(\lambda_{1},\lambda_{2})$ $\Delta_{2}(\lambda_{1},\lambda_{2})$ $\Delta_{3}(\lambda_{1},\lambda_{2})$ $y_{1}^{*}$ $y_{2}^{*}$ $x_{1,1}^{*}$ $x_{1,2}^{*}$ $x_{2,1}^{*}$ $x_{2,2}^{*}$
(i) $\leq 0$ $>0$ $>0$ $=\lambda_{1}y_{1}^{*}/\theta_{1}$ $=0$ $=0$ $=\lambda_{2}y_{2}^{*}/\theta_{2}$
(ii) $>0$ $<0$ $>0$ $>0$ $=\lambda_{1}y_{1}^{*}/\theta_{1}$ $=\lambda_{2}y_{1}^{*}/\theta_{1}$ $=\lambda_{1}y_{2}^{*}/\theta_{2}$ $=\lambda_{2}y_{2}^{*}/\theta_{2}$
(iii) $\geq 0$ $<0$ $>0$ $>0$ $=0$ $=\lambda_{2}y_{1}^{*}/\theta_{1}$ $=\lambda_{1}y_{2}^{*}/\theta_{2}$ $=\lambda_{2}y_{2}^{*}/\theta_{2}$
(iv) $\geq 0$ $\geq 0$ $>0$ $=0$ $=(\lambda_{1}-y_{1}^{*}-x_{1,2}^{*})/2$ $=\lambda_{2}y_{1}^{*}/\theta_{1}$ $=0$ $=0$
Lemma 3.4 identifies four possible types of optimal solutions, determined by the signs of three indicator functions. When $x_{i,j}^{*}=0$, the corresponding constraint in (2d) is binding; when $x_{i,j}^{\ast}=\lambda_{j}y_{i}^{\ast}/\theta_{i}$, the corresponding constraint in (2c) is binding. The first solution category corresponds to the case with no cross-matching between type 1 and type 2 agents (i.e., $x_{1,2}^{\ast}=x_{2,1}^{\ast}=0$). The second category is the fully-matched case, where all types of matchings occur and all constraints in (2c) are binding. In the third category, there is no self-matching among type 1 agents ($x_{1,1}^{\ast}=0$). The last category is characterized by $y_{2}^{\ast}=0$, where all type 2 agents are passively matched with type 1 agents. Using Lemma 3.4 and flow constraints (2b), we can explicitly give the expression of $y_{1}^{*}$ and $y_{2}^{*}$, thereby obtaining the optimal solutions in closed forms. The complete closed-form expressions are provided at the end of Appendix 6.4. Lemma 3.4 also helps us to identify the region where $c(\boldsymbol{\lambda})$ is weakly concave. Based on the results of Lemma 3.3 and Lemma 3.4, we see that if, for every $(\lambda_{1},\lambda_{2})\in\Lambda$, at least one of the inequalities $\Delta_{2}(\lambda_{1},\lambda_{2})<0$ or $\Delta_{3}(\lambda_{1},\lambda_{2})<0$ holds, then $y_{1}^{\ast},y_{2}^{\ast}>0$ is guaranteed. The following theorem further explores this condition and gives an explicit characterization of $\Lambda$ satisfying such conditions.
Theorem 3.5
Define
$\tau_{1}$ $:=c_{(1)}(\theta_{2}-3\theta_{1})+2c_{(2)}\theta_{2}-2c_{(1,2)}\theta_{2},$
$\tau_{2}$ $:=\tau_{1}^{2}-8c_{(1)}(2c_{(1,2)}-c_{(1)}-c_{(2)})\theta_{1}(\theta_{1}+\theta_{2}).$
(i) When $\tau_{1}\leq 0$ or $\tau_{2}<0$, $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$;
(ii) When $\tau_{1}>0$ and $\tau_{2}\geq 0$, $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$ if $2c_{(1,2)}-c_{(1)}-c_{(2)}\neq 0$ and
$\underline{\lambda}_{2}>\frac{\tau_{1}+\sqrt{\tau_{2}}}{4(2c_{(1,2)}-c_{(1)}-c_{(2)})}.$
Theorem 3.5 gives two sufficient conditions for weak concavity, depending on the signs of $\tau_{1}$ and $\tau_{2}$. Smaller values of $\tau_{1}$ and $\tau_{2}$ are desired to achieve weak concavity. The expressions of $\tau_{1}$ and $\tau_{2}$ lead to the following insights: (i) Since both $\tau_{1}$ and $\tau_{2}$ are decreasing in $\theta_{1}$, a larger value of $\theta_{1}$ (i.e., when $\theta_{1}$ and $\theta_{2}$ are closer since $\theta_{1}\leq\theta_{2}$) helps to achieve weak concavity in the larger region; (ii) Similarly, since both $\tau_{1}$ and $\tau_{2}$ are decreasing in $c_{(1,2)}$, a larger value of $c_{(1,2)}$ also promote weak concavity. Although Theorem 3.5 provides only sufficient conditions, we find that the resulting lower bound for $\underline{\lambda}_{2}$ is quite tight. In Appendix 7, we present a counterexample (Example 7.3) demonstrating that as soon as $\lambda_{2}$ falls just below our derived bound, weak concavity is already violated.
Note that in the second case of Theorem 3.5, we require $2c_{(1,2)}-c_{(1)}-c_{(2)}\neq 0$, or equivalently, $e_{1,2}\neq 0.5$. This condition means that the two agent types cannot have perfect matching efficiency. In the context of ridesharing, this essentially precludes the possibility that both rider types have exactly the same OD pairs. When $e_{1,2}=0.5$ and $\theta_{1}=\theta_{2}$, this is not an issue, as the platform can simply merge the two types into a single type. However, as the following proposition shows, if $\theta_{1}\neq\theta_{2}$, the region where $y_{2}^{*}=0$ becomes unbounded, making it impossible to establish any lower bound on $\underline{\boldsymbol{\lambda}}$ that guarantees $\boldsymbol{y}^{*}>\boldsymbol{0}$. Appendix 7 provides a counterexample (Example 7.4) demonstrating that weak concavity can easily be violated when $e_{1,2}=0.5$ and $\theta_{1}\neq\theta_{2}$, even for large values of $\lambda_{1}$ and $\lambda_{2}$.
Proposition 3.6
When $\tau_{1}>0$ and $02c_{(1,2)}-c_{(1)}-c_{(2)}=$ 0, given any $\underline{\boldsymbol{\lambda}}>\boldsymbol{0}$, there exists $\boldsymbol{\lambda}>\underline{\boldsymbol{\lambda}}$ such that $y_{2}^{*}=0$.
Given the complexity of the expressions for $\tau_{1}$ and $\tau_{2}$ in Theorem 3.5, we also provide the following corollary, which presents a more tractable sufficient condition in a simplified form.
Corollary 3.7
(i) When $\theta_{2}<3\theta_{1}$, $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$;
(ii) When $\theta_{2}\geq 3\theta_{1}$, $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$ if $e_{1,2}\neq 0.5$ and
$\underline{\lambda}_{2}>\frac{\theta_{2}}{4}\frac{1}{1-2e_{1,2}}.$
According to Corollary 3.7, the function is guaranteed to be weakly concave as long as $\theta_{2}<3\theta_{1}$. If the two agent types have significantly different patience levels, we require a lower bound on $\underline{\lambda}_{2}$. This bound depends on $\theta_{2}$ (the patience of agent type 2) and $e_{1,2}$ (the matching efficiency between the two agent types). A smaller $\theta_{2}$ (greater patience for both types of agents since $\theta_{1}\leq\theta_{2}$) and a lower $e_{1,2}$ (lower matching efficiency) both expand the region of $\Lambda$ where weak concavity is maintained.
3.1.1 Special case of same patience level.
We now consider a special case where two types of agents have the same patience level (that is, $\theta_{1}=\theta_{2}=\theta$). Under this condition, the structure of optimal solutions simplifies considerably compared to the general case described in Lemma 3.4, as stated in the following corollary.
Corollary 3.8
When $N=2$ and $\theta_{1}=\theta_{2}=\theta$, an optimal solution to program (2) is characterized by Table 2.
Table 2: Characterization of optimal solutions to program (2) when $N=2$ and $\theta_{1}=\theta_{2}=\theta$
Case $\Delta_{1}(\lambda_{1},\lambda_{2})$ Optimal Solutions
$y_{1}^{*}$ $y_{2}^{*}$ $x_{1,1}^{*}$ $x_{1,2}^{*}$ $x_{2,1}^{*}$ $x_{2,2}^{*}$
(i) $\leq 0$ $=\lambda_{1}\theta/(\theta+2\lambda_{1})$ $=\lambda_{2}\theta/(\theta+2\lambda_{2})$ $=\lambda_{1}y_{1}^{*}/\theta$ $=0$ $=0$ $=\lambda_{2}y_{2}^{*}/\theta$
(ii) $>0$ $=\lambda_{1}\theta/(\theta+2\lambda_{1}+2\lambda_{2})$ $=\lambda_{2}\theta/(\theta+2\lambda_{1}+2\lambda_{2})$ $=\lambda_{1}y_{1}^{*}/\theta$ $=\lambda_{2}y_{1}^{*}/\theta$ $=\lambda_{1}y_{2}^{*}/\theta$ $=\lambda_{2}y_{2}^{*}/\theta$
Corollary 3.8 states that when $\theta_{1}=\theta_{2}$, there are only two possible categories of optimal solutions, determined solely by the sign of $\Delta_{1}(\lambda_{1},\lambda_{2})$. In this special case, the last two solution categories described in Lemma 3.4 no longer arise. Intuitively, when $c_{(1,2)}$ is much larger than $c_{(1)}$ and $c_{(2)}$, the optimal strategy is to match agents only within their own types; otherwise, a mix of self- and cross-matching becomes optimal. The arrival rate and patience levels also influence the solution: higher arrival rates (i.e., larger $\lambda_{1}$ and $\lambda_{2}$) or greater patience (i.e., smaller $\theta$) both tend to reduce the extent of cross-matching.
Since $\boldsymbol{y}^{*}>\boldsymbol{0}$ always holds when $\theta_{1}=\theta_{2}$, the weak concavity of $c(\boldsymbol{\lambda})$ follows directly from Lemma 3.3. Moreover, using the explicit form of the optimal solution given by Corollary 3.8, we can further verify that $c(\boldsymbol{\lambda})$ is, in fact, concave.
Theorem 3.9
When $N=2$ and $\theta_{1}=\theta_{2}=\theta$, $c(\boldsymbol{\lambda})$ is concave in $\boldsymbol{\lambda}\in\Lambda$.
3.2 Analysis of Multiple Types of Agents
We now present results for the general case with an arbitrary number of agent types. As $N$ increases, fully characterizing the optimal solutions becomes intractable. However, by leveraging insights from previous results and their proofs, we can identify sufficient conditions for ensuring $\boldsymbol{y}^{*}>\boldsymbol{0}$ for general $N$, and once again apply Lemma 3.3 to establish weak concavity. For clarity, we assume throughout this section that $\theta_{1}\leq\theta_{2}\leq\cdots\leq\theta_{N}$.
Given any type $i$, we first sort its corresponding matching efficiency $e_{i,1},e_{i,2},\cdots,e_{i,N}$ in descending order and denote the sorted result as $e_{i,(1)}\geq e_{i,(2)}\geq\cdots\geq e_{i,(N)}$, where by definition $e_{i,(1)}=e_{i,i}=0.5$. For $k\in\mathbb{Z}^{+}$, we define the $k$-th critical matching efficiency as
$e_{(k)}=\begin{cases}\max\limits_{i\in[N]}\left\{e_{i,(k)}\right\},&\text{if }k\leq N,\\ 0,&\text{if }k>N,\end{cases}$
that is, when $k\leq N$, $e_{(k)}$ is the maximum $k$-th largest matching efficiency among all agent types. Since each agent has a self-matching efficiency of $0.5$, we always have $e_{(1)}=0.5$. When $k\geq 2$, this implies that for any agent of type $i\in[N]$, there are at most $k-2$ other agent types (excluding self-match type $i$) with a matching efficiency greater than $e_{(k)}$. For notational convenience, we also define $e_{(k)}=0$ for all $k>N$.
Based on the above definition, we now state our main results concerning concavity.
Theorem 3.10
(i) When $\theta_{1}=\theta_{2}=\cdots=\theta_{N}=\theta$, $c(\boldsymbol{\lambda})$ is concave in $\Lambda$ if $e_{(3)}\neq 0.5$ and
$\underline{\lambda}_{i}>\theta\frac{e_{(3)}}{1-2e_{(3)}},\ \forall i\in[N];$
(ii) When $\theta_{1}\neq\theta_{N}$, $c(\boldsymbol{\lambda})$ is concave in $\Lambda$ if $e_{(2)}\neq 0.5$ and
$\underline{\lambda}_{i}>\theta_{i}\frac{e_{(2)}}{1-2e_{(2)}},\ \forall i\in[N].$
Theorem 3.10 discusses two separate cases under which concavity holds for general $N$. In the first case, where all $\theta_{i},\ i\in[N]$ are equal, the lower bound for $\underline{\lambda}_{i}$ depends on $e_{(3)}$. Intuitively, when $e_{(3)}$ is small, each agent type $i$ has high matching efficiency with at most one other type $j$, while the remaining $N-2$ types have low compatibility with $i$. This structure ensures that agents of type $i$ are matched with at most one other type, allowing us to reduce the problem to a series of two-type ($N=2$) subproblems. As a result, the concavity result from Theorem 3.9 can be extended to this setting.
For the second case, where the $\theta_{i}$ may differ, we can show that each agent type is optimally matched only within its own type when the lower bound of $\underline{\lambda}_{i}$ is satisfied. This further simplifies the problem to the single-type case ($N=1$), for which concavity has been established in Lemma 3.1.
The next theorem gives the results related to weak concavity.
Theorem 3.11
(i) When $\theta_{1}=\theta_{2}=\cdots=\theta_{N}=\theta$, $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$ if $e_{(4)}\neq 0.5$ and
$\underline{\lambda}_{i}>\theta\frac{e_{(4)}}{1-2e_{(4)}},\ \forall i\in[N];$
(ii) When $\theta_{N}<2\theta_{1}$, $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$ if $e_{(3)}\neq 0.5$ and
$\underline{\lambda}_{i}>\theta_{i}\frac{e_{(3)}}{1-2e_{(3)}},\ \forall i\in[N].$
Similar to Theorem 3.10, we also require a non-zero lower bound of $\underline{\lambda}_{i}$ depending on the critical matching efficiency $e_{(3)}$ or $e_{(4)}$ in Theorem 3.11. For the first case, we continue to assume that $\theta_{i},\ i\in[N]$ are all equal, but allow the region $\Lambda$ being larger than that in Theorem 3.10. Specifically, the lower bound of $\underline{\lambda}_{i}$ is now related to $e_{(4)}$ instead of $e_{(3)}$. Under this weaker assumption, we can prove that the value function is weakly concave, if not concave. In the second case, we assume the values of $\theta_{i},\ i\in[N]$ can be different but within a ratio of 2, and also gives a condition of weak concavity. Notably, the allowable region is also enlarged compared to Theorem 3.10, since the lower bound for $\underline{\lambda}_{i}$ is determined by $e_{(3)}$ instead of $e_{(2)}$. Both results generalize the key steps used in the proof of Lemma 3.4.
From the first case of Theorem 3.11, we immediately get the following corollary for the case of $N=3$, since $e_{(4)}=0$ according to our definition.
Corollary 3.12
When $N=3$ and $\theta_{1}=\theta_{2}=\theta_{3}$, $c(\boldsymbol{\lambda})$ is weakly concave in $\boldsymbol{\lambda}\in\Lambda$.
Both Theorem 3.10 and Theorem 3.11 state that it is easier to achieve (weak) concavity when all agents have the same patience level. A higher patience level also leads to a smaller required lower bound of $\underline{\lambda}_{i}$ for each $i\in[N]$. As an extreme case, when $\theta_{i}=0$, $\forall i\in[N]$, the value function will become linear, as we previously established in Proposition 2.2.
4 Application on Pricing
In this section, we leverage the (weak) concavity of $c(\boldsymbol{\lambda})$ to design an efficient algorithm for solving the upper-level pricing problem. We assume a one-to-one mapping between $\lambda_{i}$ and $p_{i}$ for all $i\in[N]$, as well as concavity of the revenue function with respect to the arrival rate. Both assumptions are standard in the revenue management literature. {assumption} There exists a one-to-one correspondence between $\lambda_{i}$ and $p_{i}$ for all $i\in[N]$. The revenue function $\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})$ is concave in $\boldsymbol{\lambda}\in\Lambda$.
4.1 A Minorize-Maximization Algorithm
Although the revenue function is assumed to be concave, $c(\boldsymbol{\lambda})$ is in general not convex, thus, the overall program (1) is not a convex optimization problem. As shown in Example 7.5, $g(\boldsymbol{\lambda})$ is neither unimodal nor smooth. However, as analyzed in Section 3, $c(\boldsymbol{\lambda})$ tends to be concave or weakly concave. With such properties, we can treat the objective function $g(\boldsymbol{\lambda})$ as a difference of two concave functions:
$g(\boldsymbol{\lambda})=\left(\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})-\frac{1}{2}\rho\|\boldsymbol{\lambda}\|^{2}\right)-\left(c(\boldsymbol{\lambda})-\frac{1}{2}\rho\|\boldsymbol{\lambda}\|^{2}\right),$
where the value of $\rho\geq 0$ is selected such that $c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2$ is concave. We can then adapt the Minorize-Maximization (MM) algorithm to obtain the solution with a convergence guarantee. At each step of the MM algorithm, we construct a surrogate linear function to replace the concave function $c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2$. Denote $\boldsymbol{\lambda}^{(t)}$ as the value of $\boldsymbol{\lambda}$ after the $t$-th iteration. At the $(t+1)$-th iteration, the surrogate linear function is generated based on the supergradient of this concave function (analogous to the subgradient in convex functions; see Lu et al. 2014) at the point of $\boldsymbol{\lambda}^{(t)}$. Let $\boldsymbol{\gamma}^{*}=(\gamma_{i}^{*})_{i\in[N]}$ and $\boldsymbol{\eta}^{*}=(\eta_{i,j}^{*})_{i,j\in[N]}$ be the optimal dual variables corresponding to (2b) and (2c) respectively, then according to the envelop theorem (see e.g., Corollary 5 in Milgrom and Segal 2002), the supergradient, denoted as $\boldsymbol{v}=(v_{i})_{i\in[N]}$, can be given as
$\boldsymbol{v}:=\left(\gamma_{i}^{*}+\sum_{j\in[N]}y_{j}^{*}\eta_{j,i}^{*}-\rho\lambda_{i}^{(t)}\right)_{i\in[N]}\in\partial\left(c(\boldsymbol{\lambda}^{(t)})-\frac{1}{2}\rho\|\boldsymbol{\lambda}^{(t)}\|^{2}\right),$ (3)
and the surrogate linear function is given as
$c(\boldsymbol{\lambda}^{(t)})-\frac{1}{2}\rho\|\boldsymbol{\lambda}^{(t)}\|^{2}+\boldsymbol{v}^{T}\left(\boldsymbol{\lambda}-\boldsymbol{\lambda}^{(t)}\right)\geq c(\boldsymbol{\lambda})-\frac{1}{2}\rho\|\boldsymbol{\lambda}\|^{2},\quad\forall\boldsymbol{\lambda}\in\Lambda,$
where the inequality holds because $c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2$ is a concave function.
Define $Q(\boldsymbol{\lambda}\mid\boldsymbol{\lambda}^{(t)})=\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})-\rho\|\boldsymbol{\lambda}\|^{2}/2-c(\boldsymbol{\lambda}^{(t)})+\rho\|\boldsymbol{\lambda}^{(t)}\|^{2}/2-\boldsymbol{v}^{T}(\boldsymbol{\lambda}-\boldsymbol{\lambda}^{(t)})$. It can be easily verified that $Q(\boldsymbol{\lambda}\mid\boldsymbol{\lambda}^{(t)})$ is a concave function since the original non-concave component in $g(\boldsymbol{\lambda})$ have been substituted with a linear function, and we have
$Q(\boldsymbol{\lambda}\mid\boldsymbol{\lambda}^{(t)})$ $=\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})-\frac{1}{2}\rho\|\boldsymbol{\lambda}\|^{2}-c(\boldsymbol{\lambda}^{(t)})+\frac{1}{2}\rho\|\boldsymbol{\lambda}^{(t)}\|^{2}-\boldsymbol{v}^{T}(\boldsymbol{\lambda}-\boldsymbol{\lambda}^{(t)})$
$\leq\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})-\frac{1}{2}\rho\|\boldsymbol{\lambda}\|^{2}-\left(c(\boldsymbol{\lambda})-\frac{1}{2}\rho\|\boldsymbol{\lambda}\|^{2}\right)=g(\boldsymbol{\lambda}),\quad\forall\boldsymbol{\lambda}\in\Lambda,$
and
$Q(\boldsymbol{\lambda}^{(t)}\mid\boldsymbol{\lambda}^{(t)})=g(\boldsymbol{\lambda}^{(t)}).$
We update the value of $\boldsymbol{\lambda}$ by letting $\boldsymbol{\lambda}^{(t+1)}\leftarrow\arg\max_{\boldsymbol{\lambda}\in\Lambda}Q(\boldsymbol{\lambda}\mid\boldsymbol{\lambda}^{(t)})$, where the maximization of $Q(\boldsymbol{\lambda}\mid\boldsymbol{\lambda}^{(t)})$ can be solved using standard convex optimization techniques. Then we have
$g(\boldsymbol{\boldsymbol{\lambda}}^{(t+1)})\geq Q(\boldsymbol{\lambda}^{(t+1)}\mid\boldsymbol{\lambda}^{(t)})\geq Q(\boldsymbol{\lambda}^{(t)}\mid\boldsymbol{\lambda}^{(t)})=g(\boldsymbol{\lambda}^{(t)}),$
which guarantees that the objective function value is non-decreasing at each iteration. This monotonic improvement ensures monotone convergence to a local optimum (Hunter and Lange 2004).
Algorithm 1 Minorize-Maximization (MM) Algorithm
1:Input: Initial guess $\boldsymbol{\lambda}^{(0)}\in\Lambda$, $\varepsilon>0$, $\delta_{\text{MM}}>0$.
2:Initialize: $t\leftarrow 0$
3:repeat
4:     Solve $c(\boldsymbol{\lambda}^{(t)})$, obtain optimal primal variables $\boldsymbol{y}^{*}=(y_{i}^{*})_{i\in[N]}$, $\boldsymbol{x}^{*}=(x_{i,j}^{*})_{i,j\in[N]}$ and dual variables $\boldsymbol{\gamma}^{*}=(\gamma_{i}^{*})_{i\in[N]}$, $\boldsymbol{\eta}^{*}=(\eta_{i,j}^{*})_{i,j\in[N]}$.
5:     $\rho\leftarrow 0$
6:     repeat
7:         Calculate $\boldsymbol{v}=(v_{i})_{i\in[N]}$, where $v_{i}=\gamma_{i}^{*}+\sum_{j\in[N]}y_{j}^{*}\eta_{j,i}^{*}-\rho\lambda_{i}^{(t)}.$
8:         Define $Q(\boldsymbol{\lambda}\mid\boldsymbol{\lambda}^{(t)})=\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})-\rho\|\boldsymbol{\lambda}\|^{2}/2-c(\boldsymbol{\lambda}^{(t)})+\rho\|\boldsymbol{\lambda}^{(t)}\|^{2}/2-\boldsymbol{v}^{T}(\boldsymbol{\lambda}-\boldsymbol{\lambda}^{(t)})$.
9:         Update $\boldsymbol{\lambda}^{(t+1)}\leftarrow\arg\max_{\boldsymbol{\lambda}\in\Lambda}Q(\boldsymbol{\lambda}\mid\boldsymbol{\lambda}^{(t)})$.
10:         if $g(\boldsymbol{\lambda}^{(t+1)})<g(\boldsymbol{\lambda}^{(t)})$ then
11:              Update $\rho\leftarrow\rho+\delta_{\text{MM}}$.
12:         end if
13:     until $g(\boldsymbol{\lambda}^{(t+1)})\geq g(\boldsymbol{\lambda}^{(t)})$
14:     $t\leftarrow t+1$
15:until $|g(\boldsymbol{\lambda}^{(t+1)})-g(\boldsymbol{\lambda}^{(t)})|<\varepsilon$
16:Output: $\boldsymbol{\lambda}^{(t)}$
Summarizing the ideas above, we present the pseudocode of the MM algorithm in Algorithm 1. Note that, as shown in Section 3, although the cost function $c(\boldsymbol{\lambda})$ is weakly concave under certain conditions, the value of $\rho$ is not specified. To address this, we slightly modify the standard MM algorithm by introducing a search step for $\rho$. Initially, we set $\rho=0$ (Line 5) and perform one iteration to update $\boldsymbol{\lambda}$ (Lines 7–9). If the objective function does not increase after this iteration, it indicates that $c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2$ is not concave; we then increment $\rho$ by $\delta_{\text{MM}}$ and repeat the update. Since $c(\boldsymbol{\lambda})$ is weakly concave, it is guaranteed that after a finite number of increments, we will find a $\rho$ such that $c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2$ is concave, ensuring that the objective function $g(\boldsymbol{\lambda})$ increases after each iteration. By construction, each iteration yields an increase in the objective, and thus the algorithm is guaranteed to converge. It is worth noting that in the ridesharing case study presented in the next subsection, we never observe the need to increase $\rho$; Algorithm 1 consistently converges with $\rho=0$, requiring no stepsize tuning on real data. The same behavior is observed across numerous synthetically generated instances. In contrast, alternative methods for solving (2), such as the projected gradient-based approach, are highly sensitive to stepsize selection and require careful tuning, as will be demonstrated in Section 4.2.
4.2 Chicago Ridesharing Case Study
In this subsection, we conduct numerical experiments using the Chicago ridesharing dataset. The dataset, provided by Chicago Data Portal, [^1] https://data.cityofchicago.org/Transportation/Transportation-Network-Providers-Trips-2018-2022-/m6dm-c72p/about_data contains trips reported by ridesharing companies within the city of Chicago. The dataset divides the city (excluding the O’Hare International Airport) into 76 community areas, and each data point includes the passenger’s OD pair, represented by the corresponding communities’ centroid locations. Same as Yan, Yan, and Shen (2025), we focus on the typical rush-hour data (i.e., 7:30 AM - 8:30 AM on each Monday) in October and November 2019, before the COVID-19 pandemic, to conduct the experiment. In total seven weeks are selected as offline data, for the purpose of solving program (1) and obtaining the optimal static price.
To aggregate riders into distinct types, we apply the $K$-means clustering algorithm to the origin-destination (OD) pairs of all trips, with the number of clusters set to $N\in\{100,150,200\}$. For each cluster (i.e., rider type $i\in[N]$), we treat the cluster center as the representative OD for that type, and the average number of requests per hour within the cluster as the maximum possible hourly arrival rate $\overline{\lambda}_{i}$. Given the OD of each rider type, we compute the corresponding solo trip distances and pooled trip distances (in miles), denoted as $[l_{(i)}]_{i\in[N]}$ and $[l_{(i,j)}]_{i,j\in[N]}$, respectively. We let $c_{(i,j)}=cl_{(i,j)}$, where $c$ is the per-mile cost. We consider three possible values of $c\in\{0.7,0.9,1.1\}$ in the experiment, with a unit of dollar per mile. We assume in this case study that the riders’ per-mile willingness to pay is uniformly distributed in $[0,1]$, then we have $p_{i}(\lambda_{i})=l_{(i)}\left(1-\lambda_{i}/\overline{\lambda}_{i}\right)$, $\forall i\in[N]$.
We compare the following two methods to solve the pricing problem (1):
  • Projected Gradient (PG) method. We use PG as the baseline method. At each iteration, $\boldsymbol{\lambda}$ is updated along with the opposite direction of the gradient while restricting within $\Lambda$. When $g(\boldsymbol{\lambda})$ is differentiable at $\boldsymbol{\lambda}$, its gradient is given by
    $\nabla g(\boldsymbol{\lambda})=\left[l_{(i)}\left(1-\frac{2\lambda_{i}}{\overline{\lambda}_{i}}\right)-\gamma_{i}^{*}-\sum_{j\in[N]}y_{j}^{*}\eta_{j,i}^{*}\right]_{i\in[N]},$ (4)
    where $\boldsymbol{\gamma}^{*}=(\gamma_{i}^{*})_{i\in[N]}$ and $\boldsymbol{\eta}^{*}=(\eta_{i,j}^{*})_{i,j\in[N]}$ are the optimal dual variables as defined in Algorithm 1. If $g(\boldsymbol{\lambda})$ is non-differentiable, we continue to use (4) to perform the iteration, interpreting it as a subgradient or supergradient when such gradients exist. The initial step size $\delta_{\text{PG}}$ is chosen as $\{1,10,100\}$. We also dynamically adjust the step size, by checking whether the objective function increases after each iteration. If $g(\boldsymbol{\lambda}^{(t+1)})<g(\boldsymbol{\lambda}^{(t)})$, we decrease the step size by half.
  • Minorize-Maximization (MM) method. This is our Algorithm 1. Note that since we now have an explicit form of $p_{i}(\lambda_{i})$, we can give a closed form solution of $\boldsymbol{\lambda}^{(t+1)}\leftarrow\arg\max_{\boldsymbol{\lambda}\in\Lambda}Q(\boldsymbol{\lambda}\mid\boldsymbol{\lambda}^{(t)})$ (Line 9 of Algorithm 1) using its first-order condition:
    $\lambda_{i}^{(t+1)}=\max\left\{\underline{\lambda}_{i},\min\left\{\overline{\lambda}_{i},\overline{\lambda}_{i}\frac{l_{(i)}-\left(\gamma_{i}^{*}+\sum_{j\in[N]}y_{j}^{*}\eta_{j,i}^{*}\right)+\rho\lambda_{i}^{(t)}}{2l_{(i)}+\rho\overline{\lambda}_{i}}\right\}\right\}.$
    Besides, since we never observe the need to increase $\rho$ in our experiments—i.e., Algorithm 1 consistently converges with $\rho=0$—the choice of $\delta_{\text{MM}}$ in Algorithm 1 is not relevant.
For both algorithms, we set $\underline{\lambda}_{i}=10^{-3}$ for all $i\in[N]$. Initial solution $\boldsymbol{\lambda}^{(0)}$ is randomly sampled from the feasible set $\Lambda=[\underline{\boldsymbol{\lambda}},\overline{\boldsymbol{\lambda}}]$ and kept the same across all algorithms within the same run. The stopping criterion is defined as $|g(\boldsymbol{\lambda}^{(t+1)})-g(\boldsymbol{\lambda}^{(t)})|<\varepsilon=10^{-3}$. We also record the running time of each algorithm; if it exceeds 20 minutes, the algorithm terminates and returns the result from the latest iteration. We report and compare the following three metrics, averaged over three different random seeds [^2] The randomness comes from the initialization of $K$-means clustering, as well as the sampling of $\boldsymbol{\lambda}^{(0)}$ and $(\theta_{i})_{i\in[N]}$.: (i) running time (in seconds); (ii) number of iterations; and (iii) final objective value.
Table 3: Numerical results of solving (1) when $\theta_{1}=\cdots=\theta_{N}=\theta$.
$c$ (\$/mile) $0.7$ $0.9$ $1.1$
$\theta$ (min-1) $1/5$ $1/3$ $1$ $2$ $1/5$ $1/3$ $1$ $2$ $1/5$ $1/3$ $1$ $2$
$N=100$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 908 921 634 637 687 692 699 740 699 834 794 810
PG ($\delta_{\text{PG}}=10$) 332 415 375 374 304 321 298 330 704 863 657 633
PG ($\delta_{\text{PG}}=100$) 396 366 391 416 346 349 354 350 806 867 800 660
MM 64 68 59 63 63 71 75 79 124 158 194 171
Num. Iterations PG ($\delta_{\text{PG}}=1$) 70.3 72.3 47.3 48.3 57.7 59.7 60.3 57.0 56.7 70.7 67.3 68.0
PG ($\delta_{\text{PG}}=10$) 24.7 32.7 27.7 28.0 25.0 26.3 25.3 27.0 54.0 68.7 54.7 54.0
PG ($\delta_{\text{PG}}=100$) 29.3 28.7 30.0 30.3 28.3 29.0 30.0 30.0 59.0 65.3 67.0 55.3
MM 5.3 5.7 5.0 5.3 5.3 6.0 6.3 6.7 10.3 13.3 16.3 14.7
Objective Function PG ($\delta_{\text{PG}}=1$) 64.13 58.92 45.75 37.96 39.45 33.70 20.65 13.88 16.68 13.05 4.69 1.63
PG ($\delta_{\text{PG}}=10$) 64.33 59.11 45.93 38.14 39.64 33.89 20.85 14.11 13.11 10.21 4.31 1.64
PG ($\delta_{\text{PG}}=100$) 64.33 59.11 45.93 38.14 39.64 33.89 20.85 14.11 12.58 10.08 4.43 1.64
MM 64.33 59.11 45.93 38.14 39.64 33.89 20.84 14.11 19.59 14.30 4.97 1.71
$N=150$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 829 747 551 658 720 748 756 542 1200 829 919 735
PG ($\delta_{\text{PG}}=10$) 772 743 710 758 737 760 765 760 1036 959 1099 994
PG ($\delta_{\text{PG}}=100$) 1008 949 921 903 886 896 936 962 1200 1200 1024 1133
MM 132 133 133 134 159 179 170 162 411 402 446 292
Num. Iterations PG ($\delta_{\text{PG}}=1$) 30.7 27.7 20.7 22.3 25.0 27.7 28.7 20.0 44.3 30.3 33.7 25.3
PG ($\delta_{\text{PG}}=10$) 26.7 26.7 26.7 27.0 26.3 27.3 28.3 28.0 39.3 35.3 40.0 34.7
PG ($\delta_{\text{PG}}=100$) 33.7 33.3 33.3 33.3 32.0 31.7 34.3 35.0 44.7 43.7 38.0 39.7
MM 5.0 5.0 5.0 5.0 6.0 6.7 6.3 6.0 15.7 15.3 17.0 11.0
Objective Function PG ($\delta_{\text{PG}}=1$) 61.55 56.25 44.15 37.47 36.35 30.71 19.45 13.64 15.36 11.74 4.56 1.65
PG ($\delta_{\text{PG}}=10$) 61.71 56.32 44.30 37.55 36.40 30.79 19.52 13.73 12.98 10.13 4.38 1.58
PG ($\delta_{\text{PG}}=100$) 61.71 56.32 44.30 37.55 36.40 30.79 19.52 13.73 13.60 10.67 4.22 1.52
MM 61.71 56.32 44.30 37.55 36.40 30.79 19.52 13.73 16.59 12.09 4.66 1.72
$N=200$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 1069 1076 1078 1112 1001 1069 1058 1123 1200 948 1104 1015
PG ($\delta_{\text{PG}}=10$) 1183 1185 1185 1206 1137 1200 1200 1200 1200 1202 1200 1200
PG ($\delta_{\text{PG}}=100$) 1200 1200 1200 1200 1200 1200 1200 1200 1200 1200 1200 1200
MM 238 254 254 242 302 301 286 293 704 640 563 502
Num. Iterations PG ($\delta_{\text{PG}}=1$) 22.3 22.3 22.3 22.3 20.7 22.0 21.3 21.3 24.0 19.3 22.7 20.3
PG ($\delta_{\text{PG}}=10$) 25.3 25.3 25.0 24.3 23.0 25.3 24.3 24.7 24.3 24.7 24.7 23.3
PG ($\delta_{\text{PG}}=100$) 24.7 24.7 24.7 25.0 25.0 25.0 24.7 24.7 24.7 24.0 24.3 24.7
MM 5.0 5.3 5.3 5.0 6.3 6.3 6.0 6.0 15.0 13.3 11.7 10.3
Objective Function PG ($\delta_{\text{PG}}=1$) 59.78 54.82 43.69 37.29 34.39 29.31 18.78 13.47 14.59 11.08 4.35 1.52
PG ($\delta_{\text{PG}}=10$) 59.97 54.88 43.75 37.35 34.36 29.34 18.94 13.45 13.80 10.64 3.97 1.53
PG ($\delta_{\text{PG}}=100$) 55.46 50.38 31.77 29.41 32.36 21.50 6.88 -5.98 1.35 -2.16 -2.99 -3.02
MM 59.97 54.89 43.77 37.37 34.48 29.37 19.06 13.54 15.23 11.34 4.48 1.61
Table 3 presents the results of the first set of experiments, where we assume $\theta_{1}=\cdots=\theta_{N}=\theta\in\{1/5,1/3,1,2\}$ per minute. Accordingly, the mean sojourn times for riders are $\{5,3,1,0.5\}$ minutes, respectively. The results demonstrate that the MM algorithm substantially reduces both the running time and the number of iterations compared to the PG method. Compared to PG with three different step sizes ($\delta_{\text{PG}}=1,10,100$), the MM algorithm reduces average running time by 72.6%, 71.6%, and 73.8%, respectively, and reduces the average number of iterations by 77.5%, 72.9%, and 75.2%. Since the running time is capped at 20 minutes, the actual time savings may be even greater. Moreover, within the maximum allowed running time, the MM algorithm consistently achieves a higher objective value—on average, 1.05%, 2.05%, and 14.27% higher than the PG methods. We also observe that the performance of PG is sensitive to step size selection: although a step size of 1 yields the best average performance, as shown in Table 4, it is not optimal in every instance. For example, a step size of 10 performs better when $c=0.7$ or $c=0.9$. In contrast, MM demonstrates strong robustness and outperforms PG across all cases and step-size rules.
Table 4: Averaged objective values under different values of $c$, aggregated from Table 3.
$c$ (\$/mile) 0.7 0.9 1.1
PG ($\delta_{\text{PG}}=1$) 50.14 25.31 8.41
PG ($\delta_{\text{PG}}=10$) 50.28 25.42 7.36
PG ($\delta_{\text{PG}}=100$) 47.87 21.97 4.33
MM 50.28 25.45 9.02
It can also be observed that as the number of types $N$ increases, the running time of both GA and MM increases, primarily because solving $c(\boldsymbol{\lambda})$ involves $N+N^{2}$ variables at each iteration. However, the number of iterations required by the MM algorithm does not scale proportionally with $N$, and in many cases remains nearly constant at a single-digit value. Even when $N=200$, the MM algorithm remains efficient and consistently produces results within a short CPU time.
Table 5: Numerical results of solving (1) when $\theta_{i}$, $i\in[N]$ is not equal.
$c$ (\$/mile) $0.7$ $0.9$ $1.1$
$(\underline{\theta},\overline{\theta})$ (min-1) $(1/5,1/3)$ $(1,2)$ $(1/5,1/3)$ $(1,2)$ $(1/5,1/3)$ $(1,2)$
$N=100$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 226 302 226 243 327 313
PG ($\delta_{\text{PG}}=10$) 352 358 321 314 737 406
PG ($\delta_{\text{PG}}=100$) 357 358 367 367 760 682
MM 60 59 72 84 152 134
Num. Iterations PG ($\delta_{\text{PG}}=1$) 19.0 24.7 19.0 20.3 26.7 26.3
PG ($\delta_{\text{PG}}=10$) 29.3 29.7 27.0 26.3 59.0 34.3
PG ($\delta_{\text{PG}}=100$) 30.0 30.3 30.7 30.7 61.0 56.7
MM 5.0 5.0 6.0 7.0 12.7 11.0
Objective Function PG ($\delta_{\text{PG}}=1$) 61.72 41.75 36.82 17.16 15.29 2.99
PG ($\delta_{\text{PG}}=10$) 61.73 41.76 36.82 17.17 12.92 2.83
PG ($\delta_{\text{PG}}=100$) 61.75 41.76 36.82 17.17 11.39 2.79
MM 61.75 41.76 36.82 17.17 17.00 3.04
$N=150$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 981 576 594 758 1050 831
PG ($\delta_{\text{PG}}=10$) 734 734 742 757 1020 902
PG ($\delta_{\text{PG}}=100$) 921 898 841 897 1110 1102
MM 134 144 161 161 403 425
Num. Iterations PG ($\delta_{\text{PG}}=1$) 37.0 21.7 22.7 28.0 39.3 30.3
PG ($\delta_{\text{PG}}=10$) 27.3 27.7 27.3 27.7 38.0 32.7
PG ($\delta_{\text{PG}}=100$) 34.3 33.3 31.0 33.3 41.3 40.7
MM 5.0 5.3 6.0 6.0 14.3 15.0
Objective Function PG ($\delta_{\text{PG}}=1$) 58.71 40.30 33.25 16.12 13.23 2.72
PG ($\delta_{\text{PG}}=10$) 58.85 40.47 33.36 16.22 11.81 2.67
PG ($\delta_{\text{PG}}=100$) 58.85 40.47 33.36 16.22 11.79 2.65
MM 58.85 40.47 33.36 16.22 14.04 2.92
$N=200$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 1042 1086 1099 1100 1188 1200
PG ($\delta_{\text{PG}}=10$) 1210 1205 1200 1200 1200 1200
PG ($\delta_{\text{PG}}=100$) 1200 1200 1200 1200 1200 1200
MM 238 240 305 302 875 826
Num. Iterations PG ($\delta_{\text{PG}}=1$) 22.0 22.7 23.0 22.7 24.3 25.0
PG ($\delta_{\text{PG}}=10$) 25.7 25.3 25.0 25.7 25.0 25.0
PG ($\delta_{\text{PG}}=100$) 25.7 25.3 25.0 26.0 25.0 25.0
MM 5.0 5.0 6.3 6.0 18.3 17.0
Objective Function PG ($\delta_{\text{PG}}=1$) 56.90 39.78 31.31 15.48 12.15 2.25
PG ($\delta_{\text{PG}}=10$) 57.28 40.02 31.71 15.90 12.07 2.60
PG ($\delta_{\text{PG}}=100$) 53.44 33.93 26.24 14.68 5.87 -3.64
MM 57.28 40.14 31.73 15.91 13.05 2.78
Finally, Table 5 presents the experimental results for the setting where $\theta_{i}$ varies across agent types. For type-$i$ agents, $\theta_{i}$ is randomly drawn from a uniform distribution $U[\underline{\theta},\overline{\theta}]$, with two parameter choices: $(\underline{\theta},\overline{\theta})\in\{(1/5,1/3),(1,2)\}$. The results exhibit a similar trend to that observed in Table 1. Compared to three PG variants with $\delta_{\text{PG}}=1,10,100$, the MM algorithm reduces running time by 63.7%, 67.3% and 69.9%; decreases the number of iterations by 65.7%, 71.0% and 74.2%; and improves the objective value by 1.27%, 1.63% and 8.32% within the 20-minute maximum runtime, respectively. The performance of GA is also sensitive to the step size: $\delta_{\text{GA}}=10$ performs better when $c\leq 0.9$, whereas $\delta_{\text{GA}}=1$ is preferable when $c=1.1$.
5 Conclusion
Motivated by growing pricing applications in centralized matching markets, we study the behavior of the value function of the linear program (LP) introduced by Aouad and Sarıtaç (2022) for cost-minimizing dynamic stochastic matching, with a particular focus on its (weak) concavity with respect to agents’ arrival rates. We establish a key result: it is sufficient for all optimal solutions of the LP to have strictly positive unmatched rates to ensure weak concavity. Leveraging this insight, we demonstrate that weak concavity is guaranteed when there are at most three agent types sharing the same patience level. For the general case with an arbitrary number of types, (weak) concavity is more likely to hold when agents are either highly patient, have similar patience levels, or exhibit substantial heterogeneity in matching efficiencies.
We then turn our attention to the pricing application and design an efficient MM algorithm that requires little stepsize tuning. To evaluate its performance, we apply our MM algorithm to the Chicago ridesharing dataset. Experimental results demonstrate that, compared to a projected gradient-based method whose performance is highly sensitive to stepsize selection, our approach significantly reduces both running time and the number of iterations, while simultaneously achieving better optimality.
[…]
The authors thank Ying Cui from UC Berkeley for helpful discussions.
{APPENDICES}
6 Proofs
6.1 Proof of Proposition 2.2
When $\theta_{i}=0$, $\forall i\in[N]$, constraint (2c) is relaxed, and we have $y_{i}^{*}=0$, $\forall i\in[N]$—otherwise, by selecting a small enough value of $\varepsilon>0$, we can always decrease $y_{i}^{*}$ by $2\varepsilon$ and increase $x_{i,i}^{*}$ by $\varepsilon$, reducing the total cost by $c_{(i)}\varepsilon$ while still satisfying the constraint (2b).
Similarly, we can prove that there exists an optimal solution satisfying $x_{i,j}^{*}=0$, $\forall i\neq j$. Again if the opposite is true, we can decrease $x_{i,j}^{*}$ by $\varepsilon$, increase $x_{i,i}^{*}$ and $x_{j,j}^{*}$ by $\varepsilon/2$, then we can obtain a new feasible solution with the total cost smaller than or equal to the original total cost, since
$c_{(i,j)}\geq\frac{1}{2}c_{(i)}+\frac{1}{2}c_{(j)},\ \forall i,j\in[N].$
Therefore, an optimal solution can be given by $x_{i,i}^{*}=\lambda_{i}$, $\forall i\in[N]$. The optimal cost is given by $\sum_{i\in[N]}c_{(i)}\lambda_{i}/2$ and is thus linear. $\square$
6.2 Proof of Proposition 2.3
Denote the corresponding optimal solution of $c(\boldsymbol{\lambda})$ as $\boldsymbol{x}^{*}=(x_{i,j}^{*})_{i,j\in[N]}$ and $\boldsymbol{y}^{*}=(y_{i}^{*})_{i\in[N]}$. By multiplying constraints (2b) and (2c) with $\alpha$, we have
$\sum_{j\in[N]}(\alpha x_{j,i}^{*})+\sum_{j\in[N]}(\alpha x_{i,j}^{*})+(\alpha y_{i}^{*})$ $=(\alpha\lambda_{i}),$ $\forall i\in[N],$
$\theta_{i}(\alpha x_{i,j}^{*})$ $\leq(\alpha\lambda_{j})y_{i}^{*},$ $\forall i,j\in[N],$
since $\alpha>1$, $(\alpha\lambda_{j})y_{i}^{*}<(\alpha\lambda_{j})(\alpha y_{i}^{*})$, and thus $\alpha\boldsymbol{x}^{*}$ and $\alpha\boldsymbol{y}^{*}$ are feasible solution when $\alpha\boldsymbol{\lambda}$ is the arrival rate. Then we have
$\alpha c(\boldsymbol{\lambda})=\sum_{i\in[N]}\sum_{j\in[N]}c_{(i,j)}\left(\alpha x^{*}_{i,j}\right)+\sum_{i\in[N]}c_{(i)}\left(\alpha y^{*}_{i}\right)\geq c(\alpha\boldsymbol{\lambda}),$
and thus $c(\alpha\boldsymbol{\lambda})/\|\alpha\boldsymbol{\lambda}\|\leq c(\boldsymbol{\lambda})/\|\boldsymbol{\lambda}\|$. $\square$
6.3 Proof of Lemma 3.1
When $N=1$, constraint (2b) can be written as $2x_{1,1}+y_{1}=\lambda_{1}$, and constraint (2c) can be written as $\theta_{1}x_{1,1}\leq\lambda_{1}y_{1}$. We first state that constraint (2c) should be binding. Otherwise, if the optimal solution satisfies $\theta_{1}x_{1,1}^{*}<\lambda_{1}y_{1}^{*}$, we can obtain a few feasible solution by letting $\hat{x}_{1,1}=x_{1,1}^{*}+\varepsilon,\ \hat{y}_{1}=y_{1}^{*}-2\varepsilon.$ By selecting a small enough value of $\varepsilon>0$, $\theta_{1}\hat{x}_{1,1}<\lambda_{1}\hat{y}_{1}$ is still satisfied. Constraint (2b) also holds. Under this new feasible solution, the value function is decreased by $c_{(1)}\varepsilon$ , which is a contradiction.
Therefore, constraint (2c) is indeed binding. Solving $x_{1,1}$ and $y_{1}$ subject to two equations (2b) and (2c) leads to $y_{1}^{*}=\lambda_{1}\theta_{1}/(\theta_{1}+2\lambda_{1})$ and $x_{1,1}^{*}=\lambda_{1}^{2}/(\theta_{1}+2\lambda_{1})$, and the value function is
$c(\lambda_{1})=c_{(1)}\frac{\lambda_{1}(\theta_{1}+\lambda_{1})}{\theta_{1}+2\lambda_{1}}.$
Taking the second-order derivative of $c(\lambda_{1})$ leads to
$\frac{d^{2}c(\lambda_{1})}{d\lambda_{1}^{2}}=-2c_{(1)}\frac{\theta_{1}^{2}}{(\theta_{1}+2\lambda_{1})^{3}}<0,$
thus $c(\lambda_{1})$ is concave. $\square$
6.4 Proof of Lemma 3.3
The proof relies on the following three lemmas. The first lemma below establishes the connection between local weak concavity and global weak concavity. For ease of presentation, define $B(\hat{\boldsymbol{\lambda}},\varepsilon)$ as $\{\boldsymbol{\lambda}\ |\ \|\boldsymbol{\lambda}-\hat{\boldsymbol{\lambda}}\|\leq\varepsilon\}$.
Lemma 6.1
If given any $\hat{\boldsymbol{\lambda}}\in\Lambda$, there always exists $\varepsilon>0$ such that $c(\boldsymbol{\lambda})$ is weakly concave for $\boldsymbol{\lambda}\in B(\hat{\boldsymbol{\lambda}},\varepsilon)\cap\Lambda$, then $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$.
Proof 6.2
Proof. Given a value of $\hat{\boldsymbol{\lambda}}\in\Lambda$, since $c(\boldsymbol{\lambda})$ is weakly concave in $B(\hat{\boldsymbol{\lambda}},\varepsilon)\cap\Lambda$, there exist a finite value of $\rho$, denoted as $\rho(\hat{\boldsymbol{\lambda}})$, such that $c(\boldsymbol{\lambda})-\rho(\hat{\boldsymbol{\lambda}})\|\boldsymbol{\lambda}\|^{2}/2$ is concave in $B(\hat{\boldsymbol{\lambda}},\varepsilon)\cap\Lambda$. Denote
$\bar{\rho}:=\sup_{\hat{\boldsymbol{\lambda}}\in\Lambda}\rho(\hat{\boldsymbol{\lambda}})<+\infty,$
which is also a finite value because $\Lambda$ is a closed interval. Since $\bar{\rho}\geq\rho(\hat{\boldsymbol{\lambda}})$, $c(\boldsymbol{\lambda})-\bar{\rho}\|\boldsymbol{\lambda}\|^{2}/2$ is also concave in $B(\hat{\boldsymbol{\lambda}},\varepsilon)\cap\Lambda$.
The concavity of a function is equivalent to the convexity of the function’s hypograph (see e.g., Section 3.1.7 in Boyd and Vandenberghe 2004). Since $c(\boldsymbol{\lambda})-\bar{\rho}\|\boldsymbol{\lambda}\|^{2}/2$ is concave in $B(\hat{\boldsymbol{\lambda}},\varepsilon)\cap\Lambda$, the following hypograph is a convex set:
$\left\{(\boldsymbol{\lambda},t)\ \bigg{|}\ \boldsymbol{\lambda}\in B(\hat{\boldsymbol{\lambda}},\varepsilon)\cap\Lambda,t\leq c(\boldsymbol{\lambda})-\frac{\bar{\rho}\|\boldsymbol{\lambda}\|^{2}}{2}\right\}.$
According to the Tietze-Nakajima theorem, a closed, connected, and locally convex set is convex (Tietze 1928, Nakajima 1928, Bjorndahl and Karshon 2010). This implies that the following set is convex:
$\left\{(\boldsymbol{\lambda},t)\ \bigg{|}\ \boldsymbol{\lambda}\in\Lambda,t\leq c(\boldsymbol{\lambda})-\frac{\bar{\rho}\|\boldsymbol{\lambda}\|^{2}}{2}\right\},$
where this set is closed and connected because $c(\boldsymbol{\lambda})$ is bounded (since all $\boldsymbol{x}$ and $\boldsymbol{y}$ is bounded by $\overline{\boldsymbol{\lambda}}$ in (2b)) and continuous (according to the maximum theorem, Ok 2011). Again using the equivalence between a concave function and its hypograph, we know that the function $c(\boldsymbol{\lambda})-\bar{\rho}\|\boldsymbol{\lambda}\|^{2}/2$ is concave in $\Lambda$, and thus $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$. $\square$
The second lemma below is straightforward. It gives the boundedness property of a function if it can be expressed as a ratio of two polynomial functions and the denominator is always positive.
Lemma 6.3
Assume $\boldsymbol{\lambda}\in\hat{\Lambda}\subset R^{N}$, where $\hat{\Lambda}$ is any compact and convex set. Let $p(\boldsymbol{\lambda})$ and $q(\boldsymbol{\lambda})$ be two (multivariate) polynomial functions of $\boldsymbol{\lambda}\in\hat{\Lambda}$. If $q(\boldsymbol{\lambda})\geq q_{\min}>0$ always holds, then $p(\boldsymbol{\lambda})/q(\boldsymbol{\lambda})$, as well as its any order derivatives, are bounded in $\hat{\Lambda}$.
Proof 6.4
Proof. Since $p(\boldsymbol{\lambda})$ and $q(\boldsymbol{\lambda})$ are two polynomial functions defined in a compact set $\hat{\Lambda}$, they are both bounded. Denote their lower and upper bounds as $p_{\min},p_{\max},q_{\min},q_{\max}$, respectively. Since $q_{\max}\geq q_{\min}>0$, the lower bound of $p(\boldsymbol{\lambda})/q(\boldsymbol{\lambda})$ can be easily given as the minimum value among $p_{\min}/q_{\min},p_{\min}/q_{\max},p_{\max}/q_{\min},p_{\max}/q_{\max}$, and the upper bound can be given as the maximum value among these four.
Now consider the first-order partial derivative of $p(\boldsymbol{\lambda})/q(\boldsymbol{\lambda})$:
$\frac{\partial p(\boldsymbol{\lambda})/q(\boldsymbol{\lambda})}{\partial\lambda_{i}}=\frac{\frac{\partial p(\boldsymbol{\lambda})}{\partial\lambda_{i}}q(\boldsymbol{\lambda})-\frac{\partial q(\boldsymbol{\lambda})}{\partial\lambda_{i}}p(\boldsymbol{\lambda})}{[q(\boldsymbol{\lambda})]^{2}}.$
It can be seen that the numerator and denominator are still polynomial functions, and the denominator still has a positive lower bound of $q_{\min}^{2}$. Therefore, the aforementioned property still holds, and thus the first-order partial derivative is still bounded. Similarly, we can prove that any-order derivatives exist and are all bounded. $\square$
Lastly, the third lemma gives the weakly concave property for a twice differentiable function with a bounded Hessian.
Lemma 6.5
Assume $\boldsymbol{\lambda}\in\hat{\Lambda}\subset\mathbb{R}^{N}$, where $\hat{\Lambda}$ is any compact and convex set. If a function $f(\boldsymbol{\lambda})$ is twice differentiable and each element in its Hessian matrix is bounded for $\boldsymbol{\lambda}\in\hat{\Lambda}$, then $f(\boldsymbol{\lambda})$ is weakly concave in $\hat{\Lambda}$.
Proof 6.6
Proof. Let $H(f(\boldsymbol{\lambda}))$ denote the Hessian of $f(\boldsymbol{\lambda})$, and its elements are denoted as $h_{i,j}(\boldsymbol{\lambda})$, $i,j\in[N]$. Since each element is bounded, we let $\bar{h}_{i,j}$ denote the upper bound of the absolute value of $h_{i,j}(\boldsymbol{\lambda})$, that is, we have
$|h_{i,j}(\boldsymbol{\lambda})|\leq\bar{h}_{i,j},\forall\boldsymbol{\lambda}\in\hat{\Lambda},\ i,j\in[N].$
Given that $f(\boldsymbol{\lambda})$ is twice differentiable, so is $f(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}$ for any $\rho\geq 0$. Then the Hessian of $f(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2$ can be expressed as $H(f(\boldsymbol{\lambda}))-\rho\boldsymbol{I}$, where $\boldsymbol{I}$ is the identity matrix. The diagonal elements of this Hessian matrix is $h_{i,i}(\boldsymbol{\lambda})-\rho,\ i\in[N]$, and the non-diagonal elements are still $h_{i,j}(\boldsymbol{\lambda}),\ i,j\in[N]$. We select a value of $\rho$ that satisfies
$\rho>\max_{i\in[N]}\sum_{j\in[N]}\bar{h}_{i,j},$
then $h_{i,i}(\boldsymbol{\lambda})-\rho\leq\bar{h}_{i,i}-\rho\leq 0,\ i\in[N]$, and
$|h_{i,i}(\boldsymbol{\lambda})-\rho|-\sum_{j\neq i}|h_{i,j}(\boldsymbol{\lambda})|$ $=\rho-h_{i,i}(\boldsymbol{\lambda})-\sum_{j\neq i}|h_{i,j}(\boldsymbol{\lambda})|$
$\geq\rho-\sum_{j\in[N]}|h_{i,j}(\boldsymbol{\lambda})|$
$\geq\rho-\sum_{j\in[N]}\bar{h}_{i,j}$
$>0,\ \forall i\in[N].$
This means that the matrix $H(f(\boldsymbol{\lambda}))-\rho\boldsymbol{I}$ is strictly diagonally dominant and its diagonal elements are all non-positive. The matrix is thus negative semidefinite (see, e.g., Theorem 6.1.10 in Horn and Johnson 2012).
For the twice differentiable function $f(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}$, since its Hessian $H(f(\boldsymbol{\lambda}))-\rho\boldsymbol{I}$ is negative semidefinite, it is concave. Thus $f(\boldsymbol{\lambda})$ is weakly concave in $\hat{\Lambda}$. $\square$
Now we are ready to give the proof of Lemma 3.3.
Proof 6.7
Proof of Lemma 3.3.
We first write program (2) as the standard form consisting of only equality constraints and nonnegative decision variables:
$c(\boldsymbol{\lambda})=\min_{\boldsymbol{x},\boldsymbol{y}}\quad$ $\sum_{i\in[N]}\sum_{j\in[N]}c_{(i,j)}x_{i,j}+\sum_{i\in[N]}c_{(i)}y_{i},$ (5a)
s.t. $\theta_{i}x_{i,j}+\delta_{i,j}=\lambda_{j}y_{i},$ $\forall i,j\in[N],$ (5b)
$\delta_{i,j}\geq 0,$ $\forall i,j\in[N],$ (5c)
[…]
For this standard form, it involves $2N^{2}+N$ variables and $N^{2}+N$ equality constraints. It can be verified that the equality constraints are linear independent as we assumed that $\boldsymbol{\lambda}\geq\underline{\boldsymbol{\lambda}}>\boldsymbol{0}$: Constraint (2b) spans $N$ dimensions because the unique occurrence of $y_{i}$, $\forall i\in[N]$, and (5c) spans additional $N^{2}$ dimensions because $\delta_{i,j}$, $\forall i,j\in[N]$.
Given a value of $\hat{\boldsymbol{\lambda}}\in\Lambda$, assume there are in total $K\geq 1$ optimal basic feasible solutions (BFS) to this program when $\boldsymbol{\lambda}=\hat{\boldsymbol{\lambda}}$. Based on the definition of BFS and the linear independence of constraints, there must be at least $(2N^{2}+N)-(N^{2}+N)=N^{2}$ variables equal to zero in each BFS (Bertsimas and Tsitsiklis 1997). On the other hand, as we assumed, $y_{i}^{*}>0,\forall i\in[N]$ at optimality, and for any $i,j\in[N]$, $x_{i,j}^{*}$ and $\delta_{i,j}^{*}$ cannot be zero at the same time (otherwise constraint (5b) is violated), meaning that there are at most $N^{2}$ zero variables for each optimal BFS. Therefore, any optimal BFS to this program has exactly $N^{2}$ zero variables (thus is non-degenerate), and either $x_{i,j}^{*}=0$ or $\delta_{i,j}^{*}=0$ (and thus $x_{i,j}^{*}=\hat{\lambda}_{j}y_{i}^{*}/\theta_{i}$), $\forall i,j\in[N]$. Below, for the $k$-th optimal BFS at $\hat{\boldsymbol{\lambda}}$, we let $X_{i}^{(k)}$ denote $\{j|\delta_{i,j}^{*}=0\}$.
We can further derive the form of each optimal BFS and its corresponding value function at $\hat{\boldsymbol{\lambda}}$. For the $k$-th optimal BFS, substituting $x_{i,j}$ with either 0 or $\hat{\lambda}_{j}y_{i}/\theta_{i}$ in (2b), we obtain the following set of equations:
$\left(1+\frac{\sum_{j\in X_{i}^{(k)}}\hat{\lambda}_{j}}{\theta_{i}}\right)y_{i}+\sum_{j\in[N],i\in X_{j}^{k}}\frac{\hat{\lambda}_{i}}{\theta_{j}}y_{j}=\hat{\lambda}_{i},\ \forall i\in[N],\ k\in[K].$ (6)
For this set of equations, denote the coefficient matrix as $\boldsymbol{A}^{(k)}(\hat{\boldsymbol{\lambda}})=(a_{i,j})_{i,j\in[N]}$, then we have
$a_{i,i}$ $=1+\frac{\sum_{j\in X_{i}^{(k)}}\hat{\lambda}_{j}}{\theta_{i}},$ $\forall i\in[N],\ k\in[K],$
$a_{j,i}$ $=\frac{\hat{\lambda}_{j}}{\theta_{i}},$ $\forall i,j\in[N],\ j\in X_{i}^{(k)},\ j\neq i,\ k\in[K],$
$a_{j,i}$ $=0,$ $\forall i,j\in[N],\ j\notin X_{i}^{(k)},\ j\neq i,\ k\in[K].$
It can be verified that $\boldsymbol{A}^{(k)}(\hat{\boldsymbol{\lambda}})$ is strictly diagonally-dominant by column, since
$|a_{i,i}|-\sum_{j\neq i}|a_{j,i}|=a_{i,i}-\sum_{j\neq i}a_{j,i}=a_{i,i}-\sum_{j\in X_{i}^{(k)},j\neq i}a_{j,i}\geq 1,\ \forall i\in[N],\ k\in[K].$
This also gives a lower bound of $\text{det}(\boldsymbol{A}^{(k)}(\hat{\boldsymbol{\lambda}}))$, the determinant of $\boldsymbol{A}^{(k)}(\hat{\boldsymbol{\lambda}})$, as 1 (Ostrowski 1952). Therefore, $\boldsymbol{A}^{(k)}(\hat{\boldsymbol{\lambda}})$ is invertible, and the solution to equation (6), denoted as $\boldsymbol{y}^{(k)}(\hat{\boldsymbol{\lambda}})$, is given by:
$\boldsymbol{y}^{(k)}(\hat{\boldsymbol{\lambda}}):=\boldsymbol{A}^{(k)}(\hat{\boldsymbol{\lambda}})^{-1}\hat{\boldsymbol{\lambda}}=\frac{\text{adj}(\boldsymbol{A}^{(k)}(\hat{\boldsymbol{\lambda}}))}{\text{det}(\boldsymbol{A}^{(k)}(\hat{\boldsymbol{\lambda}}))}\hat{\boldsymbol{\lambda}}>\boldsymbol{0},\ \forall k\in[K],$ (7)
where $\text{adj}(\boldsymbol{A}^{(k)}(\hat{\boldsymbol{\lambda}}))$ is the adjugate matrix of $\boldsymbol{A}^{(k)}(\hat{\boldsymbol{\lambda}})$. The value function produced by this BFS, denoted as $c_{\text{BFS}}^{(k)}(\hat{\boldsymbol{\lambda}})$, is given by
$c_{\text{BFS}}^{(k)}(\hat{\boldsymbol{\lambda}}):=\sum_{i\in[N]}\left(c_{(i)}+\frac{\sum_{j\in X_{i}^{(k)}}c_{(i,j)}\hat{\lambda}_{j}}{\theta_{i}}\right)y_{i}^{(k)}(\hat{\boldsymbol{\lambda}}),\ \forall k\in[K].$ (8)
Now consider a small neighborhood of $\hat{\boldsymbol{\lambda}}$, defined as $B(\hat{\boldsymbol{\lambda}},\varepsilon)\cap\Lambda$, where $\varepsilon>0$ is a sufficiently small value. For every optimal BFS at $\hat{\boldsymbol{\lambda}}$, we can use it to produce another solution at any $\boldsymbol{\lambda}\in B(\hat{\boldsymbol{\lambda}},\varepsilon)\cap\Lambda$, by fixing the same non-basic variables to zero and solving basic variables as a function of $\boldsymbol{\lambda}$. Specifically, for the $k$-th optimal BFS at $\hat{\boldsymbol{\lambda}}$, by substituting $\hat{\boldsymbol{\lambda}}$ with $\boldsymbol{\lambda}$ in (7), we obtain the following solution at $\boldsymbol{\lambda}\in B(\hat{\boldsymbol{\lambda}},\varepsilon)\cap\Lambda$:
$\boldsymbol{y}^{(k)}(\boldsymbol{\lambda})=\boldsymbol{A}^{(k)}(\boldsymbol{\lambda})^{-1}\boldsymbol{\lambda}=\frac{\text{adj}(\boldsymbol{A}^{(k)}(\boldsymbol{\lambda}))}{\text{det}(\boldsymbol{A}^{(k)}(\boldsymbol{\lambda}))}\boldsymbol{\lambda},\ \forall k\in[K],$
and the values of $x_{i,j}$ and $\delta_{i,j}$ are obtained as:
$x_{i,j}^{(k)}(\boldsymbol{\lambda})$ $=\frac{\lambda_{j}y_{i}^{(k)}(\boldsymbol{\lambda})}{\theta_{i}},\ \delta_{i,j}^{(k)}(\boldsymbol{\lambda})=0,$ $\forall i,j\in[N],\ j\in X_{i}^{(k)},\ k\in[K],$
$x_{i,j}^{(k)}(\boldsymbol{\lambda})$ $=0,\ \delta_{i,j}^{(k)}(\boldsymbol{\lambda})=\frac{\lambda_{j}y_{i}^{(k)}(\boldsymbol{\lambda})}{\theta_{i}},$ $\forall i,j\in[N],\ j\notin X_{i}^{(k)},\ k\in[K].$
The value function $c_{\text{BFS}}^{(k)}(\boldsymbol{\lambda})$ is similarly obtained by substituting $\hat{\boldsymbol{\lambda}}$ with $\boldsymbol{\lambda}$ in (8).
For any $k\in[K]$, since we assume $\boldsymbol{y}^{(k)}(\hat{\boldsymbol{\lambda}})>\boldsymbol{0}$, by selecting a small enough value of $\varepsilon$, we can guarantee that $\boldsymbol{y}^{(k)}(\boldsymbol{\lambda})>\boldsymbol{0}$ still holds for any $\boldsymbol{\lambda}\in B(\hat{\boldsymbol{\lambda}},\varepsilon)\cap\Lambda$, and thus $x_{i,j}^{(k)}(\boldsymbol{\lambda}),\delta_{i,j}^{(k)}(\boldsymbol{\lambda})\geq 0$ also holds. Therefore, these $K$ solutions are BFS at $\hat{\boldsymbol{\lambda}}$. In addition, based on the continuity of $c_{\text{BFS}}^{(k)}(\boldsymbol{\lambda})$ and the fact that the number of BFS is finite for any LP, we can also guarantee, by selecting a small enough value of $\varepsilon$, that the best solution among the $K$ solutions produces the optimal solution at $\boldsymbol{\lambda}$. That is, we have the following results:
$\forall\hat{\boldsymbol{\lambda}}\in\Lambda,\ \exists\varepsilon>0,\text{ s.t. }c(\boldsymbol{\lambda})=\min_{k\in[K]}\{c_{\text{BFS}}^{(k)}(\boldsymbol{\lambda})\},\ \forall\boldsymbol{\lambda}\in B(\hat{\boldsymbol{\lambda}},\varepsilon)\cap\Lambda.$
Note that all elements in $\boldsymbol{A}^{(k)}(\boldsymbol{\lambda})$ are polynomial in $\boldsymbol{\lambda}$, and the same is also true for $\text{adj}(\boldsymbol{A}^{(k)}(\boldsymbol{\lambda})))$ and $\text{det}(\boldsymbol{A}^{(k)}(\boldsymbol{\lambda})))$. Therefore, $y_{i}^{(k)}(\boldsymbol{\lambda})$ can be expressed as a ratio of two polynomial functions of $\boldsymbol{\lambda}$, where its denominator $\text{det}(\boldsymbol{A}^{(k)}(\boldsymbol{\lambda}))$ is always positive and greater than 1. Based on Lemma 6.3, $y_{i}^{(k)}(\boldsymbol{\lambda})$ has infinitely many derivatives and is bounded as $\boldsymbol{\lambda}$ changes. The same results of differentiability and boundedness also hold for $c_{\text{BFS}}^{(k)}(\boldsymbol{\lambda})$.
Denote $\boldsymbol{H}(c_{\text{BFS}}^{(k)}(\boldsymbol{\lambda}))$ as the Hessian matrix of the value function $c_{\text{BFS}}^{(k)}(\boldsymbol{\lambda})$. Based on Lemma 6.3, every element of $\boldsymbol{H}(c_{\text{BFS}}^{(k)}(\boldsymbol{\lambda}))$ is still bounded as $\boldsymbol{\lambda}$ changes. Then, based on Lemma 6.5, $c_{\text{BFS}}^{(k)}(\boldsymbol{\lambda})$ is weakly concave: There exists a positive value of $\rho$, denoted as $\rho^{(k)}(\boldsymbol{\lambda})$, such that $c_{\text{BFS}}^{(k)}(\boldsymbol{\lambda})-\rho^{(k)}(\boldsymbol{\lambda})\|\boldsymbol{\lambda}\|^{2}/2$ is concave in $B(\hat{\boldsymbol{\lambda}},\varepsilon)\cap\Lambda$. Let $\bar{\rho}(\boldsymbol{\lambda})=\max_{k\in[K]}\rho^{(k)}(\boldsymbol{\lambda})$, then $c_{\text{BFS}}^{(k)}(\boldsymbol{\lambda})-\bar{\rho}(\boldsymbol{\lambda})\|\boldsymbol{\lambda}\|^{2}/2$ is still concave since $\bar{\rho}(\boldsymbol{\lambda})>\rho^{(k)}(\boldsymbol{\lambda})$. Given that $c(\boldsymbol{\lambda})=\min_{k\in[K]}\{c_{\text{BFS}}^{(k)}(\boldsymbol{\lambda})\}$, we know that $c(\boldsymbol{\lambda})-\bar{\rho}(\boldsymbol{\lambda})\|\boldsymbol{\lambda}\|^{2}/2$ because the minimization preserves concavity (see Section 3.2.3, Boyd and Vandenberghe 2004), and thus $c(\boldsymbol{\lambda})$ is (locally) weakly concave for $\boldsymbol{\lambda}\in B(\hat{\boldsymbol{\lambda}},\varepsilon)\cap\Lambda$. Lastly, using Lemma 6.1, we prove that $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$. $\square$
6.5 Proof of Lemma 3.4
For ease of presentation, we explicitly write all constraints in the case of $N=2$ as follows.
$c(\boldsymbol{\lambda})=\min_{\boldsymbol{x},\boldsymbol{y}}\quad$ $\ c_{(1)}(x_{1,1}+y_{1})+c_{(2)}(x_{2,2}+y_{2})+c_{(1,2)}(x_{1,2}+x_{2,1}),$ (9a)
s.t. $2x_{1,1}+x_{1,2}+x_{2,1}+y_{1}=\lambda_{1},$ (9b)
$2x_{2,2}+x_{1,2}+x_{2,1}+y_{2}=\lambda_{2},$ (9c)
$\theta_{1}x_{1,1}\leq\lambda_{1}y_{1},$ (9d)
$\theta_{1}x_{1,2}\leq\lambda_{2}y_{1},$ (9e)
$\theta_{2}x_{2,1}\leq\lambda_{1}y_{2},$ (9f)
$\theta_{2}x_{2,2}\leq\lambda_{2}y_{2},$ (9g)
$x_{1,1},x_{1,2},x_{2,1},x_{2,2}\geq 0,$ (9h)
$y_{1},y_{2}\geq 0.$ (9i)
The proof of Lemma 3.4 relies on the following lemmas that partially characterize the optimal solution structure.
Lemma 6.8
$y_{1}^{*}=\max\{\theta_{1}x_{1,1}^{*}/\lambda_{1},\theta_{1}x_{1,2}^{*}/\lambda_{2}\}$, and $y_{2}^{*}=\max\{\theta_{2}x_{2,1}^{*}/\lambda_{1},\theta_{2}x_{2,2}^{*}/\lambda_{2}\}$. That is, at least one of (9d) and (9e) is binding, and at least one of (9f) and (9g) is binding.
Proof 6.9
Proof. Assume by way of contradiction that the optimal solution satisfies $y_{1}^{*}\neq\max\{\theta_{1}x_{1,1}^{*}/\lambda_{1},\theta_{1}x_{1,2}^{*}/\lambda_{2}\}$. According to (9d) and (9e), $y_{1}^{*}>\theta_{1}x_{1,1}^{*}/\lambda_{1}$ and $y_{1}^{*}>\theta_{1}x_{1,2}^{*}/\lambda_{2}$. In this case, we can increase the value of $x_{1,1}^{*}$ by $\varepsilon>0$, decrease the value of $y_{1}^{*}$ by $2\varepsilon$, and keep the remaining variables unchanged. This still satisfies constraint (9b) and other constraints as long as $\varepsilon$ is small enough. However, the total cost will be decreased by $2\varepsilon-\varepsilon=\varepsilon$, which violates the assumption of optimality.
Similarly, it can be proved that $y_{2}^{*}=\max\{\theta_{2}x_{2,1}^{*}/\lambda_{1},\theta_{2}x_{2,2}^{*}/\lambda_{2}\}$. $\square$
Lemma 6.10
When $\Delta_{1}(\lambda_{1},\lambda_{2})\leq 0$, there exist an optimal solution satisfying $x_{1,2}^{*}=x_{2,1}^{*}=0$. When $\Delta_{1}(\lambda_{1},\lambda_{2})>0$, $\theta_{1}x_{1,2}^{*}=\lambda_{2}y_{1}^{*}$ and $\theta_{1}x_{2,1}^{*}=\lambda_{1}y_{2}^{*}$.
Proof 6.11
Proof. Consider the first case where $\Delta_{1}(\lambda_{1},\lambda_{2})\leq 0$. Assume, again by way of contradiction, that either $x_{1,2}^{*}>0$ or $x_{2,1}^{*}>0$. If $x_{1,2}^{*}>0$, let $\hat{x}_{1,2}=x_{1,2}^{*}-\varepsilon$, $\hat{x}_{2,1}=x_{2,1}^{*}$, and let $\hat{y}_{i}=y_{i}^{*}+\varepsilon\theta_{i}/(\theta_{i}+2\lambda_{i})$ and $\hat{x}_{i,i}=x_{i,i}^{*}+\varepsilon\lambda_{i}/(\theta_{i}+2\lambda_{i})$ for $i=1,2$. Since $\hat{y}_{i}+2\hat{x}_{i,i}+\hat{x}_{1,2}+\hat{x}_{2,1}=y_{i}^{*}+2x_{i,i}^{*}+x_{1,2}^{*}+x_{2,1}^{*}=\lambda_{i}$ ($i=1,2$), (9b) and (9c) are still satisfied. (9d) and (9g) are also satisfied because for $i=1,2$, $\theta_{i}\hat{x}_{i,i}=\theta_{i}x_{i,i}^{*}+\theta_{i}\varepsilon\lambda_{i}/(\theta_{i}+2\lambda_{i})\leq\lambda_{i}y_{i}^{*}+\lambda_{i}\theta_{i}\varepsilon/(\theta_{i}+2\lambda_{i})=\lambda_{i}\hat{y}_{i}$. (9e) and (9f) also hold because $\hat{y}_{1},\hat{y}_{2}$ are larger than $y_{1}^{*}$ and $y_{2}^{*}$, whereas $\hat{x}_{1,2},\hat{x}_{2,1}$ are smaller than $x_{1,2}^{*}$ and $x_{2,1}^{*}$. Lastly, the non-negative requirements (9h) and (9i) are satisfied by choosing a small value of $\varepsilon<x_{1,2}^{*}$. Therefore, the new solution $\boldsymbol{\hat{y}},\boldsymbol{\hat{x}}$ is feasible.
Under this feasible solution, the total cost is given by
$c_{(1)}(\hat{x}_{1,1}+\hat{y}_{1})+c_{(2)}(\hat{x}_{2,2}+\hat{y}_{2})+c_{(1,2)}(\hat{x}_{1,2}+\hat{x}_{2,1})$
$=$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*})+\varepsilon\left[c_{(1)}\frac{\theta_{1}+\lambda_{1}}{\theta_{1}+2\lambda_{1}}+c_{(2)}\frac{\theta_{2}+\lambda_{2}}{\theta_{2}+2\lambda_{2}}-c_{(1,2)}\right]$
$=$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*})+\varepsilon\Delta_{1}(\lambda_{1},\lambda_{2})$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*}),$
where (a) is because of the assumption that $\Delta_{1}(\lambda_{1},\lambda_{2})\leq 0$. Therefore, if $x_{1,2}^{*}>0$, by decreasing the value of $x_{1,2}$, we can always obtain another feasible solution without increasing the value function, meaning that $x_{1,2}^{*}=0$ is an optimal solution. Similarly, $x_{2,1}^{*}=0$.
Now consider the second case where $\Delta_{1}(\lambda_{1},\lambda_{2})>0$. In this case, we show that both (9e) and (9f) are binding at optimality. Again assume the opposite is true and consider the following three sub-cases:
(i) If both (9e) and (9f) are not binding, that is, $\theta_{1}x_{1,2}^{*}<\lambda_{2}y_{1}^{*}$ and $\theta_{2}x_{2,1}^{*}<\lambda_{1}y_{2}^{*}$, we then have $y_{1}^{*}>0$ and $y_{2}^{*}>0$. Also according to Lemma 6.8, (9d) and (9g) are binding, thus $x_{1,1}^{*},x_{2,2}^{*}>0$. We can let $\hat{x}_{1,2}=x_{1,2}^{*}+\varepsilon$, $\hat{x}_{2,1}=x_{2,1}^{*}$, and let $\hat{y}_{i}=y_{i}^{*}-\varepsilon\theta_{i}/(\theta_{i}+2\lambda_{i})$ and $\hat{x}_{i,i}=x_{i,i}^{*}-\varepsilon\lambda_{i}/(\theta_{i}+2\lambda_{i})$ for $i=1,2$, where $\varepsilon>0$. Similar to the first case, one can verify that all constraints are satisfied under this new solution when $\varepsilon$ is sufficiently small, and the new cost is given by
$c_{(1)}(\hat{x}_{1,1}+\hat{y}_{1})+c_{(2)}(\hat{x}_{2,2}+\hat{y}_{2})+c_{(1,2)}(\hat{x}_{1,2}+\hat{x}_{2,1})$
$=$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*})-\varepsilon\Delta_{1}(\lambda_{1},\lambda_{2})$
<$\stackrel{{\scriptstyle(a)}}{{<}}$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*}),$
where (a) is due to $\Delta_{1}(\lambda_{1},\lambda_{2})>0$.
(ii) If only (9f) is not binding, that is, $\theta_{1}x_{1,2}^{*}=\lambda_{2}y_{1}^{*}$ and $\theta_{2}x_{2,1}^{*}<\lambda_{1}y_{2}^{*}$, then we have $y_{2}^{*}>0$. Also according to Lemma 6.8, $\theta_{2}x_{2,2}^{*}=\lambda_{2}y_{2}^{*}$. Here we must have $y_{1}^{*}>0$, otherwise if $y_{1}^{*}=0$, according to (9d), (9e), and (9b), $x_{1,1}^{*}=x_{1,2}^{*}=0$, and $x_{2,1}^{*}=\lambda_{1}$. Thus $y_{2}^{*}>\theta_{2}x_{2,1}^{*}/\lambda_{1}=\theta_{2}$, and $x_{2,2}^{*}=\lambda_{2}y_{2}^{*}/\theta_{2}>\lambda_{2}$, which violates (9c). Therefore, $y_{1}^{*},x_{1,2}^{*}>0$.
We can then construct new feasible solutions with lower costs depending on whether $x_{1,1}^{*}$ is 0. If $x_{1,1}^{*}=0$, then $\theta_{1}x_{1,1}^{*}<\lambda_{1}y_{1}^{*}$. We construct a new feasible solution as follows: Given $\varepsilon>0$, let let $\hat{x}_{2,1}=x_{2,1}^{*}+\varepsilon$, $\hat{y}_{1}=y_{1}^{*}-\varepsilon\theta_{1}/(\theta_{1}+\lambda_{2})$, $\hat{x}_{1,1}=x_{1,1}^{*}$, $\hat{x}_{1,2}=x_{1,2}^{*}-\varepsilon\lambda_{2}/(\theta_{1}+\lambda_{2})$, $\hat{y}_{2}=y_{2}^{*}-\varepsilon\theta_{1}\theta_{2}/\left[(\theta_{1}+\lambda_{2})(\theta_{2}+2\lambda_{2})\right]$, and $\hat{x}_{2,2}=x_{2,2}^{*}-\varepsilon\theta_{1}\lambda_{2}/\left[(\theta_{1}+\lambda_{2})(\theta_{2}+2\lambda_{2})\right]$. Under the new solution $\boldsymbol{\hat{y}},\boldsymbol{\hat{x}}$, (9b) and (9c) are satisfied because:
$2\hat{x}_{1,1}+\hat{x}_{1,2}+\hat{x}_{2,1}+\hat{y}_{1}$
$=$ $2x_{1,1}^{*}+x_{1,2}^{*}+x_{2,1}^{*}+y_{1}^{*}-\frac{\lambda_{2}}{\theta_{1}+\lambda_{2}}\varepsilon+\varepsilon-\frac{\theta_{1}}{\theta_{1}+\lambda_{2}}\varepsilon$
$=$ $\lambda_{1},$
$2\hat{x}_{2,2}+\hat{x}_{1,2}+\hat{x}_{2,1}+\hat{y}_{2}$
$=$ $2x_{2,2}^{*}+x_{1,2}^{*}+x_{2,1}^{*}+y_{2}^{*}-2\frac{\theta_{1}\lambda_{2}}{(\theta_{1}+\lambda_{2})(\theta_{2}+2\lambda_{2})}\varepsilon-\frac{\lambda_{2}}{\theta_{1}+\lambda_{2}}\varepsilon+\varepsilon-\frac{\theta_{1}\theta_{2}}{(\theta_{1}+\lambda_{2})(\theta_{2}+2\lambda_{2})}\varepsilon$
$=$ $\lambda_{2},$
and the remaining constraints are also satisfied as long as $\varepsilon$ is sufficiently small. The new total cost is
$c_{(1)}(\hat{x}_{1,1}+\hat{y}_{1})+c_{(2)}(\hat{x}_{2,2}+\hat{y}_{2})+c_{(1,2)}(\hat{x}_{1,2}+\hat{x}_{2,1})$
$=$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*})$
$-\varepsilon\left[c_{(1)}\frac{\theta_{1}}{\theta_{1}+\lambda_{2}}+c_{(2)}\frac{\theta_{1}(\theta_{2}+\lambda_{2})}{(\theta_{1}+\lambda_{2})(\theta_{2}+2\lambda_{2})}-c_{(1,2)}\frac{\theta_{1}}{\theta_{1}+\lambda_{2}}\right]$
$=$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*})-\varepsilon\frac{\theta_{1}}{\theta_{1}+\lambda_{2}}\left[c_{(1)}+c_{(2)}\frac{\theta_{2}+\lambda_{2}}{\theta_{2}+2\lambda_{2}}-c_{(1,2)}\right]$
$<$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*})-\varepsilon\frac{\theta_{1}}{\theta_{1}+\lambda_{2}}\left[c_{(1)}\frac{\theta_{1}+\lambda_{1}}{\theta_{1}+2\lambda_{1}}+c_{(2)}\frac{\theta_{2}+\lambda_{2}}{\theta_{2}+2\lambda_{2}}-c_{(1,2)}\right]$
$<$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*}).$
If $x_{1,1}^{*}>0$, then we construct another new feasible solution as follows: Given $\varepsilon>0$, let $\hat{x}_{2,1}=x_{2,1}^{*}+\varepsilon$, $\hat{y}_{1}=y_{1}^{*}-\varepsilon\theta_{1}/(\theta_{1}+2\lambda_{1}+\lambda_{2})$, $\hat{x}_{1,1}=x_{1,1}^{*}-\varepsilon\lambda_{1}/(\theta_{1}+2\lambda_{1}+\lambda_{2})$, $\hat{x}_{1,2}=x_{1,2}^{*}-\varepsilon\lambda_{2}/(\theta_{1}+2\lambda_{1}+\lambda_{2})$, $\hat{y}_{2}=y_{2}^{*}-\varepsilon(\theta_{1}+2\lambda_{1})\theta_{2}/\left[(\theta_{1}+2\lambda_{1}+\lambda_{2})(\theta_{2}+2\lambda_{2})\right]$, and $\hat{x}_{2,2}=x_{2,2}^{*}-\varepsilon(\theta_{1}+2\lambda_{1})\lambda_{2}/\left[(\theta_{1}+2\lambda_{1}+\lambda_{2})(\theta_{2}+2\lambda_{2})\right]$. Under the new solution $\boldsymbol{\hat{y}},\boldsymbol{\hat{x}}$, (9b) and (9c) are satisfied because:
$2\hat{x}_{1,1}+\hat{x}_{1,2}+\hat{x}_{2,1}+\hat{y}_{1}$
$=$ $2x_{1,1}^{*}+x_{1,2}^{*}+x_{2,1}^{*}+y_{1}^{*}-2\frac{\lambda_{1}}{\theta_{1}+2\lambda_{1}+\lambda_{2}}\varepsilon-\frac{\lambda_{2}}{\theta_{1}+2\lambda_{1}+\lambda_{2}}\varepsilon+\varepsilon-\frac{\theta_{1}}{\theta_{1}+2\lambda_{1}+\lambda_{2}}\varepsilon$
$=$ $\lambda_{1},$
$2\hat{x}_{2,2}+\hat{x}_{1,2}+\hat{x}_{2,1}+\hat{y}_{2}$
$=$ $2x_{2,2}^{*}+x_{1,2}^{*}+x_{2,1}^{*}+y_{2}^{*}-2\frac{(\theta_{1}+2\lambda_{1})\lambda_{2}}{(\theta_{1}+2\lambda_{1}+\lambda_{2})(\theta_{2}+2\lambda_{2})}\varepsilon-\frac{\lambda_{2}}{\theta_{1}+2\lambda_{1}+\lambda_{2}}\varepsilon+\varepsilon-\frac{(\theta_{1}+2\lambda_{1})\theta_{2}}{(\theta_{1}+2\lambda_{1}+\lambda_{2})(\theta_{2}+2\lambda_{2})}\varepsilon$
$=$ $\lambda_{2},$
and the remaining constraints are also satisfied as long as $\varepsilon$ is sufficiently small. Under $\boldsymbol{\hat{y}},\boldsymbol{\hat{x}}$, the total cost is
$c_{(1)}(\hat{x}_{1,1}+\hat{y}_{1})+c_{(2)}(\hat{x}_{2,2}+\hat{y}_{2})+c_{(1,2)}(\hat{x}_{1,2}+\hat{x}_{2,1})$
$=$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*})$
$-\varepsilon\left[c_{(1)}\frac{\theta_{1}+\lambda_{1}}{\theta_{1}+2\lambda_{1}+\lambda_{2}}+c_{(2)}\frac{(\theta_{2}+\lambda_{2})(\theta_{1}+2\lambda_{1})}{(\theta_{2}+2\lambda_{2})(\theta_{1}+2\lambda_{1}+\lambda_{2})}-c_{(1,2)}\frac{\theta_{1}+2\lambda_{1}}{\theta_{1}+2\lambda_{1}+\lambda_{2}}\right]$
$=$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*})-\varepsilon\frac{\theta_{1}+2\lambda_{1}}{\theta_{1}+2\lambda_{1}+\lambda_{2}}\left[c_{(1)}\frac{\theta_{1}+\lambda_{1}}{\theta_{1}+2\lambda_{1}}+c_{(2)}\frac{\theta_{2}+\lambda_{2}}{\theta_{2}+2\lambda_{2}}-c_{(1,2)}\right]$
$<$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*}),$
leading to the contradiction.
(iii) If only (9e) is not binding, we can also get a contradiction in a similar way to sub-case (ii).
Therefore, both (9e) and (9f) are binding at optimality, that is, $\theta_{1}x_{1,2}^{*}=\lambda_{2}y_{1}^{*}$ and $\theta_{2}x_{2,1}^{*}=\lambda_{1}y_{2}^{*}$ when $\Delta_{1}(\lambda_{1},\lambda_{2})>0$. $\square$
Lemma 6.12
$y_{1}^{*}>0$ and $\theta_{2}x_{2,2}^{*}=\lambda_{2}y_{2}^{*}$.
Proof 6.13
Proof. We prove this lemma by contradiction. First consider the case where $y_{1}^{*}=0$. According to (9d), (9e) and (9h), $x_{1,1}^{*}=x_{1,2}^{*}=0$. Then (9b) leads to $x_{2,1}^{*}=\lambda_{1}>0$, (9f) leads to $y_{2}^{*}\geq\theta_{2}$, and (9c) is equivalent to
$x_{2,2}^{*}=\frac{\lambda_{2}-x_{2,1}^{*}-y_{2}^{*}}{2}\leq\frac{\lambda_{2}-\lambda_{1}-\theta_{2}}{2}<\lambda_{2}\leq\frac{\lambda_{2}y_{2}^{*}}{\theta_{2}},$
meaning that strict inequality holds for (9d).
For the case where $\theta_{2}x_{2,2}^{*}<\lambda_{2}y_{2}^{*}$, strict inequality also holds for (9g). We also have $y_{2}^{*}>0$. According to Lemma 6.8, (9f) must be binding, thus $x_{2,1}^{*}>0$.
Now for both cases, we try to give another feasible solution with a smaller total cost. Given a small value of $\varepsilon>0$, let $\hat{y}_{1}=y_{1}^{*}+\theta_{1}\varepsilon/(\theta_{1}+2\lambda_{1}+\lambda_{2})$, $\hat{x}_{1,1}=x_{1,1}^{*}+\lambda_{1}\varepsilon/(\theta_{1}+2\lambda_{1}+\lambda_{2})$, $\hat{x}_{1,2}=x_{1,2}^{*}+\lambda_{2}\varepsilon/(\theta_{1}+2\lambda_{1}+\lambda_{2})$, $\hat{x}_{2,1}=x_{2,1}^{*}-\varepsilon$, $\hat{y}_{2}=y_{2}^{*}-\theta_{2}\varepsilon/\lambda_{1}$, and $\hat{x}_{2,2}=x_{2,2}^{*}+\left[(\theta_{1}+2\lambda_{1})/(\theta_{1}+2\lambda_{1}+\lambda_{2})+\theta_{2}/\lambda_{1}\right]\varepsilon/2$.
For this new solution $\boldsymbol{\hat{x}},\boldsymbol{\hat{y}}$, one can verify that (9b), (9c), (9d), (9e), and (9f) still hold. For example, for (9b), we have
$2\hat{x}_{1,1}+\hat{x}_{1,2}+\hat{x}_{2,1}+\hat{y}_{1}=2x_{1,1}^{*}+x_{1,2}^{*}+x_{2,1}^{*}+y_{1}^{*}+\left[\frac{\theta_{1}+2\lambda_{1}+\lambda_{2}}{\theta_{1}+2\lambda_{1}+\lambda_{2}}-1\right]\varepsilon=\lambda_{1}.$
For (9g), since $\theta x_{2,2}^{*}<\lambda_{2}y_{2}^{*}$ holds with strict inequality, we can guarantee that $\theta\hat{x}_{2,2}<\lambda_{2}\hat{y}_{2}$ still holds by choosing a small enough value of $\varepsilon$. Similarly, non-negative constraints (9h) and (9i) hold since $x_{2,1}^{*}>0$ and $y_{2}^{*}>0$.
Under this new feasible solution, the total cost is calculated as
$c_{(1)}(\hat{x}_{1,1}+\hat{y}_{1})+c_{(2)}(\hat{x}_{2,2}+\hat{y}_{2})+c_{(1,2)}(\hat{x}_{1,2}+\hat{x}_{2,1})$
$=$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*})$
$+\varepsilon\left[c_{(1)}\frac{\theta_{1}+\lambda_{1}}{\theta_{1}+2\lambda_{1}+\lambda_{2}}+\frac{c_{(2)}}{2}\left(\frac{\theta_{1}+2\lambda_{1}}{\theta_{1}+2\lambda_{1}+\lambda_{2}}-\frac{\theta_{2}}{\lambda_{1}}\right)-c_{(1,2)}\frac{\theta_{1}+2\lambda_{1}}{\theta_{1}+2\lambda_{1}+\lambda_{2}}\right]$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*})$
$+\varepsilon\left[c_{(1)}\frac{\theta_{1}+\lambda_{1}}{\theta_{1}+2\lambda_{1}+\lambda_{2}}+\frac{c_{(2)}}{2}\left(\frac{\theta_{1}+2\lambda_{1}}{\theta_{1}+2\lambda_{1}+\lambda_{2}}-\frac{\theta_{2}}{\lambda_{1}}\right)-c_{(1)}\frac{\theta_{1}+\lambda_{1}}{\theta_{1}+2\lambda_{1}+\lambda_{2}}-c_{(2)}\frac{\lambda_{1}}{\theta_{1}+2\lambda_{1}+\lambda_{2}}\right]$
$=$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*})+\varepsilon c_{(2)}\left(\frac{\theta_{1}}{\theta_{1}+2\lambda_{1}+\lambda_{2}}-\frac{\theta_{2}}{\lambda_{1}}\right)$
<$\stackrel{{\scriptstyle(b)}}{{<}}$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*}),$
where (a) is because $c_{(1,2)}\geq c_{(1)}$ and $c_{(1,2)}\geq c_{(2)}$, and (b) is because $\theta_{1}\leq\theta_{2}$. Therefore, a new feasible solution with a lower cost is obtained, leading to a contradiction. $\square$
Lemma 6.14
$y_{2}^{*}>0$ if $\Delta_{2}(\lambda_{1},\lambda_{2})<0$ or $\Delta_{3}(\lambda_{1},\lambda_{2})<0$, and $\theta_{1}x_{1,1}^{*}=\lambda_{1}y_{1}^{*}$ if $\Delta_{2}(\lambda_{1},\lambda_{2})<0$.
Proof 6.15
Proof.
The idea of proving this lemma is the same as Lemma 6.12. Again prove this lemma by contradiction. Consider the case where $y_{2}^{*}=0$. We have $x_{2,1}^{*}=x_{2,2}^{*}=0,x_{1,2}^{*}=\lambda_{2}>0$. (9e) leads to $y_{1}^{*}\geq\theta_{1}$, and (9b) leads to
$x_{1,1}^{*}=\frac{\lambda_{1}-x_{1,2}^{*}-y_{1}^{*}}{2}\leq\frac{\lambda_{1}-\lambda_{2}-\theta_{1}}{2}=\frac{\Delta_{3}(\lambda_{1},\lambda_{2})}{2}.$
Thus if $\Delta_{3}(\lambda_{1},\lambda_{2})<0$, $x_{1,1}^{*}$ will become negative, which immediately leads to the contradiction.
When the feasibility is satisfied, we continue by following the proving process of Lemma 6.12 but swap all indexes of 1 and 2. Similarly we can define a solution $\boldsymbol{\hat{x}},\boldsymbol{\hat{y}}$ with the total cost is calculated as
$c_{(1)}(\hat{x}_{1,1}+\hat{y}_{1})+c_{(2)}(\hat{x}_{2,2}+\hat{y}_{2})+c_{(1,2)}(\hat{x}_{1,2}+\hat{x}_{2,1})$
$=$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*})$
$+\varepsilon\left[\frac{c_{(1)}}{2}\left(\frac{\theta_{2}+2\lambda_{2}}{\theta_{2}+\lambda_{1}+2\lambda_{2}}-\frac{\theta_{1}}{\lambda_{2}}\right)+c_{(2)}\frac{\theta_{2}+\lambda_{2}}{\theta_{2}+\lambda_{1}+2\lambda_{2}}-c_{(1,2)}\frac{\theta_{2}+2\lambda_{2}}{\theta_{2}+\lambda_{1}+2\lambda_{2}}\right]$
$=$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*})+\varepsilon\frac{\theta_{2}+2\lambda_{2}}{\theta_{2}+\lambda_{1}+2\lambda_{2}}\Delta_{2}(\lambda_{1},\lambda_{2}),$
thus if $\Delta_{2}(\lambda_{1},\lambda_{2})<0$, a new feasible solution with a lower cost is obtained, also leading to a contradiction. $\square$
Using the above lemmas, we are now ready to give a proof of Lemma 3.4.
Proof 6.16
Proof of Lemma 3.4.
First we notice that
$\Delta_{1}(\lambda_{1},\lambda_{2})-\Delta_{2}(\lambda_{1},\lambda_{2})$ $=c_{(1)}\frac{\theta_{1}+\lambda_{1}}{\theta_{1}+2\lambda_{1}}-\frac{c_{(1)}}{2}\left[1-\frac{\theta_{1}(\theta_{2}+\lambda_{1}+2\lambda_{2})}{\lambda_{2}(\theta_{2}+2\lambda_{2})}\right]$
$>\frac{c_{(1)}}{2}-\frac{c_{(1)}}{2}=0.$
Below we prove the four cases one by one.
Case (i): $\Delta_{1}(\lambda_{1},\lambda_{2})\leq 0$, and thus $\Delta_{2}(\lambda_{1},\lambda_{2})<0$. According to Lemmas 6.12 and 6.14, $y_{1}^{*},y_{2}^{*}>0$, and $\theta_{i}x_{i,i}^{*}=\lambda_{i}y_{i}^{*}$, $i=1,2$. According to Lemma 6.10, there exists an optimal solution satisfying $x_{1,2}^{*}=x_{2,1}^{*}=0$.
Case (ii): $\Delta_{1}(\lambda_{1},\lambda_{2})>0,\Delta_{2}(\lambda_{1},\lambda_{2})<0$. According to Lemmas 6.12 and 6.14, $y_{1}^{*},y_{2}^{*}>0$, and $\theta_{i}x_{i,i}^{*}=\lambda_{i}y_{i}^{*}$, $i=1,2$. According to Lemma 6.10, the optimal solution must also satisfy $\theta_{1}x_{1,2}^{*}=\lambda_{2}y_{1}^{*}$ and $\theta_{1}x_{2,1}^{*}=\lambda_{1}y_{2}^{*}$.
Case (iii): $\Delta_{2}(\lambda_{1},\lambda_{2})\geq 0,\Delta_{3}(\lambda_{1},\lambda_{2})<0$, and thus $\Delta_{1}(\lambda_{1},\lambda_{2})>0$. According to Lemmas 6.12 and 6.14, $y_{1}^{*},y_{2}^{*}>0$, and $\theta_{2}x_{2,2}^{*}=\lambda_{2}y_{2}^{*}$. According to Lemma 6.10, $\theta_{1}x_{1,2}^{*}=\lambda_{2}y_{1}^{*}$ and $\theta_{1}x_{2,1}^{*}=\lambda_{1}y_{2}^{*}$. Then all we need is to prove that there is an optimal solution with $x_{1,1}^{*}=0$.
Assume for contradiction that $x_{1,1}^{*}>0$. Again we can re-use the same method to construct another feasible solution as in Lemmas 6.12 and 6.14. Given a small value of $\varepsilon>0$, let $\hat{y}_{2}=y_{2}^{*}-\theta_{2}\varepsilon/(\theta_{2}+\lambda_{1}+2\lambda_{2})$, $\hat{x}_{2,2}=x_{2,2}^{*}-\lambda_{2}\varepsilon/(\theta_{2}+\lambda_{1}+2\lambda_{2})$, $\hat{x}_{2,1}=x_{2,1}^{*}-\lambda_{1}\varepsilon/(\theta_{2}+\lambda_{1}+2\lambda_{2})$, $\hat{x}_{1,2}=x_{1,2}^{*}+\varepsilon$, $\hat{y}_{1}=y_{1}^{*}+\theta_{1}\varepsilon/\lambda_{2}$, and $\hat{x}_{1,1}=x_{1,1}^{*}-\left[(\theta_{2}+2\lambda_{2})/(\theta_{2}+\lambda_{1}+2\lambda_{2})+\theta_{1}/\lambda_{2}\right]\varepsilon/2$. The new total cost is calculated as
$c_{(1)}(\hat{x}_{1,1}+\hat{y}_{1})+c_{(2)}(\hat{x}_{2,2}+\hat{y}_{2})+c_{(1,2)}(\hat{x}_{1,2}+\hat{x}_{2,1})$
$=$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*})$
$-\varepsilon\left[\frac{c_{(1)}}{2}\left(\frac{\theta_{2}+2\lambda_{2}}{\theta_{2}+\lambda_{1}+2\lambda_{2}}-\frac{\theta_{1}}{\lambda_{2}}\right)+c_{(2)}\frac{\theta_{2}+\lambda_{2}}{\theta_{2}+\lambda_{1}+2\lambda_{2}}-c_{(1,2)}\frac{\theta_{2}+2\lambda_{2}}{\theta_{2}+\lambda_{1}+2\lambda_{2}}\right]$
$=$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*})-\varepsilon\frac{\theta_{2}+2\lambda_{2}}{\theta_{2}+\lambda_{1}+2\lambda_{2}}\Delta_{2}(\lambda_{1},\lambda_{2})$
$\leq$ $c_{(1)}(x_{1,1}^{*}+y_{1}^{*})+c_{(2)}(x_{2,2}^{*}+y_{2}^{*})+c_{(1,2)}(x_{1,2}^{*}+x_{2,1}^{*}),$
thus by decreasing $x_{1,1}^{*}$, we can obtain another feasible solution with a total cost no more than the original cost. This is sufficient to prove that $x_{1,1}^{*}=0$ corresponds to an optimal solution.
Case (iv): $\Delta_{2}(\lambda_{1},\lambda_{2})\geq 0,\Delta_{3}(\lambda_{1},\lambda_{2})\geq 0$, and thus $\Delta_{1}(\lambda_{1},\lambda_{2})>0$. Lemma 6.12 leads to $y_{1}^{*}>0$ and $\theta_{2}x_{2,2}^{*}=\lambda_{2}y_{2}^{*}$. According to Lemma 6.10, $\theta_{1}x_{1,2}^{*}=\lambda_{2}y_{1}^{*}$ and $\theta_{1}x_{2,1}^{*}=\lambda_{1}y_{2}^{*}$. Then all we need is to prove that $y_{2}^{*}=0$.
Again assume $y_{2}^{*}>0$. Since $\theta_{1}x_{1,2}^{*}=\lambda_{2}y_{1}^{*}$ and from (9c), $x_{1,2}^{*}<\lambda_{2}$, we have $y_{1}^{*}<\theta_{1}$. By subtracting (9c) from (9b), we obtain
$\lambda_{1}-\lambda_{2}=(y_{1}^{*}+2x_{1,1}^{*})-(y_{2}^{*}+2x_{2,2}^{*})<\theta_{1}+2x_{1,1}^{*}.$
Since $\Delta_{3}(\lambda_{1},\lambda_{2})\geq 0$, $\lambda_{1}-\lambda_{2}\geq\theta_{1}$. Therefore, in order to satisfy $\lambda_{1}-\lambda_{2}<\theta_{1}+2x_{1,1}^{*}$, we must have $x_{1,1}^{*}>0$.
Given $y_{2}^{*},x_{2,1}^{*},x_{2,2}^{*},x_{1,1}^{*}>0$, we can immediately re-use the idea in case (iii) to construct a new feasible solution by decreasing $y_{2}^{*}$. The new feasible solution has a total cost no more than the original cost, thus $y_{2}^{*}=0$ corresponds to an optimal solution.
By fully characterizing the optimal solutions, we can explicitly determine the values of $y_{1}^{*}$ and $y_{2}^{*}$ under four different cases, thereby providing the optimal solutions in closed form.
Case (i): Solving (9b) and (9c) when (9d), (9g) are binding and $x_{1,2}^{*}=x_{2,1}^{*}=0$, we obtain
$y_{1}^{*}=\frac{\theta_{1}\lambda_{1}}{\theta_{1}+2\lambda_{1}},\ y_{2}^{*}=\frac{\theta_{2}\lambda_{2}}{\theta_{2}+2\lambda_{2}},$
and $x_{1,2}^{*}=x_{2,1}^{*}=0$, $x_{1,1}^{*}=\lambda_{1}y_{1}^{*}/\theta_{1}$, $x_{2,2}^{*}=\lambda_{2}y_{2}^{*}/\theta_{2}$.
Case (ii): Solving (9b) and (9c) when (9d), (9e), (9f) and (9g) are all binding, we obtain
$y_{1}^{*}$ $=\frac{\theta_{1}\lambda_{1}(\theta_{2}+\lambda_{1}+\lambda_{2})}{2(\lambda_{1}+\lambda_{2})^{2}+\theta_{1}(\lambda_{1}+2\lambda_{2})+\theta_{2}(\lambda_{2}+2\lambda_{1})+\theta_{1}\theta_{2}},$
$y_{2}^{*}$ $=\frac{\theta_{2}\lambda_{2}(\theta_{1}+\lambda_{1}+\lambda_{2})}{2(\lambda_{1}+\lambda_{2})^{2}+\theta_{1}(\lambda_{1}+2\lambda_{2})+\theta_{2}(\lambda_{2}+2\lambda_{1})+\theta_{1}\theta_{2}},$
and $x_{1,1}^{*}=\lambda_{1}y_{1}^{*}/\theta_{1}$,$x_{1,2}^{*}=\lambda_{2}y_{1}^{*}/\theta_{1}$,$x_{2,1}^{*}=\lambda_{1}y_{2}^{*}/\theta_{2}$,$x_{2,2}^{*}=\lambda_{2}y_{2}^{*}/\theta_{2}$.
Case (iii): Solving (9b) and (9c) subject to the binding constraints (9e), (9f), (9g), and $x_{1,1}^{*}=0$, we obtain
$y_{1}^{*}$ $=\frac{\theta_{1}\lambda_{1}(\theta_{2}+\lambda_{1}+\lambda_{2})}{2\lambda_{2}^{2}+\theta_{1}(\lambda_{1}+2\lambda_{2})+\theta_{2}\lambda_{2}+\theta_{1}\theta_{2}},$
$y_{2}^{*}$ $=\frac{\theta_{2}\lambda_{2}(\theta_{1}+\lambda_{1}+\lambda_{2})}{2\lambda_{2}^{2}+\theta_{1}(\lambda_{1}+2\lambda_{2})+\theta_{2}\lambda_{2}+\theta_{1}\theta_{2}},$
and $x_{1,1}^{*}=0$,$x_{1,2}^{*}=\lambda_{2}y_{1}^{*}/\theta_{1}$,$x_{2,1}^{*}=\lambda_{1}y_{2}^{*}/\theta_{2}$,$x_{2,2}^{*}=\lambda_{2}y_{2}^{*}/\theta_{2}$.
Case (iv): Solving (9b) and (9c) subject to $y_{2}^{*}=x_{2,1}^{*}=x_{2,2}^{*}=0$ and the binding constraint (9e), we obtain
$y_{1}^{*}=\theta_{1},\ x_{1,1}^{*}=\frac{\lambda_{1}-\lambda_{2}-\theta_{1}}{2},\ x_{1,2}^{*}=\lambda_{2}.$
$\square$
6.6 Proofs of Theorem 3.5 and Proposition 3.6
According to Lemma 3.4, $y_{1}^{*}>0$ is guaranteed, and $y_{2}^{*}>0$ if and only if $\Delta_{2}(\lambda_{1},\lambda_{2})<0$ or $\Delta_{3}(\lambda_{1},\lambda_{2})<0$ holds for all $(\lambda_{1},\lambda_{2})\in\Lambda$. Define the set $\mathcal{D}:=\{(\lambda_{1},\lambda_{2})|\Delta_{2}(\lambda_{1},\lambda_{2})\geq 0,\Delta_{3}(\lambda_{1},\lambda_{2})\geq 0,\lambda_{1},\lambda_{2}>0\}$, then we have the following two results: (i) If $\Lambda\cap\mathcal{D}=\emptyset$, then $\boldsymbol{y}^{*}>\boldsymbol{0}$ always hold for $\boldsymbol{\lambda}\in\Lambda$. According to Lemma 3.3, $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$; (ii) If $\Lambda\cap\mathcal{D}\neq\emptyset$, there exists $\boldsymbol{\lambda}\in\Lambda$, such that $y_{2}^{*}=0$.
For the set $\mathcal{D}$, note that
$\Delta_{2}(\lambda_{1},\lambda_{2})\geq 0$
[…]
where $\tau_{1}=c_{(1)}(\theta_{2}-3\theta_{1})+2c_{(2)}\theta_{2}-2c_{(1,2)}\theta_{2}$, and $\Delta_{3}(\lambda_{1},\lambda_{2})\geq 0$ is equivalent to $\lambda_{1}\geq\lambda_{2}+\theta_{1}$. Thus
$\mathcal{D}=\left\{(\lambda_{1},\lambda_{2})|\lambda_{2}+\theta_{1}\leq\lambda_{1}\leq-2\frac{2c_{(1,2)}-c_{(1)}-c_{(2)}}{c_{(1)}\theta_{1}}\lambda_{2}^{2}+\left(\frac{\tau_{1}}{c_{(1)}\theta_{1}}+1\right)\lambda_{2}-\theta_{2},\lambda_{1},\lambda_{2}>0\right\}.$
Now we prove Theorem 3.5 and Proposition 3.6 separately.
Proof 6.17
Proof of Theorem 3.5. According to Assumption 2, $2c_{(1,2)}-c_{(1)}-c_{(2)}\geq 0$. First consider the case where $02c_{(1,2)}-c_{(1)}-c_{(2)}=$ 0, then $\tau_{2}=\tau_{1}^{2}\geq 0$. If $\tau_{1}\leq 0$ in this case, $\mathcal{D}$ is equivalent to $\left\{(\lambda_{1},\lambda_{2})|\lambda_{2}+\theta_{1}\leq\lambda_{1}\leq\left(\tau_{1}/(c_{(1)}\theta_{1})+1\right)\lambda_{2}-\theta_{2},\lambda_{1},\lambda_{2}>0\right\}$, which is an empty set because $\left(\tau_{1}/(c_{(1)}\theta_{1})+1\right)\lambda_{2}-\theta_{2}\leq\lambda_{2}-\theta_{2}<\lambda_{2}+\theta_{1}$ for any $\lambda_{1},\lambda_{2}>0$. Thus $\Lambda\cap\mathcal{D}=\emptyset$ when $02c_{(1,2)}-c_{(1)}-c_{(2)}=$ 0 and $\tau_{1}\leq 0$.
For the remaining of the proof, we focus on the case of $02c_{(1,2)}-c_{(1)}-c_{(2)}>$ 0. In this case, $\lambda_{1}$ in $\mathcal{D}$ is larger than a linear function of $\lambda_{2}$ and smaller than a quadratic function with a negative leading coefficient. Then a sufficient condition of $\Lambda\cap\mathcal{D}=\emptyset$ is for any $\lambda_{2}\in[\underline{\lambda}_{2},\lambda_{\max,2}]$,
$\lambda_{2}+\theta_{1}>-2\frac{2c_{(1,2)}-c_{(1)}-c_{(2)}}{c_{(1)}\theta_{1}}\lambda_{2}^{2}+\left(\frac{\tau_{1}}{c_{(1)}\theta_{1}}+1\right)\lambda_{2}-\theta_{2}$
$\Leftrightarrow$ $-\lambda_{2}^{2}+\frac{\tau_{1}}{2(2c_{(1,2)}-c_{(1)}-c_{(2)})}\lambda_{2}-\frac{c_{(1)}\theta_{1}(\theta_{1}+\theta_{2})}{2(2c_{(1,2)}-c_{(1)}-c_{(2)})}<0$
$\Leftrightarrow$ $-\left[\lambda_{2}-\frac{\tau_{1}}{4(2c_{(1,2)}-c_{(1)}-c_{(2)})}\right]^{2}+\frac{\tau_{2}}{16(2c_{(1,2)}-c_{(1)}-c_{(2)})^{2}}<0.$
If $\tau_{2}<0$, the above inequality naturally holds for any $\lambda_{2}\in R$. If $\tau_{2}\geq 0$, the above inequality leads to
$\lambda_{2}>\frac{\tau_{1}+\sqrt{\tau_{2}}}{4(2c_{(1,2)}-c_{(1)}-c_{(2)})},\ \text{or}\ \lambda_{2}<\frac{\tau_{1}-\sqrt{\tau_{2}}}{4(2c_{(1,2)}-c_{(1)}-c_{(2)})}.$
When $\tau_{1}\leq 0$,
$\tau_{1}+\sqrt{\tau_{2}}=\tau_{1}+\sqrt{\tau_{1}^{2}-8c_{(1)}(2c_{(1,2)}-c_{(1)}-c_{(2)})\theta_{1}(\theta_{1}+\theta_{2})}\leq 0,$
meaning that the inequality always holds when $\lambda_{2}>0$. Otherwise, when $\tau_{1}>0$ and $\tau_{2}\geq 0$, a sufficient condition to make the inequality hold is when
$\underline{\lambda}_{2}>\frac{\tau_{1}+\sqrt{\tau_{2}}}{4(2c_{(1,2)}-c_{(1)}-c_{(2)})}.$
$\square$
Proof 6.18
Proof of Proposition 3.6
Similar to the first step of proving Theorem 3.5, when $02c_{(1,2)}-c_{(1)}-c_{(2)}=$ 0, $\mathcal{D}$ is equivalent to $\left\{(\lambda_{1},\lambda_{2})|\lambda_{2}+\theta_{1}\leq\lambda_{1}\leq\left(\tau_{1}/(c_{(1)}\theta_{1})+1\right)\lambda_{2}-\theta_{2},\lambda_{1},\lambda_{2}>0\right\}$. Given any $\underline{\lambda}_{1},\underline{\lambda}_{2}>0$, Let
$\lambda_{2}$ $=\max\left\{\frac{c_{(1)}\theta_{1}(\theta_{1}+\theta_{2})}{\tau_{1}},\underline{\lambda}_{1}-\theta_{1},\underline{\lambda}_{2}\right\},$
$\lambda_{1}$ $=\lambda_{2}+\theta_{1},$
then it can be verified that $\boldsymbol{\lambda}>\underline{\boldsymbol{\lambda}}$ and $\boldsymbol{\lambda}\in\mathcal{D}$, thus we find a value of $\boldsymbol{\lambda}\in\Lambda$ such that $y_{2}^{*}=0$. $\square$
6.7 Proof of Corollary 3.7
For the first case when $\theta_{2}<3\theta_{1}$, we have
$\tau_{1}=c_{(1)}(\theta_{2}-3\theta_{1})+2c_{(2)}\theta_{2}-2c_{(1,2)}\theta_{2}<2c_{(2)}\theta_{2}-2c_{(1,2)}\theta_{2}<0.$
Then according to the first case of Theorem 3.5, $c(\boldsymbol{\lambda})$ is weakly concave.
In the second case, according to the definition of $e_{1,2}$, we have
$\underline{\lambda}_{2}>\frac{\theta_{2}}{4}\frac{1}{1-2e_{1,2}}$ $=\frac{\theta_{2}}{4}\frac{c_{(1)}+c_{(2)}}{2c_{(1,2)}-c_{(1)}-c_{(2)}}\stackrel{{\scriptstyle(a)}}{{\geq}}\frac{2\theta_{2}\left(c_{(1)}+2c_{(2)}-2c_{(1,2)}\right)}{4\left(2c_{(1,2)}-c_{(1)}-c_{(2)}\right)}>\frac{2\tau_{1}}{4\left(2c_{(1,2)}-c_{(1)}-c_{(2)}\right)}.$
For (a), $c_{(1)}+c_{(2)}\geq 2\left(c_{(1)}+2c_{(2)}-2c_{(1,2)}\right)$ because $c_{(1,2)}\geq c_{(1)}$ and $c_{(1,2)}\geq c_{(2)}$. If $\tau_{1}\leq 0$ or $\tau_{2}<0$, the result holds according to the first case of Theorem 3.5. If $\tau_{1}>0$ and $\tau_{2}\geq 0$, note that
$\sqrt{\tau_{2}}=\sqrt{\tau_{1}^{2}-8c_{(1)}(2c_{(1,2)}-c_{(1)}-c_{(2)})\theta_{1}(\theta_{1}+\theta_{2})}<\sqrt{\tau_{1}^{2}}=\tau_{1},$
and thus
$\underline{\lambda}_{2}>\frac{2\tau_{1}}{4\left(2c_{(1,2)}-c_{(1)}-c_{(2)}\right)}>\frac{\tau_{1}+\sqrt{\tau_{2}}}{4\left(2c_{(1,2)}-c_{(1)}-c_{(2)}\right)}.$
According to the second case of Theorem 3.5, $c(\boldsymbol{\lambda})$ is weakly concave. $\square$
6.8 Proof of Corollary 3.8
Notice that when $\theta_{1}=\theta_{2}=\theta$,
$(\theta+2\lambda_{2})\Delta_{2}(\lambda_{1},\lambda_{2})$ $=\frac{c_{(1)}}{2}\left[(\theta+2\lambda_{2})-\frac{\theta(\theta+\lambda_{1}+2\lambda_{2})}{\lambda_{2}}\right]+c_{(2)}(\theta+\lambda_{2})-c_{(1,2)}(\theta+2\lambda_{2})$
$\leq\frac{c_{(1)}}{2}\left[(\theta+2\lambda_{2})-\frac{\theta(\theta+\lambda_{1}+2\lambda_{2})}{\lambda_{2}}\right]+c_{(2)}(\theta+\lambda_{2})-c_{(1)}\lambda_{2}-c_{(2)}(\theta+\lambda_{2})$
$=\frac{c_{(1)}\theta}{2}\left(1-\frac{\theta+\lambda_{1}+2\lambda_{2}}{\lambda_{2}}\right)$
$<0,$
thus $\Delta_{2}(\lambda_{1},\lambda_{2})<0$ always hold and we do not need to consider the last two cases in Lemma 3.4.
For the first two cases in Lemma 3.4, we can solve (9b) and (9c) with respect to $y_{1}^{*}$ and $y_{2}^{*}$, and obtain the optimal solution. This leads to the results stated in this proposition. $\square$
6.9 Proof of Theorem 3.9
Lemma 6.19
Given any 2-dimensional vector $\boldsymbol{\alpha}$ and any nonnegative vector $\boldsymbol{\beta}\geq 0$, the function $(\boldsymbol{\alpha}^{T}\boldsymbol{\lambda})^{2}/(\theta+\boldsymbol{\beta}^{T}\boldsymbol{\lambda})$ is convex with respect to $\boldsymbol{\lambda}\geq\boldsymbol{0}$.
Proof 6.20
Proof. Define the function $f(a,b)=a^{2}/b$, we can verify that $f(a,b)$ is convex when $b>0$, based on its positive semidefinite Hessian matrix $\begin{bmatrix}2a^{2}/b^{3}&-2a/b^{2}\\ -2a/b^{2}&2/b\end{bmatrix}$: the diagonal elements are positive, and the determinant is 0. Then the function $f((\boldsymbol{\alpha}^{T}\boldsymbol{\lambda})^{2},\theta+\boldsymbol{\beta}^{T}\boldsymbol{\lambda})$ is also convex since $\theta+\boldsymbol{\beta}^{T}\boldsymbol{\lambda}>\boldsymbol{0}$ and an affine mapping preserves convexity (see Section 3.2.2, Boyd and Vandenberghe 2004).
Proof 6.21
Proof of Theorem 3.9. Based on the results of Corollary 3.8, define two cost functions as follows:
$c^{(1)}(\boldsymbol{\lambda})$ $=\frac{c_{(1)}\lambda_{1}(\theta+\lambda_{1})}{\theta+2\lambda_{1}}+\frac{c_{(2)}\lambda_{2}(\theta+\lambda_{2})}{\theta+2\lambda_{2}},$
$c^{(2)}(\boldsymbol{\lambda})$ $=\frac{c_{(1)}\lambda_{1}(\theta+\lambda_{1})+c_{(2)}\lambda_{2}(\theta+\lambda_{2})+2c_{(1,2)}\lambda_{1}\lambda_{2}}{\theta+2\lambda_{1}+2\lambda_{2}}.$
where the first function $c^{(1)}(\boldsymbol{\lambda})$ corresponds to the optimal cost function when $\Delta_{1}(\lambda_{1},\lambda_{2})\leq 0$, and $c^{(2)}(\boldsymbol{\lambda})$ is the optimal cost when $\Delta_{1}(\lambda_{1},\lambda_{2})>0$.
For $c^{(1)}(\boldsymbol{\lambda})$, notice that
$c^{(1)}(\boldsymbol{\lambda})$ $=\frac{c_{(1)}\lambda_{1}(\theta+\lambda_{1})}{\theta+2\lambda_{1}}+\frac{c_{(2)}\lambda_{2}(\theta+\lambda_{2})}{\theta+2\lambda_{2}}$
$=c_{(1)}\lambda_{1}+c_{(2)}\lambda_{2}-c_{(1)}\frac{\lambda_{1}^{2}}{\theta+2\lambda_{1}}-c_{(2)}\frac{\lambda_{2}^{2}}{\theta+2\lambda_{2}}.$
Using Lemma 6.19, by choosing $\boldsymbol{\alpha}=(1,0)^{T}$ and $\boldsymbol{\beta}=(2,0)^{T}$, we obtain the result that $\lambda_{1}^{2}/(\theta+2\lambda_{1})$ is convex. Similarly, by choosing $\boldsymbol{\alpha}=(0,1)^{T}$ and $\boldsymbol{\beta}=(0,2)^{T}$, $\lambda_{2}^{2}/(\theta+2\lambda_{2})$ is convex. $c^{(1)}(\boldsymbol{\lambda})$ is expressed as a linear function minus convex functions, and is therefore concave.
Note that if $c_{(1)}+c_{(2)}-c_{(1,2)}\leq 0$,
$\Delta_{1}(\lambda_{1},\lambda_{2})=c_{(1)}\frac{\theta+\lambda_{1}}{\theta+2\lambda_{1}}+c_{(2)}\frac{\theta+\lambda_{2}}{\theta+2\lambda_{2}}-c_{(1,2)}<c_{(1)}+c_{(2)}-c_{(1,2)}\leq 0,$
thus in this case, $c(\boldsymbol{\lambda})=c^{(1)}(\boldsymbol{\lambda})$ and is concave. The following analysis considers the case when $c_{(1)}+c_{(2)}-c_{(1,2)}>0$.
For $c^{(2)}(\boldsymbol{\lambda})$, we can reformulate the function as
$c^{(2)}(\boldsymbol{\lambda})=$ $\frac{c_{(1)}\lambda_{1}(\theta+\lambda_{1})+c_{(2)}\lambda_{2}(\theta+\lambda_{2})+2c_{(1,2)}\lambda_{1}\lambda_{2}}{\theta+2\lambda_{1}+2\lambda_{2}}$
$=$ $c_{(1)}\lambda_{1}+c_{(2)}\lambda_{2}-\frac{c_{(1)}\lambda_{1}^{2}+c_{(2)}\lambda_{2}^{2}+2[c_{(1)}+c_{(2)}-c_{(1,2)}]\lambda_{1}\lambda_{2}}{\theta+2\lambda_{1}+2\lambda_{2}}$
$=$ $c_{(1)}\lambda_{1}+c_{(2)}\lambda_{2}-[c_{(1,2)}-c_{(2)}]\frac{\lambda_{1}^{2}}{\theta+2\lambda_{1}+2\lambda_{2}}-[c_{(1,2)}-c_{(1)}]\frac{\lambda_{2}^{2}}{\theta+2\lambda_{1}+2\lambda_{2}}$
$-[c_{(1)}+c_{(2)}-c_{(1,2)}]\frac{(\lambda_{1}+\lambda_{2})^{2}}{\theta+2\lambda_{1}+2\lambda_{2}}.$
Again, according to Lemma 6.19, we can prove that $\lambda_{1}^{2}/(\theta+2\lambda_{1}+2\lambda_{2})$, $\lambda_{2}^{2}/(\theta+2\lambda_{1}+2\lambda_{2})$, and $(\lambda_{1}+\lambda_{2})^{2}/(\theta+2\lambda_{1}+2\lambda_{2})$ are all convex, by choosing $\boldsymbol{\beta}=(2,2)^{T}$, $\boldsymbol{\alpha}=(1,0)^{T},(0,1)^{T}$ and $(1,1)^{T}$ respectively. Since $c_{(1,2)}\geq c_{(1)}$, $c_{(1,2)}\geq c_{(2)}$, and $c_{(1,2)}<c_{(1)}+c_{(2)}$, thus $c^{(1)}(\boldsymbol{\lambda})$ can also be expressed as a linear function minus convex functions, and is therefore concave.
Since two types of optimal solutions given in Corollary 3.8 remain feasible for any $\lambda_{1},\lambda_{2}\geq 0$, we can re-write the optimal cost function $c(\boldsymbol{\lambda})$ as the minimization of two functions:
$c(\boldsymbol{\lambda})$ $=\min\{c^{(1)}(\boldsymbol{\lambda}),c^{(2)}(\boldsymbol{\lambda})\}.$
Given that the minimization preserves concavity, $c(\boldsymbol{\lambda})$ is concave. $\square$
6.10 Proofs of Theorem 3.10 and Theorem 3.11
Our proof relies on the following lemma characterizing the optimal solution structure.
Lemma 6.22
$\forall i,j\in[N]$, if
$c_{(i)}\frac{\theta_{i}+\lambda_{i}}{\theta_{i}+2\lambda_{i}}+c_{(j)}\frac{\theta_{j}+\lambda_{j}}{\theta_{j}+2\lambda_{j}}-c_{(i,j)}<0,$
the optimal solution must have $x_{i,j}^{*}=x_{j,i}^{*}=0$.
Proof 6.23
Proof. This lemma is a direct generalization of the first part of Lemma 6.10. Assume by contradiction that there exist $i,j\in[N]$, such that either $x_{i,j}^{*}>0$ or $x_{j,i}^{*}>0$ when $c_{(i)}(\theta_{i}+\lambda_{i})/(\theta_{i}+2\lambda_{i})+c_{(j)}(\theta_{j}+\lambda_{j})/(\theta_{j}+2\lambda_{j})-c_{(i,j)}<0$. If $x_{i,j}^{*}>0$, we obtain a new solution by letting
$\hat{x}_{i,j}$ $=x_{i,j}^{*}-\varepsilon,$
$\hat{y}_{i}$ $=y_{i}^{*}+\varepsilon\frac{\theta_{i}}{\theta_{i}+2\lambda_{i}},$
$\hat{x}_{i,i}$ $=x_{i,i}^{*}+\varepsilon\frac{\lambda_{i}}{\theta_{i}+2\lambda_{i}},$
$\hat{y}_{j}$ $=y_{j}^{*}+\varepsilon\frac{\theta_{j}}{\theta_{j}+2\lambda_{j}},$
$\hat{x}_{j,j}$ $=x_{j,j}^{*}+\varepsilon\frac{\lambda_{j}}{\theta_{j}+2\lambda_{j}},$
and keeping the remaining variables unchanged, where $\varepsilon>0$ is a sufficiently small positive value. Similar to the way in Lemma 6.10, we can verify this is still a feasible solution to (2), since we guarantee that $2\hat{x}_{i,i}+\hat{x}_{i,j}+\hat{y}_{i}=2x_{i,i}^{*}+x_{i,j}^{*}+y_{i}^{*}$ and $2\hat{x}_{j,j}+\hat{x}_{i,j}+\hat{y}_{j}=2x_{j,j}^{*}+x_{i,j}^{*}+y_{j}^{*}$. Compared to the original solution, the difference in cost under this new feasible solution is given by
$\left[c_{(i)}(\hat{x}_{i,i}+\hat{y}_{i})+c_{(j)}(\hat{x}_{j,j}+\hat{y}_{j})+c_{(i,j)}\hat{x}_{i,j}\right]-\left[c_{(i)}(x_{i,i}^{*}+y_{i}^{*})+c_{(j)}(x_{j,j}^{*}+y_{j}^{*})+c_{(i,j)}x_{i,j}^{*}\right]$
$=$ $\varepsilon\left[c_{(i)}\frac{\theta_{i}+\lambda_{i}}{\theta_{i}+2\lambda_{i}}+c_{(j)}\frac{\theta_{j}+\lambda_{j}}{\theta_{j}+2\lambda_{j}}-c_{(i,j)}\right]$
$<$ $0$
Therefore, if $x_{i,j}^{*}>0$, by decreasing the value of $x_{i,j}$, we can always obtain another feasible solution with lower cost, meaning that $x_{i,j}^{*}=0$ must be an optimal solution. Similarly, $x_{j,i}^{*}=0$. $\square$
We now give the proof of Theorem 3.10.
Proof 6.24
Proof of Theorem 3.10. We prove each part of Theorem 3.10 separately.
(i) Consider the first case where $\theta_{1}=\theta_{2}=\cdots=\theta_{N}=\theta$ and $\underline{\lambda}_{i}>\theta e_{(3)}/(1-2e_{(3)}),\ \forall i\in[N].$ For any $i,j\in[N]\ (i\neq j)$, we have
$c_{(i)}\frac{\theta+\lambda_{i}}{\theta+2\lambda_{i}}+c_{(j)}\frac{\theta+\lambda_{j}}{\theta+2\lambda_{j}}-c_{(i,j)}$
$=$ $c_{(i)}+c_{(j)}-c_{(i)}\frac{\lambda_{i}}{\theta+2\lambda_{i}}-c_{(j)}\frac{\lambda_{j}}{\theta+2\lambda_{j}}-c_{(i,j)}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $c_{(i)}+c_{(j)}-c_{(i)}\frac{\underline{\lambda}_{i}}{\theta+2\underline{\lambda}_{i}}-c_{(j)}\frac{\underline{\lambda}_{j}}{\theta+2\underline{\lambda}_{j}}-c_{(i,j)}$
<$\stackrel{{\scriptstyle(b)}}{{<}}$ $\left(c_{(i)}+c_{(j)}\right)\left(1-e_{(3)}\right)-c_{(i,j)}$
$=$ $\left(c_{(i)}+c_{(j)}\right)\left(1-e_{(3)}-\frac{c_{(i,j)}}{c_{(i)}+c_{(j)}}\right)$
$=$ $\left(c_{(i)}+c_{(j)}\right)\left(e_{i,j}-e_{(3)}\right),$
where (a) and (b) are because the function $\lambda_{i}/(\theta+2\lambda_{i})$ is increasing in $\lambda_{i}>0$.
Denote
$\hat{j}(i)={\arg\max}_{j\in[N],j\neq i}e_{i,j},$
that is, $\hat{j}(i)$ is the “neighbor” of agent type $i$ with the highest matching efficiency (excluding agent type $i$ itself). Then based on the definition of $e_{(3)}$, $e_{i,j}>e_{(3)}\geq\max\{e_{i,(3)},e_{j,(3)}\}\ (i\neq j)$ is possible only if $j=\hat{j}(i)$ and $i=\hat{i}(j)$. Then according to Lemma 6.22, for $i,j\in[N],i\neq j$, $x_{i,j}^{*}>0$ or $x_{j,i}^{*}>0$ is possible only if $j=\hat{j}(i)$ and $i=\hat{i}(j)$. and for all other pairs $(i,j)$, we must have $x_{i,j}^{*}=x_{j,i}^{*}=0$. This result allows us to reduce program (2) into the following form:
$c(\boldsymbol{\lambda})=\min_{\boldsymbol{x},\boldsymbol{y}}\quad$ $\sum_{i,j\in[N],j=\hat{j}(i),i=\hat{i}(j)}c_{(i,j)}x_{i,j}+\sum_{i\in[N]}c_{(i)}\left(y_{i}+x_{i,i}\right),$
s.t. $\sum_{j\in[N],j=\hat{j}(i),i=\hat{i}(j)}\left(x_{i,j}+x_{j,i}\right)+x_{i,i}+y_{i}=\lambda_{i},$ $\forall i\in[N],$
[…]
This reduced formulation is further decomposable, since all $N$ types of agents can be formed into groups of at most 2, and the matching only happens within the group. Specifically, when pair $(i,j)$ satisfies $j=\hat{j}(i)$ and $i=\hat{i}(j)$, they might be matched together and thus form a group of 2. The remaining agent types (i.e., $i\neq\hat{i}(\hat{j}(i))$) will not be matched with any other agent types. Let $c_{i}(\lambda_{i})$ denote the optimal value function assuming only type-$i$ agent exists, then $c_{i}(\lambda_{i})$ can be given by Lemma 3.1 and is concave. Let $c_{i,j}(\lambda_{i},\lambda_{j})$ denote the optimal value function assuming only type-$i$ and type-$j$ agents exist, then $c_{i,j}(\lambda_{i},\lambda_{j})$ can be given by Corollary 3.8 and is also concave. By decomposition, we have
$c(\boldsymbol{\lambda})=\sum_{i,j\in[N],j=\hat{j}(i),i=\hat{i}(j)}c_{i,j}(\lambda_{i},\lambda_{j})+\sum_{i\in[N],i\neq\hat{i}(\hat{j}(i))}c_{i}(\lambda_{i}),$
and is thus concave.
(ii) For the second case, we have $\underline{\lambda}_{i}>\theta e_{(2)}/(1-2e_{(2)}),\ \forall i\in[N].$ Similar to the first case, for any $i,j\in[N]\ (i\neq j)$, we have
$c_{(i)}\frac{\theta+\lambda_{i}}{\theta+2\lambda_{i}}+c_{(j)}\frac{\theta+\lambda_{j}}{\theta+2\lambda_{j}}-c_{(i,j)}<\left(c_{(i)}+c_{(j)}\right)\left(e_{i,j}-e_{(2)}\right).$
This means that, according to Lemma 6.22, for any agent type $i\in[N]$, it can only be matched with itself. Thus the program (2) can be reduced into:
$c(\boldsymbol{\lambda})=\sum_{i\in[N]}c_{i}(\lambda_{i}),$
which is concave according to Lemma 3.1. $\square$
Before giving the proof of Theorem 3.11, we give another lemma. For $i\in[N]$, we denote $X_{i}=\{j|\theta_{i}x_{i,j}^{*}=\lambda_{j}y_{i}^{*}\}$.
Lemma 6.25
If
$\sum_{j\in X_{i}-\{i\}}\lambda_{j}\frac{\theta_{j}}{\theta_{j}+2\lambda_{j}}-\theta_{i}<0,$
we must have $i\in X_{i}$.
Proof 6.26
Proof. This lemma is a generalization of the second statements in Lemmas 6.12 and 6.14. Again, assume by contradiction that $i\notin X_{i}$ (i.e.,$\theta_{i}x_{i,i}^{*}<\lambda_{i}y_{i}^{*}$), then we must have $y_{i}^{*}>0$. Now we try to give another feasible solution with a smaller total cost. Given a small value of $\varepsilon>0$, let
$\hat{y}_{j}$ $=y_{j}^{*}+\varepsilon\lambda_{j}\frac{\theta_{j}}{\theta_{j}+2\lambda_{j}},$ $\forall j\in X_{i}-\{i\},$
$\hat{x}_{j,j}$ $=x_{j,j}^{*}+\varepsilon\lambda_{j}\frac{\lambda_{j}}{\theta_{j}+2\lambda_{j}},$ $\forall j\in X_{i}-\{i\},$
$\hat{x}_{i,j}$ $=x_{i,j}^{*}-\varepsilon\lambda_{j},$ $\forall j\in X_{i}-\{i\},$
$\hat{y}_{i}$ $=y_{i}^{*}-\varepsilon\theta_{i},$
$\hat{x}_{i,i}$ $=x_{i,i}^{*}+\frac{\varepsilon}{2}\left(\theta_{i}+\sum_{j\in X_{i}-\{i\}}\lambda_{j}\right),$
and keep the remaining variables unchanged. Similar to Lemmas 6.12 and 6.14, since $y_{i}^{*}>0$ and $x_{i,j}^{*}=\lambda_{j}y_{i}^{*}/\theta_{i}>0,\ \forall j\in X_{i}$, we can also guarantee that the new solution is still feasible. Then compared to the original solution, the difference in cost under this new feasible solution is given by
$\left[c_{(i)}(\hat{x}_{i,i}+\hat{y}_{i})+\sum_{j\in X_{i}-\{i\}}c_{(j)}(\hat{x}_{j,j}+\hat{y}_{j})+\sum_{j\in X_{i}-\{i\}}c_{(i,j)}(\hat{x}_{i,j}+\hat{x}_{j,i})\right]$
$-\left[c_{(i)}(x_{i,i}^{*}+y_{i}^{*})+\sum_{j\in X_{i}-\{i\}}c_{(j)}(x_{j,j}^{*}+y_{j}^{*})+\sum_{j\in X_{i}-\{i\}}c_{(i,j)}(x_{i,j}^{*}+x_{j,i}^{*})\right]$
$=$ $\varepsilon\left[\frac{c_{(i)}}{2}\left(\sum_{j\in X_{i}-\{i\}}\lambda_{j}-\theta_{i}\right)+\sum_{j\in X_{i}-\{i\}}\left(c_{(j)}\lambda_{j}\frac{\theta_{j}+\lambda_{j}}{\theta_{j}+2\lambda_{j}}-c_{(i,j)}\lambda_{j}\right)\right]$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $\varepsilon\left[\frac{c_{(i)}}{2}\left(\sum_{j\in X_{i}-\{i\}}\lambda_{j}-\theta_{i}\right)-\sum_{j\in X_{i}-\{i\}}\left(c_{(i,j)}\lambda_{j}\frac{\lambda_{j}}{\theta_{j}+2\lambda_{j}}\right)\right]$
$\stackrel{{\scriptstyle(b)}}{{\leq}}$ $\varepsilon\left[\frac{c_{(i)}}{2}\left(\sum_{j\in X_{i}-\{i\}}\lambda_{j}\frac{\theta_{j}}{\theta_{j}+2\lambda_{j}}-\theta_{i}\right)\right]$
$<$ $0,$
where (a) is due to $c_{(i,j)}\geq c_{(j)}$ and (b) is due to $c_{(i,j)}\geq c_{(i)}$. Therefore, we can obtain another feasible solution with a lower cost, leading to the contradiction. $\square$
Proof 6.27
Proof of Theorem 3.11. (i) Consider the first case where $\theta_{1}=\theta_{2}=\cdots=\theta_{N}=\theta$ and $\underline{\lambda}_{i}>\theta e_{(4)}/(1-2e_{(4)}),\ \forall i\in[N]$. Similar to the proof in Theorem 3.10, for any $i,j\in[N]\ (i\neq j)$, we have
$c_{(i)}\frac{\theta+\lambda_{i}}{\theta+2\lambda_{i}}+c_{(j)}\frac{\theta+\lambda_{j}}{\theta+2\lambda_{j}}-c_{(i,j)}<\left(c_{(i)}+c_{(j)}\right)\left(e_{i,j}-e_{(4)}\right).$
This means that, according Lemma 6.22, for a given type $i\in[N]$, the number of other agent types $j\in[N]$ satisfying $x_{i,j}^{*}>0$ or $x_{j,i}^{*}>0$ is at most 2.
Now, consider the following expression
$\sum_{j\in X_{i}-\{i\}}\lambda_{j}\frac{\theta}{\theta+2\lambda_{j}}-\theta=\theta\sum_{j\in X_{i}-\{i\}}\frac{\lambda_{j}}{\theta+2\lambda_{j}}-\theta\stackrel{{\scriptstyle(a)}}{{<}}0,$
where (a) is due to the size of the set $X_{i}-\{i\}$ is at most 2 and $\lambda_{j}/(\theta+2\lambda_{j})<1/2$. Then based on Lemma 6.25, we have $i\in X_{i}$, that is $\theta x_{i,i}^{*}=\lambda_{i}y_{i}^{*},\ \forall i\in[N]$. From constraint (2b), we have $x_{i,i}^{*}+y_{i}^{*}\leq\lambda_{i}$ and thus
$y_{i}^{*}\leq\frac{\lambda_{i}\theta}{\theta+2\lambda_{i}},\ \forall i\in[N].$ (10)
Then we show that $y_{i}^{*}>0$ for all $i\in[N]$. Assume by contradiction that there exists $i\in[N]$, $y_{i}^{*}=0$. Based on (2c), we have $x_{i,j}^{*}=0,\ \forall j\in[N]$. Then constraint (2b) for type $i$ leads to $\sum_{j\in[N],j\neq i}x_{j,i}^{*}=\lambda_{i}$. For $x_{j,i}^{*}$, we have
$x_{j,i}^{*}\leq\frac{\lambda_{i}y_{j}^{*}}{\theta}\stackrel{{\scriptstyle(a)}}{{\leq}}\lambda_{i}\frac{\lambda_{j}}{\theta+2\lambda_{j}}<\frac{\lambda_{i}}{2},$
where (a) is because of the inequality (10). Note that there are at most 2 agent types $j\neq i$ such that $x_{j,i}^{*}>0$, and if $x_{j,i}^{*}>0$, then we have $\sum_{j\in[N],j\neq i}x_{j,i}^{*}<2\lambda_{i}/2=\lambda_{i}$, which contradicts with constraint (2b).
Therefore, we must have $y_{i}^{*}>0$ for all $i\in[N]$. Based on Lemma 3.3, $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$.
(iii) Consider the second case where $\theta_{N}\leq 2\theta_{1}$ and $\underline{\lambda}_{i}>\theta_{i}e_{(3)}/(1-2e_{(3)}),\ \forall i\in[N]$. Similarly, we have
$c_{(i)}\frac{\theta_{i}+\lambda_{i}}{\theta_{i}+2\lambda_{i}}+c_{(j)}\frac{\theta_{i}+\lambda_{j}}{\theta_{i}+2\lambda_{j}}-c_{(i,j)}<\left(c_{(i)}+c_{(j)}\right)\left(e_{i,j}-e_{(3)}\right),$
thus for a given type $i\in[N]$, the number of other agent types $j\in[N]$ satisfying $x_{i,j}^{*}>0$ or $x_{j,i}^{*}>0$ is at most 1.
For any agent type $i\in[N]$, we have
$\sum_{j\in X_{i}-\{i\}}\lambda_{j}\frac{\theta_{j}}{\theta_{j}+2\lambda_{j}}-\theta_{i}\stackrel{{\scriptstyle(a)}}{{\leq}}\theta_{N}\sum_{j\in X_{i}-\{i\}}\frac{\lambda_{j}}{\theta_{j}+2\lambda_{j}}-\theta_{1}\stackrel{{\scriptstyle(b)}}{{<}}\frac{1}{2}\theta_{N}-\theta_{1}\leq 0,$
where (a) is due to the assumption that $\theta_{1}\leq\theta_{2}\leq\cdots\leq\theta_{N}$, and (b) is due to the size of the set $X_{i}-\{i\}$ is at most 1 and $\lambda_{j}/(\theta_{j}+2\lambda_{j})<1/2$. So again based on Lemma 6.25, we have $i\in X_{i}$. Repeating the remaining steps in the second case, we can prove that $y_{i}^{*}>0$ for all $i\in[N]$, and thus $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$. $\square$
7 Examples
Example 7.1
(Non-concavity of $c(\boldsymbol{\lambda})$ when $N=2,\theta_{1}\neq\theta_{2}$) Let $N=2$, $\theta_{1}=1$, $\theta_{2}=2$, $c_{(1)}=c_{(2)}=1$, and $c_{(1,2)}=c_{(2,1)}=1.01$. The function $c(\boldsymbol{\lambda})$ is non-concave at the point $\boldsymbol{\lambda}=(0.1,0.1)$—the function is smooth at this point but the largest eigenvalue is $0.03>0$.
Example 7.2
(Non-concavity of $c(\boldsymbol{\lambda})$ when $N=4$, $\theta_{1}=\theta_{2}=\theta_{3}=\theta_{4}$). Consider the case when $N=4,\theta_{1}=\cdots=\theta_{4}=1$. Let $[c_{(i,j)}]_{i,j\in[4]}$ be $\begin{bmatrix}0.70&0.77&0.83&0.92\\ 0.77&0.40&0.62&0.74\\ 0.83&0.62&0.50&0.86\\ 0.92&0.74&0.86&0.70\end{bmatrix}$. The function $c(\boldsymbol{\lambda})$ is non-concave at the point $\boldsymbol{\lambda}=(0.80,1.20,1.20,0.01)$—the function is smooth at this point but the largest eigenvalue is $0.04>0$.
Example 7.3
(Violation of weak concavity of $c(\boldsymbol{\lambda})$ when $N=2,\theta_{1}\neq\theta_{2}$) Let $N=2$, $\theta_{1}=1$, $\theta_{2}=8$, $c_{(1)}=c_{(2)}=1$, and $c_{(1,2)}=c_{(2,1)}=1.05$. According to Theorem 3.5, we have $\tau_{1}=4.2$ and $\tau_{2}=10.44$, and $c(\boldsymbol{\lambda})$ is weak concave when $\underline{\lambda}_{2}>18.57$. We can show that this bound is quite tight. For example, consider the point $(\lambda_{1},\lambda_{2})=(19.5,18.5)$, we can numerically calculate the left and right partial derivatives with respect to $\lambda_{1}$ at this point:
$\frac{\partial c(\boldsymbol{\lambda)}}{\partial\lambda_{1}^{-}}\approx 0.49986<\frac{\partial c(\boldsymbol{\lambda)}}{\partial\lambda_{1}^{+}}=0.5.$
For any $\rho>0$, a necessary condition for $c(\boldsymbol{\lambda)}-\rho\|\boldsymbol{\lambda}\|^{2}/2$ to be concave is that for any point $\boldsymbol{\lambda}\in\Lambda$, the left partial derivative is always larger than or equal to the right partial derivative (see e.g., Beckenbach 1948). However, the left partial derivative $c(\boldsymbol{\lambda)}-\rho\|\boldsymbol{\lambda}\|^{2}/2$ is always smaller than its right partial derivative in our example at $(\lambda_{1},\lambda_{2})$, since
$\frac{\partial c(\boldsymbol{\lambda)}-\rho\|\boldsymbol{\lambda}\|^{2}/2}{\partial\lambda_{1}^{-}}=\frac{\partial c(\boldsymbol{\lambda)}}{\partial\lambda_{1}^{-}}-\rho\lambda_{1}<\frac{\partial c(\boldsymbol{\lambda)}}{\partial\lambda_{1}^{+}}-\rho\lambda_{1}=\frac{\partial c(\boldsymbol{\lambda)}-\rho\|\boldsymbol{\lambda}\|^{2}/2}{\partial\lambda_{1}^{+}}.$
Therefore, in our example, $c(\boldsymbol{\lambda)}$ is not weakly concave.
Example 7.4
(Violation of weak concavity of $c(\boldsymbol{\lambda})$ when $N=2,\theta_{1}\neq\theta_{2}$ and $e_{1,2}=0.5$) Let $N=2$, $\theta_{1}=1$, $\theta_{2}=8$, $c_{(1)}=c_{(2)}=c_{(1,2)}=c_{(2,1)}=1$, then we have $e_{1,2}=0.5$ in this case, and $\tau_{1}=5>0$.
We can show that in this example, weak concavity is violated even for large values of $\lambda_{1}$ and $\lambda_{2}$, by again comparing left and right partial derivatives. Let $(\lambda_{1},\lambda_{2})=(11,10)$, we have
$\frac{\partial c(\boldsymbol{\lambda)}}{\partial\lambda_{1}^{-}}\approx 0.43574<\frac{\partial c(\boldsymbol{\lambda)}}{\partial\lambda_{1}^{+}}=0.5.$
Let $(\lambda_{1},\lambda_{2})=(101,100)$, we have
$\frac{\partial c(\boldsymbol{\lambda)}}{\partial\lambda_{1}^{-}}\approx 0.48837<\frac{\partial c(\boldsymbol{\lambda)}}{\partial\lambda_{1}^{+}}=0.5.$
Let $(\lambda_{1},\lambda_{2})=(1001,1000)$, we have
$\frac{\partial c(\boldsymbol{\lambda)}}{\partial\lambda_{1}^{-}}\approx 0.49876<\frac{\partial c(\boldsymbol{\lambda)}}{\partial\lambda_{1}^{+}}=0.5.$
Let $(\lambda_{1},\lambda_{2})=(10001,10000)$, we have
$\frac{\partial c(\boldsymbol{\lambda)}}{\partial\lambda_{1}^{-}}\approx 0.49988<\frac{\partial c(\boldsymbol{\lambda)}}{\partial\lambda_{1}^{+}}=0.5.$
Similar to Example 7.3, $c(\boldsymbol{\lambda})$ is not weakly concave at these points.
Example 7.5
(Multimodality of $g(\boldsymbol{\lambda})$) Let $N=2$, $\theta_{1}=\theta_{2}=0.3$, $c_{(1)}=c_{(2)}=1.1$, and $c_{(1,2)}=c_{(2,1)}=1.65$. The price-conversion function is given by $p_{i}(\lambda_{i})=1-\lambda_{i}$, $i=1,2$. Then $g(\boldsymbol{\lambda})$ is multimodal and not always differentiable, as shown in Figure 1.
Figure 1: Example showing the multimodality and non-differentiability of $g(\boldsymbol{\lambda})$ when $N=2$.