### v2

# Chen, Yan, and Jiang

[…]
Chen, Yan, and Jiang
[…]
On the LP for Dynamic Stochastic Matching and Its Application on Pricing
[…]
On the Linear Programming Model for Dynamic Stochastic Matching and Its Application on Pricing
[…]
Important pricing problems in centralized matching markets—such as carpooling and food delivery platforms—often exhibit a bi-level structure. At the upper level, the platform sets prices for different types of agents (e.g., riders with various origins and destinations, or delivery orders from diverse restaurants to customer locations). The lower level then matches converted agents to minimize operational costs, such as pooling riders with similar itineraries into the same vehicle or consolidating multiple delivery orders into a single trip. Motivated by these applications, we study the optimal value (cost) function of a linear programming model with respect to agent arrival rates, originally proposed by Aouad and Sarıtaç (2022) for cost-minimizing dynamic stochastic matching under limited time. In particular, we investigate the concavity of this cost function. We show that weak concavity holds when there are fewer than or equal to three agent types with identical patience level. For an arbitrary number of agent types, (weak) concavity is more likely to hold when agents have high or similar patience, or when agents are more heterogeneous in their matching efficiencies. Building on these theoretical insights, we develop a Minorize-Maximization (MM) algorithm for the pricing problem that requires little stepsize tuning, and demonstrate substantial performance improvements over projected gradient-based methods on a large-scale, real-world ridesharing dataset. This makes it a compelling algorithmic choice for solving such pricing problems in practice.
[…]
dynamic stochastic matching, pricing, fluid relaxation, weak concavity, minorize-maximization
1 Introduction
Centralized platforms that perform dynamic, non-bipartite matching are increasingly prevalent in modern marketplaces, where different types of agents arrive stochastically and can be matched with other agents within a limited time. A representative example is shared ride services offered by ridesharing platforms such as Uber, Lyft and Didi, which match passengers with similar origin-destination (OD) pairs into shared trips. Due to limited patience, unmatched passengers may eventually be served individually (Yan, Yan, and Shen 2025). Another example is the online food delivery platform with batching decisions, where two or more orders with similar customer locations can be matched together and delivered by the same courier (Ma, Ma, and Romero 2025).
This work is motivated by the pricing challenges faced by such platforms—for example, in setting upfront fares for shared rides, where passengers have different ODs. Pricing and matching are inherently interdependent (see, e.g., Feng, Niazadeh, and Saberi 2024), naturally forming a bi-level optimization problem: at the upper level, the platform determines optimal prices to maximize profit, often expressed as a concave revenue component minus a total cost component; whereas at the lower level, the platform matches agents to minimize total costs. For a detailed discussion of this shared rides pricing application, see Yan, Yan, and Shen (2025).
Problem Formulation.
Consider a system with $N$ possible agent types, where $[N]$ denotes the set ${1,2,\ldots,N}$. Potential agents of type $i\in[N]$ observe a quoted price $p_{i}$ and decide whether to enter the platform. Each type-$i$ agent arrives according to an independent Poisson process with arrival rate $\lambda_{i}>0$, which depends on the quoted price. Once admitted, a type-$i$ agent’s sojourn time in the system is exponentially distributed with rate $\theta_{i}$. Here, $\theta_{i}$ characterizes the patience of type-$i$ agents—a smaller $\theta_{i}$ indicates a greater willingness to wait for a match.
The platform’s objective is to maximize total profit by optimally setting prices and conducting matching for all agents. Following common assumptions in the literature, we assume that for each agent type $i\in[N]$, there is a one-to-one correspondence between the arrival rate $\lambda_{i}$ and the quoted price $p_{i}$, allowing us to express the price as a function of the arrival rate: $p_{i}(\lambda_{i})$. We further assume that the revenue function, $\lambda_{i}p_{i}(\lambda_{i})$, is concave in $\lambda_{i}$. Thus, the pricing problem reduces to finding the optimal vector $\boldsymbol{\lambda}=(\lambda_{i})_{i\in[N]}$ that maximizes total profit, formulated as
$\max_{\boldsymbol{\lambda}\in\Lambda}~{}g(\boldsymbol{\lambda})=\max_{\boldsymbol{\lambda}\in\Lambda}{\sum_{i\in[N]}~{}\lambda_{i}p_{i}(\lambda_{i})-c(\boldsymbol{\lambda})},$ (1)
where $\Lambda$ denotes the feasible domain of $\boldsymbol{\lambda}$, and $c(\boldsymbol{\lambda})$ is the optimal value (cost) function of the lower-level matching problem.
The lower-level matching problem has received a lot of attention in the literature recently. Existing studies often formulate it as a Markov Decision Process (MDP), under either a continuous-time setting (Ashlagi et al. 2019, Aouad and Sarıtaç 2022) or a discrete-time setting (Ashlagi et al. 2023, Gupta 2024, Eom and Toriello 2024, Kerimov, Ashlagi, and Gurvich 2024, 2025). In the MDP framework, the state of the system is defined as the number of agents of each type waiting in the system to be matched. When an agent arrives, the platform can (i) immediately match her with another existing agent, or (ii) keep her waiting in the system. However, the optimal policies are in general intractable to obtain due to the curse of dimensionality. Therefore, many studies use fluid relaxations based on linear programs (LPs) to give bounds for the optimal value function. For example, for the dynamic non-bipartite matching problem, Eom and Toriello (2024), Gupta (2024) and Kerimov, Ashlagi, and Gurvich (2024, 2025) give reward-maximizing LP formulations under infinite agent patience. Collina et al. (2020) and Aouad and Sarıtaç (2022) formulate LPs incorporating limited patience. In this work, we adopt the cost-minimizing LP by Aouad and Sarıtaç (2022) to model $c(\boldsymbol{\lambda})$.
While existing studies focus on developing near-optimal matching policies using these LPs, our contribution lies in understanding how the optimal cost function $c(\boldsymbol{\lambda})$ varies as a function of the arrival rate, which is a crucial aspect of solving the pricing problem (1). This, to our best knowledge, has not been covered in existing studies where arrival rates are typically treated as fixed inputs.
With infinite patience (as in the LP formulation of Eom and Toriello 2024), it can be shown that $c(\boldsymbol{\lambda})$ is convex, making the pricing problem (1) a well-behaved concave maximization. However, when limited patience is taken into account, characterizing $c(\boldsymbol{\lambda})$ becomes highly nontrivial. As we will show in Section 2, $\boldsymbol{\lambda}$ appears in both the coefficients and the right-hand sides of the LP constraints, departing from known results in classical parametric LP literature (Gal and Nedoma 1972, Adler and Monteiro 1992). This makes the pricing problem (1) non-convex in general under limited agent patience.
Interestingly, in sharp contrast to the convexity of the optimal cost function under infinite agent patience, we find that $c(\boldsymbol{\lambda})$ often exhibits weak concavity when agent patience is limited. In particular, we show that weak concavity holds when there are fewer than or equal to three agent types with identical patience level. For an arbitrary number of agent types, (weak) concavity is more likely to hold when agents have high or similar patience, or when agents are more heterogeneous in their matching efficiencies.
Motivated by this observation, we develop a Minorize-Maximization (MM) algorithm (Hunter and Lange 2004), which provides convergence guarantees for concave-minus-concave maximization problems. Specifically, in program (1), the revenue term $\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})$ is concave, and when $c(\boldsymbol{\lambda})$ also exhibits concave properties, the MM algorithm can be effectively applied. Our proposed MM algorithm requires little stepsize tuning, making it easy to implement. A case study using the Chicago ridesharing dataset demonstrates that the MM algorithm significantly outperforms the projected gradient-based method, achieving both shorter running times and better optimality. Moreover, while the performance of the projected gradient method is highly sensitive to step-size tuning, the MM algorithm delivers consistently robust results, making it a compelling choice for solving such pricing problems in practice.
Organization. The remainder of the paper is organized as follows. We introduce the model and preliminaries in Section 2. We perform concave analysis on the value function in Section 3 and discuss its application on price optimization in Section 4. We conclude in Section 5. All proofs, auxiliary results, and additional examples are provided in the appendix.
2 Preliminaries
First, we introduce some additional terminology and notation. Following Aouad and Sarıtaç (2022), when two agents are matched, we refer to the agent who arrives earlier as the active agent and the one who arrives later as the passive agent. When an active agent of type $i\in[N]$ is matched with a passive agent of type $j\in[N]$, a cost $c_{(i,j)}>0$ is incurred. If an agent of type $i\in[N]$ remains unmatched when her patience runs out, a cost $c_{(i)}>0$ is incurred. Consistent with Yan, Yan, and Shen (2025), we make the following assumptions about these costs. {assumption} $c_{(i)}=c_{(i,i)}\leq c_{(i,j)}=c_{(j,i)},\ \forall i,j\in[N]$.
Assumption 2 states that when two agents of the same type $i\in[N]$ are matched, the resulting cost equals that of serving a single unmatched agent of type $i$. In the context of ridesharing, this means that matching two identical trips incurs the same cost as serving either trip individually. The condition $c_{(i,j)}=c_{(j,i)}$ imposes symmetry and requires that the order of arrival—that is, which agent is active or passive—does not affect the matched cost. Additionally, $c_{(i,j)}\geq c_{(i)}$ requires that the cost of serving two matched agents is at least as great as serving a single agent alone. This is a natural assumption in ridesharing, where a pooled trip must cover at least the distance required by one of the solo trips.
We also define the matching efficiency between two types of agents as follows.
Definition 2.1
The matching efficiency between agent types $i\in[N]$ and $j\in[N]$ is defined as $e_{i,j}=1-c_{(i,j)}/(c_{(i)}+c_{(j)})$.
Matching efficiency quantifies the proportion of cost saved when two types of agents are matched. A low matching efficiency indicates greater heterogeneity between the agents—for example, when their OD pairs differ significantly in ridesharing. By Assumption 2, $c_{(i,j)}\geq c_{(i)}$ and $c_{(i,j)}\geq c_{(j)}$, so $e_{i,j}$ is bounded above by $0.5$. Additionally, since $c_{(i,i)}=c_{(i)}$, we have $e_{i,i}=0.5$ for all $i\in[N]$, implying that self-matching always yields the highest efficiency. The lower bound of $e_{i,j}$ can be negative when $c_{(i,j)}>c_{(i)}+c_{(j)}$, meaning that matching these two types of agents would result in higher costs than serving them individually—rendering such matches impractical in any scenario.
The dynamic stochastic matching problem can be formulated as an MDP. In our study, instead of directly focusing on solving the MDP, we analyze the following LP:
$c(\boldsymbol{\lambda})=\min_{\boldsymbol{x},\boldsymbol{y}}\quad$ $\sum_{i\in[N]}\sum_{j\in[N]}c_{(i,j)}x_{i,j}+\sum_{i\in[N]}c_{(i)}y_{i}$ (2a)
s.t. $\sum_{j\in[N]}x_{j,i}+\sum_{j\in[N]}x_{i,j}+y_{i}=\lambda_{i},$ $\forall i\in[N],$ (2b)
$\theta_{i}x_{i,j}\leq\lambda_{j}y_{i},$ $\forall i,j\in[N],$ (2c)
$x_{i,j}\geq 0,$ $\forall i,j\in[N],$ (2d)
$y_{i}\geq 0,$ $\forall i\in[N],$ (2e)
where $\boldsymbol{\lambda}:=(\lambda_{i})_{i\in[N]}$, $\boldsymbol{x}:=(x_{i,j})_{i,j\in[N]}$, and $\boldsymbol{y}:=(y_{i})_{i\in[N]}$.
This LP model is often called the fluid relaxation of the MDP model. Variable $x_{i,j}$ represents the average match rate between type $i\in[N]$ active agent and type $j\in[N]$ passive agent, and $y_{i}$ represents the rate of type $i\in[N]$ agent unmatched. The objective function (2a) is the average total cost per unit of time. Constraints (2b) can be viewed as the flow balance equation, that is, in the steady state, the arrival rate of type-$i$ agent should be the summation of its active match rate, passive match rate, and unmatched rate. Constraints (2c) give the lower bound of the unmatched rate $y_{i}$ in terms of $\lambda_{j}$, $x_{i,j}$, and $\theta_{i}$. Intuitively, this reflects that a type $i$-th agent may leave the system before the next type $j$-th agent arrives, with a probability of $\theta_{i}/(\theta_{i}+\lambda_{j})$, thus $y_{i}$ should be a factor of at least $\theta_{i}/\lambda_{j}$ of the match rate $x_{i,j}$. Lastly, (2d) and (2e) require that the variables are nonnegative.
This LP was originally introduced in Aouad and Sarıtaç (2022), where it was shown that its optimal value provides a lower bound on the expected cost under the optimal matching policy. Correspondingly, using this LP, the optimal profit given in (1) provides an upper bound for the profit achievable by any dynamic matching policy. While Aouad and Sarıtaç (2022) treat $\boldsymbol{\lambda}$ as exogenously given, we instead view this LP as a parametric programming model in $\boldsymbol{\lambda}$. Formally, we consider $c(\boldsymbol{\lambda})$ as a function of $\boldsymbol{\lambda}$ defined over a compact interval $\Lambda=[\underline{\boldsymbol{\lambda}},\overline{\boldsymbol{\lambda}}]$. These lower and upper bounds can be interpreted as minimum and maximum prices imposed on different agent types. For convenience, we assume $\underline{\boldsymbol{\lambda}}>\boldsymbol{0}$, where all vector inequalities in this paper are interpreted component-wise.
{assumption}
$\boldsymbol{\lambda}\in\Lambda=[\underline{\boldsymbol{\lambda}},\overline{\boldsymbol{\lambda}}]$, where $\underline{\boldsymbol{\lambda}}>\boldsymbol{0}$.
We first observe that, as an extreme case, when all agents have infinite patience (i.e., $\theta_{i}=0$ for all $i\in[N]$), constraints (2c) become redundant, and the resulting LP closely resembles the formulation in Eom and Toriello (2024). In this scenario, $\boldsymbol{\lambda}$ appears only on the right-hand side of constraint (2b), and it is well known that $c(\boldsymbol{\lambda})$ becomes a convex, piecewise linear function (Adler and Monteiro 1992). Moreover, under Assumption 2, we can show that $c(\boldsymbol{\lambda})$ is, in fact, linear.
Proposition 2.2
When $\theta_{i}=0$, $\forall i\in[N]$, $c(\boldsymbol{\lambda})$ is linear on $\Lambda$.
However, when constraint (2c) is imposed, analyzing $c(\boldsymbol{\lambda})$ becomes significantly more challenging. In this case, $\boldsymbol{\lambda}$ also appears in the coefficients of constraint (2c), making the LP fundamentally different from the classical parametric LP literature, where the parameter appears only on the right-hand side of the constraints (Gal and Nedoma 1972, Adler and Monteiro 1992). In fact, as shown in the following proposition, the cost function exhibits economies of scale, implying that $c(\boldsymbol{\lambda})$ tends to be concave rather than convex.
Proposition 2.3
[…]
Proposition 2.3 states that the average cost per agent decreases when the arrival rates are scaled proportionally. This implies that $c(\boldsymbol{\lambda})$ tends to bend downward (or remain linear) as the arrival rate increases—a behavior typically associated with concavity rather than convexity. However, since $c(\boldsymbol{\lambda})$ is a multivariate function, establishing concavity requires a more rigorous analysis.
3 Concavity of Value Function
In this section, we examine the concavity of $c(\boldsymbol{\lambda})$. We start with the simplest case when there is only one type of agent (i.e., $N=1$). It is straightforward to derive the explicit solution and show that the value function is concave under this case. Note that the optimal solutions to (2) depend on $\boldsymbol{\lambda}$ and can be written as $\boldsymbol{x}^{*}(\boldsymbol{\lambda})=(x_{i,j}^{*}(\boldsymbol{\lambda}))_{i,j\in[N]}$ and $\boldsymbol{y}^{*}(\boldsymbol{\lambda})=(y_{i}^{*}(\boldsymbol{\lambda}))_{i\in[N]}$, but for notational simplicity, we just use the abbreviations $\boldsymbol{x}^{*}=(x_{i,j}^{*})_{i,j\in[N]}$ and $\boldsymbol{y}^{*}=(y_{i}^{*})_{i\in[N]}$ throughout this section.
Lemma 3.1
When $N=1$, the optimal solution to (2) is given by $y_{1}^{*}=\lambda_{1}\theta_{1}/(\theta_{1}+2\lambda_{1})$ and $x_{1,1}^{*}=\lambda_{1}^{2}/(\theta_{1}+2\lambda_{1})$. The value function $c(\lambda_{1})$ is concave.
To prove Lemma 3.1, it suffices to show that the inequality in (2c) is always binding at optimality, implying that the platform should match as much as possible. While Lemma 3.1 establishes concavity in the simple case of $N=1$, this result does not directly extend to larger $N$. When $N\geq 2$, different agent types interact, and the total cost involves a sum of heterogeneous pairwise terms $(c_{(i,j)})_{i,j\in[N]}$, making the structure of $c(\boldsymbol{\lambda})$ much more complex—and potentially non-smooth. This complexity invalidates the straightforward concavity observed for $N=1$. For explicit counterexamples of non-concavity, see Example 7.1 (for $N=2$ with $\theta_{1}\neq\theta_{2}$) and Example 7.2 (for $N>2$ with $\theta_{1}=\theta_{2}=\cdots=\theta_{N}$), both in Appendix 7.
Given these counterexamples for $N\geq 2$, it may be unrealistic to rely solely on the standard definition of concavity. Therefore, we also consider the notion of weak concavity, a relaxed alternative defined below. See also Proposition 4.3 in Vial (1983) for an equivalent definition.
Definition 3.2
(Weak concavity) $c(\boldsymbol{\lambda})$ is said to be weakly concave on $\Lambda$ if there exists a finite value $\rho\geq 0$, such that $c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2$ is concave on $\Lambda$.
If $c(\boldsymbol{\lambda})$ is weakly concave, it can be written as the sum of a concave function, $c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2$, and a convex function, $\rho\|\boldsymbol{\lambda}\|^{2}/2$. This decomposition allows us to express the objective as $g(\boldsymbol{\lambda})=(\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})-\rho\|\boldsymbol{\lambda}\|^{2}/2)-(c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2)$, thereby restoring a concave-minus-concave structure that is useful for algorithm design in Section 4. We also emphasize that the weakly concave property is nontrivial, even for bounded functions; for example, one can verify that $|x|$ (for $-1\leq x\leq 1$) and $-\sqrt{x}$ (for $0\leq x\leq 1$) are not weakly concave.
We now state a fundamental lemma, which provides a sufficient condition under which the value function $c(\boldsymbol{\lambda})$ is weakly concave.
Lemma 3.3
If $\forall\boldsymbol{\lambda}\in\Lambda$, all optimal solutions satisfy $\boldsymbol{y}^{*}>\boldsymbol{0}$, then $c(\boldsymbol{\lambda})$ is weakly concave on $\Lambda$.
When proving Lemma 3.3, we first establish the connection between global weak concavity and local weak concavity (see Lemma 6.1 in Appendix 6.4), that is, it is sufficient for us to prove that $c(\boldsymbol{\lambda})$ is weakly concave in a small neighborhood around any point $\hat{\boldsymbol{\lambda}}\in\Lambda$. To prove the local weak concavity, we examine the optimal basic feasible solution (BFS) of program (2) at $\hat{\boldsymbol{\lambda}}$ (Bertsimas and Tsitsiklis 1997). Using the assumption that $\boldsymbol{y}^{*}>\boldsymbol{0}$, we can prove that any optimal BFS still produces a basic feasible solution within a sufficiently small neighborhood around $\hat{\boldsymbol{\lambda}}$, and the corresponding value function remains both bounded and weakly concave. Using the fact that $c(\boldsymbol{\lambda})$ can be seen as a minimization of all candidate BFS, this local concavity property extends to $c(\boldsymbol{\lambda})$ itself.
Given Lemma 3.3, to establish the weak concavity of $c(\boldsymbol{\lambda})$, it suffices to identify conditions under which $\boldsymbol{y}^{*}>\boldsymbol{0}$. This allows us to bypass the complexity and potential non-smoothness of $c(\boldsymbol{\lambda})$ when $N\geq 2$. Moreover, restricting our attention to the case where $\boldsymbol{y}^{*}>\boldsymbol{0}$ is also practically justifiable—the case $y_{i}^{*}=0$ is an artifact of the LP relaxation and cannot be realized in the original MDP framework, as the unmatched rate should never be exactly zero when agents have limited patience (i.e., $\theta_{i}>0$ for all $i\in[N]$).
Leveraging Lemma 3.3, we present results for the case $N=2$ in Section 3.1, and extend to the general case for arbitrary $N$ in Section 3.2.
3.1 Analysis of Two Types of Agents
In this subsection, we examine the case where $N=2$. We show that the optimal solution to the LP can be fully characterized, and that weak concavity can be established by identifying the region where $\boldsymbol{y}^{*}>\boldsymbol{0}$. Without loss of generality, we assume $\theta_{1}\leq\theta_{2}$ throughout this subsection.
It is well known that whenever an LP admits an optimal solution, there always exists an optimal BFS (Bertsimas and Tsitsiklis 1997). In any such optimal BFS, a subset of the inequality constraints—specifically, those in (2c), (2d), and (2e)—will be binding. We leverage this fact to characterize the structure of optimal solutions, that is, to determine which constraints are binding for different values of $\lambda_{1}$ and $\lambda_{2}$. For ease of exposition, we define:
$\Delta_{1}(\lambda_{1},\lambda_{2})$ $:=c_{(1)}\frac{\theta_{1}+\lambda_{1}}{\theta_{1}+2\lambda_{1}}+c_{(2)}\frac{\theta_{2}+\lambda_{2}}{\theta_{2}+2\lambda_{2}}-c_{(1,2)},$
$\Delta_{2}(\lambda_{1},\lambda_{2})$ $:=\frac{c_{(1)}}{2}\left[1-\frac{\theta_{1}(\theta_{2}+\lambda_{1}+2\lambda_{2})}{\lambda_{2}(\theta_{2}+2\lambda_{2})}\right]+c_{(2)}\frac{\theta_{2}+\lambda_{2}}{\theta_{2}+2\lambda_{2}}-c_{(1,2)},$
$\Delta_{3}(\lambda_{1},\lambda_{2})$ $:=\lambda_{1}-\lambda_{2}-\theta_{1}.$
Then the following proposition describes the optimal solution to (2).
Lemma 3.4
When $N=2$, an optimal solution to (2) is characterized by Table 1. In addition,
(i) all optimal solutions to (2) satisfy $y_{1}^{*}>0$;
(ii) all optimal solutions to (2) satisfy $y_{2}^{*}>0$ if and only if $\Delta_{2}(\lambda_{1},\lambda_{2})<0$ or $\Delta_{3}(\lambda_{1},\lambda_{2})<0$.
Table 1: Characterization of optimal solutions to program (2) when $N=2$.
Case Conditions Optimal Solutions
$\Delta_{1}(\lambda_{1},\lambda_{2})$ $\Delta_{2}(\lambda_{1},\lambda_{2})$ $\Delta_{3}(\lambda_{1},\lambda_{2})$ $y_{1}^{*}$ $y_{2}^{*}$ $x_{1,1}^{*}$ $x_{1,2}^{*}$ $x_{2,1}^{*}$ $x_{2,2}^{*}$
(i) $\leq 0$ $>0$ $>0$ $=\lambda_{1}y_{1}^{*}/\theta_{1}$ $=0$ $=0$ $=\lambda_{2}y_{2}^{*}/\theta_{2}$
(ii) $>0$ $<0$ $>0$ $>0$ $=\lambda_{1}y_{1}^{*}/\theta_{1}$ $=\lambda_{2}y_{1}^{*}/\theta_{1}$ $=\lambda_{1}y_{2}^{*}/\theta_{2}$ $=\lambda_{2}y_{2}^{*}/\theta_{2}$
(iii) $\geq 0$ $<0$ $>0$ $>0$ $=0$ $=\lambda_{2}y_{1}^{*}/\theta_{1}$ $=\lambda_{1}y_{2}^{*}/\theta_{2}$ $=\lambda_{2}y_{2}^{*}/\theta_{2}$
(iv) $\geq 0$ $\geq 0$ $>0$ $=0$ $=(\lambda_{1}-y_{1}^{*}-x_{1,2}^{*})/2$ $=\lambda_{2}y_{1}^{*}/\theta_{1}$ $=0$ $=0$
Lemma 3.4 identifies four possible types of optimal solutions, determined by the signs of three indicator functions. When $x_{i,j}^{*}=0$, the corresponding constraint in (2d) is binding; when $x_{i,j}^{\ast}=\lambda_{j}y_{i}^{\ast}/\theta_{i}$, the corresponding constraint in (2c) is binding. The first solution category corresponds to the case with no cross-matching between type 1 and type 2 agents (i.e., $x_{1,2}^{\ast}=x_{2,1}^{\ast}=0$). The second category is the fully-matched case, where all types of matchings occur and all constraints in (2c) are binding. In the third category, there is no self-matching among type 1 agents ($x_{1,1}^{\ast}=0$). The last category is characterized by $y_{2}^{\ast}=0$, where all type 2 agents (with low patience) are passively matched with type 1 agents (with high patience). Using Lemma 3.4 and flow constraints (2b), we can explicitly give the expression of $y_{1}^{*}$ and $y_{2}^{*}$, thereby obtaining the optimal solutions in closed forms. The complete closed-form expressions are provided at the end of Appendix 6.4. Lemma 3.4 also helps us to identify the region where $c(\boldsymbol{\lambda})$ is weakly concave. Based on the results of Lemma 3.3 and Lemma 3.4, we see that if, for every $(\lambda_{1},\lambda_{2})\in\Lambda$, at least one of the inequalities $\Delta_{2}(\lambda_{1},\lambda_{2})<0$ or $\Delta_{3}(\lambda_{1},\lambda_{2})<0$ holds, then $y_{1}^{\ast},y_{2}^{\ast}>0$ is guaranteed. The following theorem further explores this condition and gives an explicit characterization of $\Lambda$ satisfying such conditions.
Theorem 3.5
Define
$\tau_{1}$ $:=c_{(1)}(\theta_{2}-3\theta_{1})+2c_{(2)}\theta_{2}-2c_{(1,2)}\theta_{2},$
$\tau_{2}$ $:=\tau_{1}^{2}-8c_{(1)}(2c_{(1,2)}-c_{(1)}-c_{(2)})\theta_{1}(\theta_{1}+\theta_{2}),$
then $c(\boldsymbol{\lambda})$ is weakly concave on $\Lambda$ under any of the following conditions:
(i) $\tau_{1}\leq 0$ or $\tau_{2}<0$;
(ii) $\tau_{1}>0$, $\tau_{2}\geq 0$, $2c_{(1,2)}-c_{(1)}-c_{(2)}\neq 0$ and $\underline{\lambda}_{2}>(\tau_{1}+\sqrt{\tau_{2}})/\left[4(2c_{(1,2)}-c_{(1)}-c_{(2)})\right];$
(iii) $\tau_{1}>0$, $\tau_{2}\geq 0$, and $\underline{\lambda}_{2}>\overline{\lambda}_{1}-\theta_{1}$.
Theorem 3.5 gives three sufficient conditions for weak concavity, depending on the signs of $\tau_{1}$ and $\tau_{2}$. For the first case when $\tau_{1}\leq 0$ or $\tau_{2}<0$, $c(\boldsymbol{\lambda})$ is weakly concave on any closed interval $\Lambda$. For the second case, we require a lower bound of $\underline{\lambda}_{2}$, and for the third case, we pose restrictions on both $\underline{\lambda}_{2}$ and $\overline{\lambda}_{1}$. In general, smaller values of $\tau_{1}$ and $\tau_{2}$ are desired to achieve weak concavity. Since both $\tau_{1}$ and $\tau_{2}$ are decreasing in $\theta_{1}$, a larger value of $\theta_{1}$ (i.e., when $\theta_{1}$ and $\theta_{2}$ are closer since $\theta_{1}\leq\theta_{2}$) helps to achieve weak concavity in the larger region. Similarly, since both $\tau_{1}$ and $\tau_{2}$ are decreasing in $c_{(1,2)}$, a larger value of $c_{(1,2)}$ (i.e., lower matching efficiency) also promotes weak concavity. Although Theorem 3.5 provides only sufficient conditions, we find that the resulting bounds are quite tight. In Appendix 7, we also present a counterexample (Example 7.3) demonstrating that as soon as $\lambda_{2}$ falls just below our derived bound in the second case, weak concavity is already violated.
Also note that in the second case of Theorem 3.5, we require $2c_{(1,2)}-c_{(1)}-c_{(2)}\neq 0$, or equivalently, $e_{1,2}\neq 0.5$. This condition means that the two agent types cannot have perfect matching efficiency. In the context of ridesharing, this essentially precludes the possibility that both rider types have exactly the same OD pairs. When $e_{1,2}=0.5$ and $\theta_{1}=\theta_{2}$, this is not an issue, as the platform can simply merge the two types into a single type. However, as the following proposition shows, if $\theta_{1}\neq\theta_{2}$, the region where $y_{2}^{*}=0$ becomes unbounded, making it impossible to establish any lower bound on $\underline{\boldsymbol{\lambda}}$ that guarantees $\boldsymbol{y}^{*}>\boldsymbol{0}$, without restricting $\overline{\boldsymbol{\lambda}}$. In Appendix 7, we also provide a counterexample (Example 7.4) demonstrating that weak concavity can be violated when $e_{1,2}=0.5$ and $\theta_{1}\neq\theta_{2}$, even for large values of $\lambda_{1}$ and $\lambda_{2}$.
Proposition 3.6
When $\tau_{1}>0$ and $02c_{(1,2)}-c_{(1)}-c_{(2)}=$ 0, given any $\underline{\boldsymbol{\lambda}}>\boldsymbol{0}$, there exists $\boldsymbol{\lambda}>\underline{\boldsymbol{\lambda}}$ such that $y_{2}^{*}=0$.
Given the complexity of the expressions for $\tau_{1}$ and $\tau_{2}$ in Theorem 3.5, we also provide the following corollary, which presents a more tractable sufficient condition in a simplified form.
Corollary 3.7
$c(\boldsymbol{\lambda})$ is weakly concave on $\Lambda$ under any of the following conditions:
(i) $\theta_{2}/\theta_{1}<3$;
(ii) $\theta_{2}/\theta_{1}\geq 3$, $e_{1,2}\neq 0.5$ and $\underline{\lambda}_{2}>\theta_{2}/\left[4(1-2e_{1,2})\right]$;
(iii) $\theta_{2}/\theta_{1}\geq 3$, and $\overline{\lambda}_{1}\leq\theta_{1}$.
According to Corollary 3.7, the function is guaranteed to be weakly concave as long as $\theta_{2}<3\theta_{1}$. If the two agent types have significantly different patience levels, we require a lower bound on $\underline{\lambda}_{2}$ or an upper bound on $\overline{\lambda}_{1}$. The lower bound of $\underline{\lambda}_{2}$ depends on $\theta_{2}$ (the patience of agent type 2) and $e_{1,2}$ (the matching efficiency between the two agent types). A smaller $\theta_{2}$ (greater patience for both types of agents since $\theta_{1}\leq\theta_{2}$) and a lower $e_{1,2}$ (lower matching efficiency) both expand the region of $\Lambda$ where weak concavity is maintained. The upper bound of $\overline{\lambda}_{1}$ depends on $\theta_{1}$, the patience of agent type 1, and when $\theta_{1}$ is larger (i.e., when $\theta_{1}$ and $\theta_{2}$ are closer since $\theta_{1}\leq\theta_{2}$), weak concavity is easier to achieve.
3.1.1 Special case of same patience level.
We now consider a special case where two types of agents have the same patience level (that is, $\theta_{1}=\theta_{2}=\theta$). Under this condition, the structure of optimal solutions simplifies considerably compared to the general case described in Lemma 3.4, as stated in the following corollary.
Corollary 3.8
When $N=2$ and $\theta_{1}=\theta_{2}=\theta$, an optimal solution to program (2) is characterized by Table 2.
Table 2: Characterization of optimal solutions to program (2) when $N=2$ and $\theta_{1}=\theta_{2}=\theta$.
Case $\Delta_{1}(\lambda_{1},\lambda_{2})$ Optimal Solutions
$y_{1}^{*}$ $y_{2}^{*}$ $x_{1,1}^{*}$ $x_{1,2}^{*}$ $x_{2,1}^{*}$ $x_{2,2}^{*}$
(i) $\leq 0$ $=\lambda_{1}\theta/(\theta+2\lambda_{1})$ $=\lambda_{2}\theta/(\theta+2\lambda_{2})$ $=\lambda_{1}y_{1}^{*}/\theta$ $=0$ $=0$ $=\lambda_{2}y_{2}^{*}/\theta$
(ii) $>0$ $=\lambda_{1}\theta/(\theta+2\lambda_{1}+2\lambda_{2})$ $=\lambda_{2}\theta/(\theta+2\lambda_{1}+2\lambda_{2})$ $=\lambda_{1}y_{1}^{*}/\theta$ $=\lambda_{2}y_{1}^{*}/\theta$ $=\lambda_{1}y_{2}^{*}/\theta$ $=\lambda_{2}y_{2}^{*}/\theta$
Corollary 3.8 states that when $\theta_{1}=\theta_{2}$, there are only two possible categories of optimal solutions, determined solely by the sign of $\Delta_{1}(\lambda_{1},\lambda_{2})$. In this special case, the last two solution categories described in Lemma 3.4 no longer arise. Intuitively, when $c_{(1,2)}$ is much larger than $c_{(1)}$ and $c_{(2)}$, the optimal strategy is to match agents only within their own types; otherwise, a mix of self- and cross-matching becomes optimal. The arrival rate and patience levels also influence the solution: higher arrival rates (i.e., larger $\lambda_{1}$ and $\lambda_{2}$) or greater patience (i.e., smaller $\theta$) both tend to reduce the extent of cross-matching.
Since $\boldsymbol{y}^{*}>\boldsymbol{0}$ always holds when $\theta_{1}=\theta_{2}$, the weak concavity of $c(\boldsymbol{\lambda})$ follows directly from Lemma 3.3. Moreover, using the explicit form of the optimal solution given by Corollary 3.8, we can further verify that $c(\boldsymbol{\lambda})$ is, in fact, concave.
Theorem 3.9
When $N=2$ and $\theta_{1}=\theta_{2}=\theta$, $c(\boldsymbol{\lambda})$ is concave on $\Lambda$.
3.2 Analysis of Multiple Types of Agents
We now present results for the general case with an arbitrary number of agent types. As $N$ increases, fully characterizing the optimal solutions becomes intractable. However, by leveraging insights from previous results and their proofs, we can identify sufficient conditions for ensuring $\boldsymbol{y}^{*}>\boldsymbol{0}$ for general $N$, and once again apply Lemma 3.3 to establish weak concavity. For clarity, we assume throughout this subsection that $\theta_{1}\leq\theta_{2}\leq\cdots\leq\theta_{N}$. We also define $\nu:=\theta_{N}/\theta_{1}\geq 1$, that is, a larger $\nu$ represents a larger difference in agents’ patience levels.
Given any type $i$, we sort its corresponding matching efficiency $e_{i,1},e_{i,2},\cdots,e_{i,N}$ in descending order and denote the sorted result as $e_{i,(1)}\geq e_{i,(2)}\geq\cdots\geq e_{i,(N)}$, where by definition $e_{i,(1)}=e_{i,i}=0.5$. For $k\in\mathbb{Z}^{+}$, we define the $k$-th critical matching efficiency as
$e_{(k)}:=\begin{cases}\max\limits_{i\in[N]}\left\{e_{i,(k)}\right\},&\text{if }k\leq N,\\ 0,&\text{if }k>N,\end{cases}$
that is, when $k\leq N$, $e_{(k)}$ is the maximum $k$-th largest matching efficiency among all agent types. Since each agent has a self-matching efficiency of $0.5$, we always have $e_{(1)}=0.5$. When $k\geq 2$, this implies that for any agent of type $i\in[N]$, there are at most $k-2$ other agent types (excluding self-match type $i$) with a matching efficiency greater than $e_{(k)}$. For notational convenience, we also define $e_{(k)}=0$ for all $k>N$. It can be easily verified that $e_{(k)}$ is non-increasing in $k\in\mathbb{Z}^{+}$.
Based on the above definition, we now state our main results concerning concavity.
Theorem 3.10
$c(\boldsymbol{\lambda})$ is concave on $\Lambda$ under either of the following conditions:
(i) $\nu=1$, $e_{(3)}\neq 0.5$ and $\underline{\lambda}_{i}>\theta_{i}e_{(3)}/(1-2e_{(3)}),\ \forall i\in[N];$
(ii) $\nu>1$, $e_{(2)}\neq 0.5$ and $\underline{\lambda}_{i}>\theta_{i}e_{(2)}/(1-2e_{(2)}),\ \forall i\in[N].$
Theorem 3.10 discusses two separate cases under which concavity holds for general $N$. In the first case, where all $\theta_{i},\ i\in[N]$ are equal, the lower bound for $\underline{\lambda}_{i}$ depends on $e_{(3)}$. Intuitively, when $e_{(3)}$ is small, each agent type $i$ has high matching efficiency with at most one other type $j$, while the remaining $N-2$ types have low compatibility with $i$. This structure ensures that agents of type $i$ are matched with at most one other type, allowing us to reduce the problem to a series of two-type ($N=2$) subproblems. As a result, the concavity result from Theorem 3.9 can be extended to this setting.
For the second case, where the $\theta_{i}$ may differ, we can show that each agent type is optimally matched only within its own type when the lower bound of $\underline{\lambda}_{i}$ is satisfied. This further simplifies the problem to the single-type case ($N=1$), for which concavity has been established in Lemma 3.1.
The next theorem gives the results related to weak concavity.
Theorem 3.11
$c(\boldsymbol{\lambda})$ is weakly concave on $\Lambda$ under either of the following conditions:
(i) there exists $K\in\mathbb{Z}^{+}$, such that $K\nu\leq 2$, $e_{(K+2)}\neq 0.5$, and
$\underline{\lambda}_{i}>\theta_{i}\frac{e_{(K+2)}}{1-2e_{(K+2)}},\ \forall i\in[N];$
(ii) there exists $K\in\mathbb{Z}^{+}$, such that $K\nu>2$, $e_{(K+2)}\neq 0.5$ and
$\theta_{i}\frac{e_{(K+2)}}{1-2e_{(K+2)}}<\underline{\lambda}_{i}<\overline{\lambda}_{i}<\theta_{i}\max\left\{\frac{1}{K},\frac{1}{K\nu-2}\right\}\left(1+\frac{1}{\nu}\frac{e_{(K+2)}}{1-2e_{(K+2)}}\right),\ \forall i\in[N].$
Theorem 3.11 provides two sufficient conditions of weak concavity, depending on the choice of $K$ and the value of $\nu$. The parameter $K$ determines the maximum number of other agent types that can be matched with agent type $i\in[N]$, and $\nu$ measures the difference in patience levels across agent types. When $K\nu\leq 2$, it suffices to impose only lower bounds $\underline{\lambda}_{i}$, $i\in[N]$, whereas when $K\nu>2$, we pose both lower and upper bounds. As $K$ increases, $e_{(K+2)}$ decreases, leading to lower values for the bounds—thus shifting the feasible region $\Lambda$. The value of $K$ can therefore be selected by comparing these feasible regions and identifying the one that is most practically useful. On the other hand, as $\nu$ decreases—indicating more similar patience levels among agents—the upper bounds (when $K\nu>2$) become less restrictive, expanding the allowable region. Therefore, a $\nu$ close to 1 helps to achieve weak concavity.
From Theorem 3.11, we can get the following corollary, by letting $K=2$ when $\nu=1$ and letting $K=1$ when $21<\nu<$ 2.
Corollary 3.12
$c(\boldsymbol{\lambda})$ is weakly concave on $\Lambda$ under either of the following conditions:
(i) $\nu=1$, $e_{(4)}\neq 0.5$ and $\underline{\lambda}_{i}>\theta_{i}e_{(4)}/(1-2e_{(4)}),\ \forall i\in[N];$
(ii) $21<\nu<$ 2, $e_{(3)}\neq 0.5$ and $\underline{\lambda}_{i}>\theta_{i}e_{(3)}/(1-2e_{(3)}),\ \forall i\in[N].$
Interestingly, Corollary 3.12 closely resembles Theorem 3.10. For the first case, we continue to assume that $\theta_{i},\ i\in[N]$ are all equal, but allow the region $\Lambda$ being larger than that in Theorem 3.10. Specifically, the lower bound of $\underline{\lambda}_{i}$ is now related to $e_{(4)}$ instead of $e_{(3)}$. Under this weaker assumption, we can prove that the value function is weakly concave, if not concave. In the second case, we assume the values of $\theta_{i},\ i\in[N]$ can be different but within a ratio of 2, and also give a condition of weak concavity. Notably, the allowable region is also enlarged compared to Theorem 3.10, since the lower bound for $\underline{\lambda}_{i}$ is determined by $e_{(3)}$ instead of $e_{(2)}$.
From the first case of Corollary 3.12, we can also get the following corollary for the case of $N=3$, since $e_{(4)}=0$ according to our definition. Therefore, weak concavity always holds when there are fewer than or equal to three agent types with identical patience levels.
Corollary 3.13
When $N=3$ and $\theta_{1}=\theta_{2}=\theta_{3}$, $c(\boldsymbol{\lambda})$ is weakly concave on $\Lambda$.
To summarize, Theorems 3.10 and 3.11, along with Corollary 3.12, all suggest that it is easier to achieve (weak) concavity when agents have similar or identical patience levels (i.e., when $\nu$ is close to or equal to $1$). Additionally, a higher overall patience level or a greater heterogeneity in matching efficiencies also leads to a smaller required lower bound of $\underline{\lambda}_{i}$ for each $i\in[N]$. These findings are consistent with the results established in Section 3.1 when $N=2$. Lastly, as an extreme case, when $\theta_{i}=0$, $\forall i\in[N]$, the value function will become linear, as we previously established in Proposition 2.2.
4 Application on Pricing
In this section, we leverage the (weak) concavity of $c(\boldsymbol{\lambda})$ to design an efficient algorithm for solving the upper-level pricing problem. We assume a one-to-one mapping between $\lambda_{i}$ and $p_{i}$ for all $i\in[N]$, as well as concavity of the revenue function with respect to the arrival rate. Both assumptions are standard in the revenue management literature. {assumption} There exists a one-to-one correspondence between $\lambda_{i}$ and $p_{i}$ for all $i\in[N]$. The revenue function $\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})$ is concave on $\Lambda$.
4.1 A Minorize-Maximization Algorithm
Although the revenue function is assumed to be concave, $c(\boldsymbol{\lambda})$ is in general not convex, thus, the overall program (1) is not a convex optimization problem. As shown in Example 7.5, $g(\boldsymbol{\lambda})$ is neither unimodal nor smooth. However, as analyzed in Section 3, $c(\boldsymbol{\lambda})$ tends to be concave or weakly concave. With such properties, we can treat the objective function $g(\boldsymbol{\lambda})$ as a difference of two concave functions:
$g(\boldsymbol{\lambda})=\left(\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})-\frac{1}{2}\rho\|\boldsymbol{\lambda}\|^{2}\right)-\left(c(\boldsymbol{\lambda})-\frac{1}{2}\rho\|\boldsymbol{\lambda}\|^{2}\right),$
where the value of $\rho\geq 0$ is selected such that $c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2$ is concave. We can then adapt the Minorize-Maximization (MM) algorithm to obtain the solution with a convergence guarantee. At each step of the MM algorithm, we construct a surrogate linear function to replace the concave function $c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2$. Denote $\boldsymbol{\lambda}^{(t)}$ as the value of $\boldsymbol{\lambda}$ after the $t$-th iteration. At the $(t+1)$-th iteration, the surrogate linear function is generated based on the supergradient of this concave function (analogous to the subgradient in convex functions; see Lu et al. 2014) at the point of $\boldsymbol{\lambda}^{(t)}$. Let $\boldsymbol{\gamma}^{*}=(\gamma_{i}^{*})_{i\in[N]}$ and $\boldsymbol{\eta}^{*}=(\eta_{i,j}^{*})_{i,j\in[N]}$ be the optimal dual variables corresponding to (2b) and (2c) respectively, then according to the envelop theorem (see e.g., Corollary 5 in Milgrom and Segal 2002), the supergradient, denoted as $\boldsymbol{v}=(v_{i})_{i\in[N]}$, can be given as
$\boldsymbol{v}:=\left(\gamma_{i}^{*}+\sum_{j\in[N]}y_{j}^{*}\eta_{j,i}^{*}-\rho\lambda_{i}^{(t)}\right)_{i\in[N]}\in\partial\left(c(\boldsymbol{\lambda}^{(t)})-\frac{1}{2}\rho\|\boldsymbol{\lambda}^{(t)}\|^{2}\right),$ (3)
and the surrogate linear function is given as
$c(\boldsymbol{\lambda}^{(t)})-\frac{1}{2}\rho\|\boldsymbol{\lambda}^{(t)}\|^{2}+\boldsymbol{v}^{T}\left(\boldsymbol{\lambda}-\boldsymbol{\lambda}^{(t)}\right)\geq c(\boldsymbol{\lambda})-\frac{1}{2}\rho\|\boldsymbol{\lambda}\|^{2},\quad\forall\boldsymbol{\lambda}\in\Lambda,$
where the inequality holds because $c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2$ is a concave function.
Define $Q(\boldsymbol{\lambda}\mid\boldsymbol{\lambda}^{(t)})=\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})-\rho\|\boldsymbol{\lambda}\|^{2}/2-c(\boldsymbol{\lambda}^{(t)})+\rho\|\boldsymbol{\lambda}^{(t)}\|^{2}/2-\boldsymbol{v}^{T}(\boldsymbol{\lambda}-\boldsymbol{\lambda}^{(t)})$. It can be easily verified that $Q(\boldsymbol{\lambda}\mid\boldsymbol{\lambda}^{(t)})$ is a concave function since the original non-concave component in $g(\boldsymbol{\lambda})$ have been substituted with a linear function, and we have
$Q(\boldsymbol{\lambda}\mid\boldsymbol{\lambda}^{(t)})$ $=\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})-\frac{1}{2}\rho\|\boldsymbol{\lambda}\|^{2}-c(\boldsymbol{\lambda}^{(t)})+\frac{1}{2}\rho\|\boldsymbol{\lambda}^{(t)}\|^{2}-\boldsymbol{v}^{T}(\boldsymbol{\lambda}-\boldsymbol{\lambda}^{(t)})$
$\leq\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})-\frac{1}{2}\rho\|\boldsymbol{\lambda}\|^{2}-\left(c(\boldsymbol{\lambda})-\frac{1}{2}\rho\|\boldsymbol{\lambda}\|^{2}\right)=g(\boldsymbol{\lambda}),\quad\forall\boldsymbol{\lambda}\in\Lambda,$
and
$Q(\boldsymbol{\lambda}^{(t)}\mid\boldsymbol{\lambda}^{(t)})=g(\boldsymbol{\lambda}^{(t)}).$
We update the value of $\boldsymbol{\lambda}$ by letting $\boldsymbol{\lambda}^{(t+1)}\leftarrow\arg\max_{\boldsymbol{\lambda}\in\Lambda}Q(\boldsymbol{\lambda}\mid\boldsymbol{\lambda}^{(t)})$, where the maximization of $Q(\boldsymbol{\lambda}\mid\boldsymbol{\lambda}^{(t)})$ can be solved using standard convex optimization techniques. Then we have
$g(\boldsymbol{\boldsymbol{\lambda}}^{(t+1)})\geq Q(\boldsymbol{\lambda}^{(t+1)}\mid\boldsymbol{\lambda}^{(t)})\geq Q(\boldsymbol{\lambda}^{(t)}\mid\boldsymbol{\lambda}^{(t)})=g(\boldsymbol{\lambda}^{(t)}),$
which guarantees that the objective function value is non-decreasing at each iteration. This monotonic improvement ensures monotone convergence to a local optimum (Hunter and Lange 2004).
Algorithm 1 Minorize-Maximization (MM) Algorithm
1:Input: Initial guess $\boldsymbol{\lambda}^{(0)}\in\Lambda$, $\varepsilon>0$, $\delta_{\text{MM}}>0$.
2:Initialize: $t\leftarrow 0$
3:repeat
4:     Solve $c(\boldsymbol{\lambda}^{(t)})$, obtain optimal primal variables $\boldsymbol{y}^{*}=(y_{i}^{*})_{i\in[N]}$, $\boldsymbol{x}^{*}=(x_{i,j}^{*})_{i,j\in[N]}$ and dual variables $\boldsymbol{\gamma}^{*}=(\gamma_{i}^{*})_{i\in[N]}$, $\boldsymbol{\eta}^{*}=(\eta_{i,j}^{*})_{i,j\in[N]}$.
5:     $\rho\leftarrow 0$
6:     repeat
7:         Calculate $\boldsymbol{v}=(v_{i})_{i\in[N]}$, where $v_{i}=\gamma_{i}^{*}+\sum_{j\in[N]}y_{j}^{*}\eta_{j,i}^{*}-\rho\lambda_{i}^{(t)}.$
8:         Define $Q(\boldsymbol{\lambda}\mid\boldsymbol{\lambda}^{(t)})=\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})-\rho\|\boldsymbol{\lambda}\|^{2}/2-c(\boldsymbol{\lambda}^{(t)})+\rho\|\boldsymbol{\lambda}^{(t)}\|^{2}/2-\boldsymbol{v}^{T}(\boldsymbol{\lambda}-\boldsymbol{\lambda}^{(t)})$.
9:         Update $\boldsymbol{\lambda}^{(t+1)}\leftarrow\arg\max_{\boldsymbol{\lambda}\in\Lambda}Q(\boldsymbol{\lambda}\mid\boldsymbol{\lambda}^{(t)})$.
10:         if $g(\boldsymbol{\lambda}^{(t+1)})<g(\boldsymbol{\lambda}^{(t)})$ then
11:              Update $\rho\leftarrow\rho+\delta_{\text{MM}}$.
12:         end if
13:     until $g(\boldsymbol{\lambda}^{(t+1)})\geq g(\boldsymbol{\lambda}^{(t)})$
14:     $t\leftarrow t+1$
15:until $|g(\boldsymbol{\lambda}^{(t+1)})-g(\boldsymbol{\lambda}^{(t)})|<\varepsilon$
16:Output: $\boldsymbol{\lambda}^{(t)}$
Summarizing the ideas above, we present the pseudocode of the MM algorithm in Algorithm 1. Note that, as shown in Section 3, although the cost function $c(\boldsymbol{\lambda})$ is weakly concave under certain conditions, the value of $\rho$ is not specified. To address this, we slightly modify the standard MM algorithm by introducing a search step for $\rho$. Initially, we set $\rho=0$ (Line 5) and perform one iteration to update $\boldsymbol{\lambda}$ (Lines 7–9). If the objective function does not increase after this iteration, it indicates that $c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2$ is not concave; we then increment $\rho$ by $\delta_{\text{MM}}$ and repeat the update. Since $c(\boldsymbol{\lambda})$ is weakly concave, it is guaranteed that after a finite number of increments, we will find a $\rho$ such that $c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2$ is concave, ensuring that the objective function $g(\boldsymbol{\lambda})$ increases after each iteration. By construction, each iteration yields an increase in the objective, and thus the algorithm is guaranteed to converge. It is worth noting that in the ridesharing case study presented in the next subsection, we never observe the need to increase $\rho$; Algorithm 1 consistently converges with $\rho=0$, requiring no stepsize tuning on real data. The same behavior is observed across numerous synthetically generated instances. In contrast, alternative methods for solving (2), such as the projected gradient-based approach, are highly sensitive to stepsize selection and require careful tuning, as will be demonstrated in Section 4.2.
4.2 Chicago Ridesharing Case Study
In this subsection, we conduct numerical experiments using the Chicago ridesharing dataset. The dataset, provided by Chicago Data Portal, [^1] https://data.cityofchicago.org/Transportation/Transportation-Network-Providers-Trips-2018-2022-/m6dm-c72p/about_data contains trips reported by ridesharing companies within the city of Chicago. The dataset divides the city (excluding the O’Hare International Airport) into 76 community areas, and each data point includes the passenger’s OD pair, represented by the corresponding communities’ centroid locations. We use typical rush-hour data (i.e., 7:30 AM - 8:30 AM on Mondays), spanning 7 weeks between October 7 and November 18, 2019, prior to the COVID-19 pandemic, to conduct the experiment.
To aggregate riders into distinct types, we apply the $K$-means clustering algorithm to the origin-destination (OD) pairs of all trips, with the number of clusters set to $N\in\{100,150,200\}$. For each cluster (i.e., rider type $i\in[N]$), we treat the cluster center as the representative OD for that type, and the average number of requests per hour within the cluster as the maximum possible hourly arrival rate $\overline{\lambda}_{i}$. Given the OD of each rider type, we compute the corresponding solo trip distances and pooled trip distances (in miles), denoted as $[l_{(i)}]_{i\in[N]}$ and $[l_{(i,j)}]_{i,j\in[N]}$, respectively. We let $c_{(i,j)}=cl_{(i,j)}$, where $c$ is the per-mile cost. We consider three possible values of $c\in\{0.7,0.9,1.1\}$ in the experiment, with a unit of dollar per mile. We assume in this section that the riders’ per-mile willingness to pay is uniformly distributed in $[0,1]$, then we have $p_{i}(\lambda_{i})=l_{(i)}\left(1-\lambda_{i}/\overline{\lambda}_{i}\right)$, $\forall i\in[N]$. To justify the robustness of our results, we also test an alternative exponential demand model in Appendix 8.1.
We compare the following two methods to solve the pricing problem (1):
  • Projected Gradient (PG) method. We use PG as the baseline method. At each iteration, $\boldsymbol{\lambda}$ is updated along with the opposite direction of the gradient while restricting within $\Lambda$. When $g(\boldsymbol{\lambda})$ is differentiable at $\boldsymbol{\lambda}$, its gradient is given by
    $\nabla g(\boldsymbol{\lambda})=\left[l_{(i)}\left(1-\frac{2\lambda_{i}}{\overline{\lambda}_{i}}\right)-\gamma_{i}^{*}-\sum_{j\in[N]}y_{j}^{*}\eta_{j,i}^{*}\right]_{i\in[N]},$ (4)
    where $\boldsymbol{\gamma}^{*}=(\gamma_{i}^{*})_{i\in[N]}$ and $\boldsymbol{\eta}^{*}=(\eta_{i,j}^{*})_{i,j\in[N]}$ are the optimal dual variables as defined in Algorithm 1. If $g(\boldsymbol{\lambda})$ is non-differentiable, we continue to use (4) to perform the iteration, interpreting it as a subgradient or supergradient when such gradients exist. The initial step size $\delta_{\text{PG}}$ is chosen as $\{1,10,100\}$. We also dynamically adjust the step size, by checking whether the objective function increases after each iteration. If $g(\boldsymbol{\lambda}^{(t+1)})<g(\boldsymbol{\lambda}^{(t)})$, we decrease the step size by half.
  • Minorize-Maximization (MM) method. This is our Algorithm 1. Note that since we now have an explicit form of $p_{i}(\lambda_{i})$, we can give a closed form solution of $\boldsymbol{\lambda}^{(t+1)}\leftarrow\arg\max_{\boldsymbol{\lambda}\in\Lambda}Q(\boldsymbol{\lambda}\mid\boldsymbol{\lambda}^{(t)})$ (Line 9 of Algorithm 1) using its first-order condition:
    $\lambda_{i}^{(t+1)}=\max\left\{\underline{\lambda}_{i},\min\left\{\overline{\lambda}_{i},\overline{\lambda}_{i}\frac{l_{(i)}-\left(\gamma_{i}^{*}+\sum_{j\in[N]}y_{j}^{*}\eta_{j,i}^{*}\right)+\rho\lambda_{i}^{(t)}}{2l_{(i)}+\rho\overline{\lambda}_{i}}\right\}\right\}.$
    Besides, since we never observe the need to increase $\rho$ in our experiments—i.e., Algorithm 1 consistently converges with $\rho=0$—the choice of $\delta_{\text{MM}}$ in Algorithm 1 is not relevant.
For both algorithms, we set $\underline{\lambda}_{i}=10^{-3}$ for all $i\in[N]$. Initial solution $\boldsymbol{\lambda}^{(0)}$ is randomly sampled from the feasible set $\Lambda=[\underline{\boldsymbol{\lambda}},\overline{\boldsymbol{\lambda}}]$ and kept the same across all algorithms within the same run. The stopping criterion is defined as $|g(\boldsymbol{\lambda}^{(t+1)})-g(\boldsymbol{\lambda}^{(t)})|<\varepsilon=10^{-3}$. We also record the running time of each algorithm; if it exceeds 20 minutes, the algorithm terminates and returns the result from the latest iteration. We report and compare the following three metrics, averaged over three different random seeds [^2] The randomness comes from the initialization of $K$-means clustering, as well as the sampling of $\boldsymbol{\lambda}^{(0)}$ and $(\theta_{i})_{i\in[N]}$.: (i) running time (in seconds); (ii) number of iterations; and (iii) final objective value.
Table 3: Numerical results of solving (1) when $\theta_{1}=\cdots=\theta_{N}=\theta$.
$c$ (\$/mile) $0.7$ $0.9$ $1.1$
$\theta$ (min-1) $1/5$ $1/3$ $1$ $2$ $1/5$ $1/3$ $1$ $2$ $1/5$ $1/3$ $1$ $2$
$N=100$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 1036 905 970 674 891 887 666 666 1151 1167 889 598
PG ($\delta_{\text{PG}}=10$) 333 263 281 287 273 292 263 300 608 681 1115 435
PG ($\delta_{\text{PG}}=100$) 371 362 380 400 342 347 358 377 1048 778 723 414
MM 59 59 58 59 66 70 77 74 121 119 187 220
Num. Iterations PG ($\delta_{\text{PG}}=1$) 86.3 76.7 82.7 57.7 76.7 76.0 57.3 57.0 98.3 98.3 76.0 51.0
PG ($\delta_{\text{PG}}=10$) 27.3 22.3 24.0 24.3 23.3 25.0 22.7 25.7 52.0 57.7 95.0 37.0
PG ($\delta_{\text{PG}}=100$) 31.0 30.7 32.3 34.0 29.3 29.7 31.0 32.3 89.3 66.3 61.7 35.3
MM 5.0 5.0 5.0 5.0 5.7 6.0 6.7 6.3 10.3 10.0 16.0 18.7
Objective Function PG ($\delta_{\text{PG}}=1$) 64.05 58.93 45.88 37.99 39.55 33.88 20.67 13.83 16.70 13.22 4.77 0.92
PG ($\delta_{\text{PG}}=10$) 64.26 59.13 46.00 38.11 39.62 33.98 20.81 14.07 12.46 10.28 4.50 1.57
PG ($\delta_{\text{PG}}=100$) 64.26 59.13 46.00 38.10 39.65 33.97 20.85 14.07 13.23 10.50 4.28 1.59
MM 64.26 59.13 46.00 38.11 39.65 33.98 20.85 14.07 19.66 14.42 4.95 1.69
$N=150$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 1200 1200 869 917 1038 720 1000 993 1200 1200 1004 1172
PG ($\delta_{\text{PG}}=10$) 722 770 746 775 701 702 698 727 1185 992 1063 1048
PG ($\delta_{\text{PG}}=100$) 899 891 836 868 873 866 900 980 1203 969 1200 1128
MM 132 133 130 134 159 177 176 159 378 433 541 314
Num. Iterations PG ($\delta_{\text{PG}}=1$) 45.7 45.7 33.3 34.7 39.0 27.3 38.0 38.0 46.0 45.3 38.3 44.0
PG ($\delta_{\text{PG}}=10$) 27.3 29.0 28.7 29.3 26.0 26.7 26.3 27.3 45.3 38.0 40.3 39.7
PG ($\delta_{\text{PG}}=100$) 34.0 33.7 32.0 33.0 33.0 33.0 34.0 36.7 46.0 37.0 45.3 43.0
MM 5.0 5.0 5.0 5.0 6.0 6.7 6.7 6.0 14.3 16.3 20.7 11.7
Objective Function PG ($\delta_{\text{PG}}=1$) 61.95 56.48 44.57 37.76 36.61 30.70 19.47 13.63 15.01 11.46 4.37 1.52
PG ($\delta_{\text{PG}}=10$) 62.42 56.95 44.76 37.96 36.85 31.08 19.72 13.89 14.01 10.36 4.36 1.09
PG ($\delta_{\text{PG}}=100$) 62.42 56.95 44.65 37.96 36.85 31.08 19.72 13.90 13.48 10.31 4.28 1.10
MM 62.42 56.95 44.76 37.96 36.85 31.08 19.72 13.90 16.75 12.13 4.75 1.74
$N=200$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 1132 1128 838 1156 1148 1131 1162 1146 1200 1134 1099 1133
PG ($\delta_{\text{PG}}=10$) 1200 1200 1200 1200 1200 1200 1200 1200 1200 1179 1200 1200
PG ($\delta_{\text{PG}}=100$) 1200 1200 1200 1200 1200 1200 1200 1200 1200 1200 1200 1200
MM 232 234 232 236 285 311 297 286 666 642 661 499
Num. Iterations PG ($\delta_{\text{PG}}=1$) 24.3 24.3 18.0 24.7 24.3 24.3 25.0 24.3 25.3 24.3 23.7 24.0
PG ($\delta_{\text{PG}}=10$) 26.0 26.0 26.0 26.0 25.7 26.0 26.0 25.7 25.3 25.3 26.0 26.0
PG ($\delta_{\text{PG}}=100$) 25.7 26.0 26.0 26.0 25.7 26.0 26.0 25.7 25.7 25.7 26.0 26.0
MM 5.0 5.0 5.0 5.0 6.0 6.7 6.3 6.0 14.0 13.7 14.0 10.3
Objective Function PG ($\delta_{\text{PG}}=1$) 60.25 54.99 18.79 37.66 34.62 29.32 19.01 13.42 14.15 10.87 4.06 1.19
PG ($\delta_{\text{PG}}=10$) 60.75 55.47 44.22 37.81 34.81 29.58 19.27 13.68 13.32 10.26 4.14 1.33
PG ($\delta_{\text{PG}}=100$) 56.67 54.40 43.22 36.91 33.27 29.14 18.37 11.62 10.20 7.16 2.57 -6.12
MM 60.77 55.51 44.27 37.86 34.88 29.59 19.29 13.71 15.28 11.37 4.58 1.66
Table 3 presents the results of the first set of experiments, where we assume $\theta_{1}=\cdots=\theta_{N}=\theta\in\{1/5,1/3,1,2\}$ per minute. Accordingly, the mean sojourn times for riders are $\{5,3,1,0.5\}$ minutes, respectively. The results demonstrate that the MM algorithm substantially reduces both the running time and the number of iterations compared to the PG method. Compared to PG with three different step sizes ($\delta_{\text{PG}}=1,10,100$), the MM algorithm reduces average running time by 76.3%, 70.9%, and 73.0%, respectively, and reduces the average number of iterations by 81.6%, 73.0%, and 75.7%. Since the running time is capped at 20 minutes, the actual time savings may be even greater. Moreover, within the maximum allowed running time, the MM algorithm consistently achieves a higher objective value—on average, 4.31%, 2.17%, and 5.00% higher than the PG methods. We also observe that the performance of PG is sensitive to step size selection: although a step size of 10 yields the best average performance, as shown in Table 4, it is not optimal in every instance: A step size of 1 performs better when $c=1.1$. In contrast, MM demonstrates strong robustness and outperforms PG across all cases and step-size rules.
Table 4: Averaged objective values under different values of $c$, aggregated from Table 3.
$c$ (\$/mile) 0.7 0.9 1.1
PG ($\delta_{\text{PG}}=1$) 48.28 25.39 8.19
PG ($\delta_{\text{PG}}=10$) 50.65 25.61 7.31
PG ($\delta_{\text{PG}}=100$) 50.06 25.21 6.05
MM 50.67 25.63 9.08
It can also be observed that as the number of types $N$ increases, the running time of both PG and MM increases, primarily because solving $c(\boldsymbol{\lambda})$ involves $N+N^{2}$ variables at each iteration. However, the number of iterations required by the MM algorithm does not scale proportionally with $N$, and in many cases remains nearly constant at a single-digit value. Even when $N=200$, the MM algorithm remains efficient and consistently produces results within a short CPU time.
Table 5: Numerical results of solving (1) when $\theta_{i}$, $i\in[N]$ is not equal.
$c$ (\$/mile) $0.7$ $0.9$ $1.1$
$(\underline{\theta},\overline{\theta})$ (min-1) $(1/5,1/3)$ $(1,2)$ $(1/5,1/3)$ $(1,2)$ $(1/5,1/3)$ $(1,2)$
$N=100$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 585 559 554 562 581 678
PG ($\delta_{\text{PG}}=10$) 321 282 302 286 831 662
PG ($\delta_{\text{PG}}=100$) 366 394 347 378 1120 709
MM 58 58 70 82 138 183
Num. Iterations PG ($\delta_{\text{PG}}=1$) 50.7 48.7 47.7 48.3 50.3 58.7
PG ($\delta_{\text{PG}}=10$) 27.7 24.3 25.7 24.3 72.0 57.7
PG ($\delta_{\text{PG}}=100$) 31.7 34.0 29.3 32.0 97.7 61.3
MM 5.0 5.0 6.0 7.0 12.0 15.7
Objective Function PG ($\delta_{\text{PG}}=1$) 61.49 41.58 36.64 16.99 15.08 2.97
PG ($\delta_{\text{PG}}=10$) 61.66 41.73 36.79 17.11 12.46 2.62
PG ($\delta_{\text{PG}}=100$) 61.66 41.70 36.79 17.15 11.74 2.02
MM 61.66 41.73 36.79 17.15 16.96 3.09
$N=150$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 899 974 815 842 981 898
PG ($\delta_{\text{PG}}=10$) 764 819 691 659 1051 836
PG ($\delta_{\text{PG}}=100$) 856 857 874 893 1200 1125
MM 132 130 155 166 414 305
Num. Iterations PG ($\delta_{\text{PG}}=1$) 34.0 37.3 31.7 32.7 37.7 34.3
PG ($\delta_{\text{PG}}=10$) 28.7 31.0 26.3 25.3 41.0 32.0
PG ($\delta_{\text{PG}}=100$) 32.7 32.7 33.0 34.3 47.0 42.3
MM 5.0 5.0 6.0 6.3 15.7 11.7
Objective Function PG ($\delta_{\text{PG}}=1$) 59.40 40.88 33.64 16.35 13.38 2.73
PG ($\delta_{\text{PG}}=10$) 59.66 41.15 33.93 16.49 12.41 2.84
PG ($\delta_{\text{PG}}=100$) 59.66 41.16 33.93 16.66 10.97 2.94
MM 59.66 41.16 33.93 16.68 14.36 3.08
$N=200$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 1200 1031 1137 1148 1200 1197
PG ($\delta_{\text{PG}}=10$) 1200 1200 1200 1200 1200 1200
PG ($\delta_{\text{PG}}=100$) 1200 1200 1200 1200 1200 1200
MM 235 237 278 280 722 479
Num. Iterations PG ($\delta_{\text{PG}}=1$) 26.3 22.0 25.0 24.7 26.0 25.3
PG ($\delta_{\text{PG}}=10$) 26.3 25.7 26.3 26.0 26.7 25.3
PG ($\delta_{\text{PG}}=100$) 26.7 25.3 26.3 26.0 26.0 25.3
MM 5.0 5.0 6.0 6.0 15.3 10.0
Objective Function PG ($\delta_{\text{PG}}=1$) 57.51 40.15 31.91 16.10 12.63 2.60
PG ($\delta_{\text{PG}}=10$) 58.08 40.90 32.19 16.40 11.70 2.32
PG ($\delta_{\text{PG}}=100$) 55.01 36.09 31.25 12.72 4.90 -0.54
MM 58.13 40.96 32.20 16.42 13.35 2.93
Table 5 presents the experimental results for the setting where $\theta_{i}$ varies across agent types. For type-$i$ agents, $\theta_{i}$ is randomly drawn from a uniform distribution $U[\underline{\theta},\overline{\theta}]$, with two parameter choices: $(\underline{\theta},\overline{\theta})\in\{(1/5,1/3),(1,2)\}$. The results exhibit a similar trend to that observed in Table 1. Compared to three PG variants with $\delta_{\text{PG}}=1,10,100$, the MM algorithm reduces running time by 74.0%, 72.0% and 74.7%; decreases the number of iterations by 77.7%, 74.2% and 77.7%; and improves the objective value by 1.63%, 1.95% and 7.23% within the 20-minute maximum runtime, respectively. The performance of PG is still sensitive to the step size: $\delta_{\text{PG}}=10$ performs better when $c\leq 0.9$, whereas $\delta_{\text{PG}}=1$ is preferable when $c=1.1$.
We also conduct simulation studies by implementing the optimal prices given by our MM algorithm and comparing with the baseline pricing scheme when agents’ finite patience is not incorporated (i.e., when (2c) is relaxed, as in the LP formulation of Eom and Toriello 2024). We find that our pricing results consistently outperform the baseline. For details of the simulation, please refer to Appendix 8.2.
5 Conclusion
Motivated by growing pricing applications in centralized matching markets, we study the behavior of the value function of the LP introduced by Aouad and Sarıtaç (2022) for cost-minimizing dynamic stochastic matching, with a particular focus on its (weak) concavity with respect to agents’ arrival rates. We establish a key result: it is sufficient for all optimal solutions of the LP to have strictly positive unmatched rates to ensure weak concavity. Leveraging this insight, we demonstrate that weak concavity is guaranteed when there are at most three agent types sharing the same patience level. For the general case with an arbitrary number of types, (weak) concavity is more likely to hold when agents are either highly patient, have similar patience levels, or exhibit substantial heterogeneity in matching efficiencies.
We then turn our attention to the pricing application and design an efficient MM algorithm that requires little stepsize tuning. To evaluate its performance, we apply our MM algorithm to the Chicago ridesharing dataset. Experimental results demonstrate that, compared to a projected gradient-based method whose performance is highly sensitive to stepsize selection, our approach significantly reduces both running time and the number of iterations, while simultaneously achieving better optimality.
[…]
The authors thank Ying Cui from UC Berkeley for helpful discussions. Chen acknowledges the support of the INFORMS Transportation Science and Logistics Society. Yan also acknowledges the support of the National Science Foundation under Award Number 2517861.